### v1

# Rethinking Approximate Gaussian Inference in Classification

Abstract
In classification tasks, softmax functions are ubiquitously used as output activations to produce predictive probabilities. Such outputs only capture aleatoric uncertainty. To capture epistemic uncertainty, approximate Gaussian inference methods have been proposed, which output Gaussian distributions over the logit space. Predictives are then obtained as the expectations of the Gaussian distributions pushed forward through the softmax. However, such softmax Gaussian integrals cannot be solved analytically, and Monte Carlo (MC) approximations can be costly and noisy. We propose a simple change in the learning objective which allows the exact computation of predictives and enjoys improved training dynamics, with no runtime or memory overhead. This framework is compatible with a family of output activation functions that includes the softmax, as well as element-wise normCDF and sigmoid. Moreover, it allows for approximating the Gaussian pushforwards with Dirichlet distributions by analytic moment matching. We evaluate our approach combined with several approximate Gaussian inference methods (Laplace, HET, SNGP) on large- and small-scale datasets (ImageNet, CIFAR-10), demonstrating improved uncertainty quantification capabilities compared to softmax MC sampling. Code is available at github.com/bmucsanyi/probit.
Deep learning, Uncertainty quantification, Probabilistic modeling, Bayesian inference
[…]
[ linewidth=1pt, roundcorner=3pt, innerleftmargin=4pt, innerrightmargin=4pt, innertopmargin=-5pt, innerbottommargin=4pt, leftmargin=0pt, rightmargin=0pt, skipabove=2pt, skipbelow=0pt, ]eqbox
1 Introduction
Figure 1: Mean KL divergence between the ‘true’ predictive (approximated with a $10000$ sample MC approximation) and different predictive approximations on a synthetic data set. This synthetic dataset consists of 100 logit means $\bm{\mu}$ and standard deviations $\bm{\sigma}$, whose components are i.i.d. uniformly distributed. For a fair comparison we scale the logit dataset for the different activations to match the activations to first order, so that $\mu_{c}\in[-1,1]$, $\sigma_{c}\in[0,1]$ for sigmoid, $\mu_{c}\in[-1/2-\log 2,1/2-\log 2]$, $\sigma_{c}\in[0,1/2]$ for softmax, $\mu_{c}\in[-\sqrt{\pi/8},\sqrt{\pi/8}]$, $\sigma_{c}\in[0,\pi/8]$ for normCDF. The MC approximations are capped at a fixed computational budget of $10000$ class samples, i.e. $\lceil 10000/C\rceil$ samples. We see that, even without enforcing the constraint Equation 18, our approximations are the only ones that do not become worse with the number of classes. See Appendix C for theoretical analyses of the quality of the different approximations.
Given an input space $\mathcal{X}$, $C$ classes and training data $(x_{n},c_{n})_{n=1}^{N}\subset\mathcal{X}\times\{1,\dots,C\}$, the goal of probabilistic classification is to learn a function $\bm{h}\colon\mathcal{X}\to\Delta^{C-1}$, where
$\Delta^{C-1}=\{(p_{1},\dots,p_{C})\in[0,1]^{C}:p_{1}+\dots+p_{C}=1\}$ (1)
is the $(C-1)$-dimensional probability simplex. The model $\bm{h}$ outputs the probability of inputs $x\in\mathcal{X}$ belonging to each class, $h_{c}(x)=p(c\mid x)$. To learn such a map to the simplex, $\bm{h}$ is typically defined as a composition
$\bm{h}\colon\mathcal{X}\xrightarrow{\quad\bm{f}\quad}\mathbb{R}^{C}\xrightarrow{\operatorname{\mathbf{softmax}}}\Delta^{C-1}$ (2)
where a function $\bm{f}$, learned from the training data, is mapped through the softmax activation function. In this case, $\mathbb{R}^{C}$ is called the logit space, and, for $x\in\mathcal{X}$, $\bm{f}(x)$ is a logit.
With the usual assumption of i.i.d. data, the likelihood under the model $\bm{h}$ is given by ($\delta$ is the Kronecker delta)
$p((c_{n})_{n=1}^{N}\mid(x_{n})_{n=1}^{N})$ $=\prod_{n=1}^{N}p(c_{n}\mid x_{n})$ (3)
$=\prod_{n=1}^{N}\prod_{c=1}^{C}h_{c}(x_{n})^{\delta_{c_{n},c}}.$
This yields a natural loss function for classification problems, the cross-entropy (CE) loss
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ $=-\log(p((c_{n})_{n=1}^{N}\mid(x_{n})_{n=1}^{N}))$ (4)
$=-\sum_{n=1}^{N}\sum_{c=1}^{C}\delta_{c_{n},c}\log(h_{c}(x_{n})).$
For typical models $\bm{h}$ trained through such approximate maximum likelihood estimation, an output probability vector $\bm{h}(x)\in\Delta^{C-1}$ should be interpreted as an estimate of $p_{\text{gen}}(\bm{y}\mid x)$, the probability under the generative model. Notably, the output probability does not take into account the uncertainty of the model due to the finite nature of the data. In the uncertainty quantification formalism, such models can only estimate aleatoric uncertainty and disregard epistemic uncertainty (Hüllermeier & Waegeman, 2021).
The probabilistic way to capture epistemic uncertainty is to require the model to output a second-order distribution (as in “a distribution over probability distributions”). That is, for each input $x\in\mathcal{X}$, the model should output a probability measure over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$, as opposed to a point estimate in the simplex $\bm{h}(x)\in\Delta^{C-1}$ (Sensoy et al., 2018; Charpentier et al., 2020).
A number of methods for such distributional uncertainty quantification in classification rely on approximate Gaussian inference in logit space. That is, $\mathrm{h}$ is written as a composition
$\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\operatorname{\mathbf{softmax}}_{*}}\mathcal{P}(\Delta^{C-1})$ (5)
where $f$ is learned from the training data, $\mathcal{G}(\mathbb{R}^{C})$ is the set of Gaussian measures on the logit space $\mathbb{R}^{C}$, and $\operatorname{\mathbf{softmax}}_{*}$ pushes forward the Gaussian probability measures through the softmax. For example, Heteroscedastic Classifiers (HET) (Collier et al., 2021) learn the output means and covariances of $\mathrm{f}$ explicitly with a neural network. In linearised Laplace approximations (Daxberger et al., 2021), only the means are learned explicitly with a neural network, while the covariances are obtained post-hoc from that neural network by approximating its parameter posterior distribution with a Gaussian, and then pushing it forward to logit space by linearising the network in the parameters. Last-layer Laplace approximations (Kristiadi et al., 2020) are a variant of linearised Laplace approximations where only the posterior over the last-layer parameters is Gaussian-approximated, requiring no linearisation of the network. Spectral-Normalized Gaussian processes (SNGP) (Liu et al., 2020) are neural networks with last layers that are approximately Gaussian processes, through random feature expansions and Laplace approximations.
To obtain predictive probabilities from a distributional classifier as in Equation 5, we marginalise it out by the measure-theoretic change of variables, using $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$:
$\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]$ $=\int_{\Delta^{C-1}}p\;\mathrm{h}(x)(dp)$ (6)
$=\int_{\mathbb{R}^{C}}\operatorname{\mathbf{softmax}}(y)\;\mathrm{f}(x)(dy).$
In addition to predictive probabilities, the distributional framework allows for the acquisition of other quantities of interest, such as variances, entropy, and information-theoretic decompositions of the predictive uncertainty into aleatoric and epistemic parts (Depeweg et al., 2018; Mukhoti et al., 2023; Wimmer et al., 2023). These are usually obtained by Monte Carlo (MC) sampling from the Gaussian measure $\mathrm{f}(x)$. Indeed, while the probability density of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ can be obtained in closed form using a change of variables formula (Atchison & Shen, 1980), this does not seem of practical use. Even predictives cannot be computed analytically. On the other hand, given a variance, the runtime and memory cost required for an MC approximation to achieve an estimator of said variance grows linearly with the number of classes. Crucially, this computational cost is not limited to training time, but also to inference time, when the computational budget may be much more limited. This limits the application of such Gaussian inference methods in classification tasks with large number of classes, such as those common in computer vision and natural language processing.
1.1 Proposed Recipe
We can summarise our framework as a 4-step recipe:
  1. 1.
    Choose an approximate Gaussian inference method: a method that outputs logit covariances.
  2. 2.
    Choose an output activation for which the corresponding Gaussian integral can be solved (Section 2).
  3. 3.
    Train with an objective that enforces a certain normalisation constraint (Section 4).
  4. 4.
    At inference time, use our analytic formulas (see Section 2 for the predictives, Section 3 for other quantities of interest).
2 Gaussian Inference with Analytic Predictives
2.1 Motivation from Softmax Models
A key difficulty in obtaining exact or closed-form approximate predictive probabilities of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ is that pushforwards of Gaussian distributions through the softmax are intractable. The softmax function is the composition $\bm{\operatorname{\mathbf{softmax}}}=\bm{n}\circ\bm{\exp}$, where $\bm{n}$ is the normalisation function given by $\bm{n}(\bm{q}):=\bm{q}/(q_{1}+\dots+q_{C})$ and $\bm{\exp}$ is applied element-wise. Thus, $\mathrm{h}$ can be written as the composition (c.f. Equation 5)
$\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\bm{\exp}_{*}}\mathcal{P}(\mathbb{R}_{>0}^{C})\xrightarrow{\bm{n}_{*}}\mathcal{P}(\Delta^{C-1}).$ (7)
Now it is easy to see that (Section A.1)
$\int_{\mathbb{R}}\exp(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)=\exp\left(\mu+\frac{\sigma^{2}}{2}\right)$ (8)
and thus
$\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}[\bm{Q}]$ $=\int_{\mathbb{R}^{C}}\bm{\exp}(\bm{y})\;\mathrm{f}(x)(d\bm{y})$ (9)
$=\bm{\exp}\left(\bm{\mu}(x)+\frac{\bm{\sigma}^{2}(x)}{2}\right)$
where $\bm{\mu}(x)$ and $\bm{\sigma}^{2}(x)$ are, respectively, the mean and the diagonal of the covariance of $\mathrm{f}(x)$, and the vector division is element-wise. Note that no approximations are made so far.
We can use this to approximate the predictive of the model in closed form:
$\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]$ $=\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}\left[\frac{\bm{Q}}{\sum_{c=1}^{C}Q_{c}}\right]$ (10)
$\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}[\bm{Q}]}{\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}\left[\sum_{c=1}^{C}Q_{c}\right]}$
$=\frac{\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}[\bm{Q}]}{\sum_{c=1}^{C}\mathbb{E}_{Q_{c}\sim\exp_{*}\mathrm{f}_{c}(x)}\left[Q_{c}\right]}$
$=\frac{\bm{\exp}\left(\bm{\mu}(x)+\frac{\bm{\sigma}^{2}(x)}{2}\right)}{\sum_{c=1}^{C}\exp\left(\mu_{c}(x)+\frac{\sigma^{2}_{c}(x)}{2}\right)}.$
2.2 The General Framework
In fact, we can apply the previous argumentation to a general family of output activation functions: suppose we can write the output activation as $\bm{n}\circ\bm{\varphi}$, where $\bm{\varphi}$ is the element-wise application of some activation function $\varphi\colon\mathbb{R}\to\mathbb{R}_{>0}$. So we have (c.f. Equation 7)
$\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\bm{\varphi}_{*}}\mathcal{P}(\mathbb{R}_{>0}^{C})\xrightarrow{\bm{n}_{*}}\mathcal{P}(\Delta^{C-1}).$ (11)
If we can solve the one-dimensional integrals (c.f. Equation 8)
$\int_{\mathbb{R}}\varphi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)$ (12)
then we can compute the approximate predictive (c.f. Equation 10)
$\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{Q}]}{\sum_{c=1}^{C}\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}\left[Q_{c}\right]}.$ (13)
For instance, taking $\varphi=\Phi$, the standard Normal cumulative distribution function (normCDF) [^1] The resulting model is distinct from a multinomial probit model, see Appendix B., we have by a classical result (Section A.2),
$\int_{\mathbb{R}}\Phi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{1+\sigma^{2}}}\right)$ (14)
and thus
$\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]\approx\frac{\bm{\Phi}\left(\frac{\bm{\mu}(x)}{\sqrt{1+\bm{\sigma}^{2}(x)}}\right)}{\sum_{c=1}^{C}\Phi\left(\frac{\mu_{c}(x)}{\sqrt{1+\sigma^{2}_{c}(x)}}\right)}.$ (15)
Taking instead $\varphi=\rho$, the logistic sigmoid, we can approximate Equation 12 with the probit approximation (Section A.3)
$\int_{\mathbb{R}}\rho(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)$ (16)
yielding in this case
$\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]\approx\frac{\bm{\rho}\left(\frac{\bm{\mu}(x)}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}(x)}}\right)}{\sum_{c=1}^{C}\rho\left(\frac{\mu_{c}(x)}{\sqrt{1+\frac{\pi}{8}\sigma^{2}_{c}(x)}}\right)}.$ (17)
2.3 Making the Approximations Exact
At this point, the approximate Equation 13 may seem ad-hoc. However, note that Equation 13 is exact when $\sum_{c=1}^{C}Q_{c}$ is a constant random variable, as can be seen from the derivation in Equation 10. Thus, we would like to enforce a constraint in the model of the form
$\sum_{c=1}^{C}Q_{c}$ $=s(x)$ $\text{where }\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)$ (18)
where $s\colon\mathcal{X}\to\mathbb{R}$ is a deterministic function. In other words, the law of $\sum_{c=1}^{C}Q_{c}$ is a Dirac measure for all $x$.
In neural networks, constraints can be learned through appropriate choices of learning objectives during training, ensuring that they are satisfied within the training data distribution. In Section 4, we describe how this can be achieved depending on $\bm{\varphi}$ and the method that is used to obtain the mapping to Gaussian distributions $\mathrm{f}$.
Given the constraint of Equation 18 is satisfied, Equations 10 and 15 ($\varphi=\exp\text{ or }\Phi$) both become exact. For Equation 17 ($\varphi=\rho$), the only error incurred is the one from the probit approximation (Equation 16) per each class. In the synthetic experiment (Figure 1), we see that this induces no significant effect on the quality of the predictive. We justify this observation with the following result:
Theorem 2.1.
Let $\varphi=\rho$. Suppose that the logit means and variances $(\mu_{c},\sigma^{2}_{c})$ lie in some compact set $\mathcal{K}\subset\mathbb{R}\times\mathbb{R}_{>0}$ for each class $c$. Then, assuming the constraint Equation 18 is perfectly satisfied, the KL divergence between the true and approximate predictive Equation 17 is bounded by some constant $M(\mathcal{K})$ only depending on $\mathcal{K}$, not on $C$.
Section C.2 provides a more precise statement, a formula for $M(\mathcal{K})$, and a proof. Importantly, this result states that the approximation error does not increase with the number of classes.
Finally, observe that, when enforcing Equation 18, we obtain exact predictives with Equation 13, which only uses the diagonal of $\mathrm{f}(x)$’s covariance. Enforcing this constraint thus allows one to ignore non-diagonal terms in the logit covariances. In Section 5, we discuss how this can be leveraged for computational gains.
3 Distributional Approximations of the Gaussian Pushforwards
3.1 Dirichlet Matching
While in many applications, predictive probabilities are all that is needed from a model, we set out to build a model that outputs a tractable probability distribution over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$. As previously discussed, such a distribution allows, for instance, closed-form decompositions of aleatoric and epistemic uncertainties.
The formulation of Equation 11 does not yet allow this. However, we will now see that we can obtain good Dirichlet approximations by constructing a tractable approximation of Equation 11 of the form
$\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\quad\mathrm{a}\quad}\mathcal{D}(\Delta^{C-1}),$ (19)
where $\mathcal{D}(\Delta^{C-1})$ is the space of Dirichlet distributions on $\Delta^{C-1}$. The family of Dirichlets presents an ideal choice of distributions over the simplex as, for such distributions, all quantities of interest can be computed analytically (c.f. Appendix G). The map $\mathrm{a}$ is constructed by moment matching: we match the first two moments of the Dirichlet distribution to the moments of the Gaussian pushforward $\bm{\varphi}_{*}\mathrm{f}(x)$.
Conveniently, just as the first moments described in Section 2, we can obtain second moments for $\exp$
$\int_{\mathbb{R}}\exp(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)=\exp(2\mu+2\sigma^{2})$ (20)
and for $\Phi$ (Owen (1980, Eq. 20,010.4))
$\int_{\mathbb{R}}\Phi(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)$ (21)
$\quad=\Phi\left(\frac{\mu}{\sqrt{1+\sigma^{2}}}\right)-2T\left(\frac{\mu}{\sqrt{1+\sigma^{2}}},\frac{1}{\sqrt{1+2\sigma^{2}}}\right),$
where $T$ is Owen’s T function. For $\rho$, we have the following approximation (Daunizeau (2017, Equation 23))
$\int_{\mathbb{R}}\rho(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)$ (22)
$-\frac{\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\left(1-\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\right)}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}.$
This enables us to compute the second moment of the pushforward distribution via
$\mathbb{E}_{\bm{P}\sim\bm{n}_{*}\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{P}^{2}]$ $=\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}\left[\frac{\bm{Q}^{2}}{\left(\sum_{c=1}^{C}Q_{c}\right)^{2}}\right]$ (23)
$\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{Q}^{2}]}{\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}\left[\sum_{c=1}^{C}Q_{c}\right]^{2}}.$
As in Equation 13, the approximation in Equation 23 is exact when $\sum_{c=1}^{C}Q_{c}$ is a constant random variable, as under the constraint of Equation 18.
If we were to directly attempt to infer the parameters of the Dirichlet $\mathrm{h}(x)$ from Equations 13 and 23, we would obtain an overparametrised system of equations with $2C$ equations and $C$ unknowns. Thus, we instead use a classical Dirichlet method of moments (Minka, 2000, Eqs. 19 & 23), which reduces these equations, as we will now describe.
The sum of the Dirichlet parameters $\sum_{c=1}^{C}\gamma_{c}$ may be estimated by
$\frac{\mathbb{E}[P_{c}]-\mathbb{E}[P^{2}_{c}]}{\mathbb{E}[P^{2}_{c}]-\mathbb{E}[P_{c}]^{2}}$ (24)
for any $1\leq c\leq C$. So, a natural estimate of $\sum_{c=1}^{C}\gamma_{c}$ that uses all $P_{c}$ is the geometric mean of Equation 24. Since the mean of the Dirichlet is $\bm{\gamma}/\sum_{c=1}^{C}\gamma_{c}$, we obtain using Equation 24 an expression for the Dirichlet parameters
$\left(\prod_{c=1}^{C}\frac{\mathbb{E}[P_{c}]-\mathbb{E}[P^{2}_{c}]}{\mathbb{E}[P^{2}_{c}]-\mathbb{E}[P_{c}]^{2}}\right)^{1/C}\mathbb{E}[\bm{P}].$ (25)
Hence, using Equations 13 and 23, we obtain the computationally feasible Dirichlet parameters
$\bm{\gamma}:=\left(\prod_{c=1}^{C}\frac{\mathbb{E}[Q_{c}]\cdot S-\mathbb{E}[Q^{2}_{c}]}{\mathbb{E}[Q^{2}_{c}]-\mathbb{E}[Q_{c}]^{2}}\right)^{1/C}\left(\frac{\mathbb{E}[\bm{Q}]}{\sum_{c=1}^{C}\mathbb{E}[Q_{c}]}\right)$ (26)
for $\mathrm{h}(x)$, where $S=\max(\sum_{c=1}^{C}\mathbb{E}[Q_{c}],1)$. Taking the maximum with $1$ in the expression for $S$ is an additional approximation to ensure $\bm{\gamma}\geq\bm{0}$, and is justified when taking $s(x)=1$ for all $x$ in Equation 18. Note that performing such moment matching to obtain the Dirichlet parameters in Equation 26, the mean of this Dirichlet matches precisely the approximate predictive Equation 13. That is, the approximate predictive of the exact distributional model (Equation 11) matches the exact predictive of the approximate distributional model (Equation 19).
3.2 Intermediate Beta Matching
Figure 2: Approximating Gaussian pushforwards $\Phi_{*}\mathrm{f}(x)$ and $\rho_{*}\mathrm{f}(x)$ through normCDF and sigmoid respectively by Beta distributions $\mathrm{p}(\mathrm{f}(x))$ with moment matching. With normCDF, we can match exact moments, whereas with sigmoid, we match approximate moments. Moreover, when $\mathrm{f}(x)$ is the standard Normal $\mathcal{N}(0,1)$ (on the left in blue), the approximation for the pushforward through normCDF is exact (Section D.1).
When $\varphi=\Phi\text{ or }\rho$, noting that their range is $(0,1)$, we can decompose the approximate distributional model (Equation 19) further:
$\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\mathrm{p}}\mathcal{B}((0,1))^{C}\xrightarrow{\mathrm{n}}\mathcal{D}(\Delta^{C-1}),$ (27)
where $\mathcal{B}((0,1))$ is the space of Beta distributions on $(0,1)$. Like the output Dirichlets, we obtain the intermediate Beta distributions by moment matching (Appendix D). Figure 2 indicates that such Beta approximations tend to be of high quality. In Section D.1, we give an information-geometric motivation for this, which shows that the choice of activation $\Phi$ is ideal for approximating the Gaussian pushforwards with Beta distributions.
The quality of such Beta approximations is encouraging for that of the output Dirichlet approximations: observe that taking $s(x)=1$ for all $x$ in Equation 18, $\bm{n}$ becomes the identity (as it normalises by 1), and thus all $\mathrm{n}$ does is combining the $C$ Beta approximations into one Dirichlet.
4 Learning the Constraint
As discussed in Sections 2.3 and 3, the quality of our approximations relies on the enforcement of the constraint $\sum_{c=1}^{C}Q_{c}=s(x)$ for some deterministic function $s\colon\mathcal{X}\to\mathbb{R}$ (Equation 18). In this work, we simply choose $s(x)=1$ for all $x$. We thus make use of the degree of freedom in the logits to obtain exact predictives in Equation 13. In Appendix E, we make the connection with the energy-based model view of classifiers (Grathwohl et al., 2019), which instead makes use of this degree of freedom to turn discriminative models into generative models. Future work could combine both frameworks by learning $s(x)\propto p_{\operatorname{gen}}(x)$.
4.1 Explicit Constraint
Up to this point, we have treated the model $\mathrm{f}$ as a black box. However, to discuss how to impose Equation 18, we need to think about the specifics of how such models are trained and how Gaussian distributions are then obtained.
In linearised Laplace approximations and SNGP, a neural network $\bm{f}\colon\mathcal{X}\to\mathbb{R}^{C}$ is trained for the mean of $\mathrm{f}$, while the covariance is obtained post-hoc. Therefore, to enforce Equation 18 on the training data distribution, we can train for $\sum_{c=1}^{C}\varphi(f_{c}(x_{n}))=1$ for all $1\leq n\leq N$.
This can be achieved, for example, by adding an appropriate regulariser to the loss. Namely,
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ (28)
=~$\quad=\tilde{\mathcal{L}}((x_{n},c_{n})_{n=1}^{N})+\lambda\sum_{n=1}^{N}\left(\sum_{c=1}^{C}\varphi(f_{c}(x_{n}))-1\right)^{2}$
for some tuned regularisation constant $\lambda$, where $\tilde{\mathcal{L}}$ is the original unregularised loss.
In contrast to the previously mentioned methods, HET directly trains not only for the mean $\bm{f}$ but also for the covariance matrix $\bm{\Sigma}$ of $\mathrm{f}$ through the predictives. Thus, in this case, we can regularise with the means of the Gaussian pushforwards through $\varphi$:
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ =~$=\tilde{\mathcal{L}}((x_{n},c_{n})_{n=1}^{N})$ (29)
$+\lambda\sum_{n=1}^{N}\left(\sum_{c=1}^{C}\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}[Q_{c}]-1\right)^{2}.$
In Section 6, we will see that training with such a learning objective does not always outperform vanilla CE training. For $\varphi=\exp$, for instance, we observe that the regulariser can take on extreme values, dominating the data term, slowing down training.
We shall now see that, when $\varphi$ has range $(0,1)$, we can enforce the constraint with the learning objective in a more natural, implicit way.
4.2 Implicit Constraint
In the case of $\varphi=\Phi\text{ or }\rho$, $\sum_{c=1}^{C}\varphi(f_{c}(x_{n}))=1$ can be achieved more advantageously: since the range of $\Phi$ and $\rho$ is $(0,1)$, under the condition $\sum_{c=1}^{C}\varphi(f_{c}(x_{n}))=1$, we can view $\varphi(f_{c}(x_{n}))$ as estimates of $p(c\mid x_{n})$. This motivates the use of class-wise cross-entropy losses (class-wise CE).
For linearised Laplace approximations and SNGP, which solely train a mean $\bm{f}$, the class-wise CE is given by
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ $=-\sum_{n=1}^{N}\sum_{c=1}^{C}\big{(}\delta_{c_{n},c}\log(\varphi(f_{c}(x)))$ (30)
$\quad+(1-\delta_{c_{n},c})\log(1-\varphi(f_{c}(x))\big{)}.$
The labels being one-hot vectors and the binary cross-entropy loss being a (negative) strictly proper scoring rule (c.f. Appendix I) together ensure that Equation 30 enforces the soft constraint $\sum_{c=1}^{C}\varphi\circ\mu_{c}=1$ on the training data, and hence in the training data distribution (Figure 3).
We note that the loss Equation 30 is only used for training and not for inference or evaluation. Indeed, methods such as Laplace approximations and SNGP – which rely on negative log-likelihood losses to obtain an approximate Gaussian distribution – can then use the actual log-likelihood loss of the model, that is, the cross-entropy loss of $\bm{n}\circ\bm{\varphi}\circ\bm{f}$.
Similarly, for HET, we use the class-wise CE with the pushforward means
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ $=-\sum_{n=1}^{N}\sum_{c=1}^{C}\big{(}\delta_{c_{n},c}\log(\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}[Q_{c}])$ (31)
$+(1-\delta_{c_{n},c})\log(1-\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}[Q_{c}])\big{)}.$
Enforcing the constraint in this implicit way is powerful since, in contrast to the regularised CE objective (Equation 28), the class-wise CE (Equation 30) fully aligns with the task. In Section 6 we shall see that training on class-wise CE generally outperforms softmax CE training.
Figure 3: Evolution of the constraint on validation data for normCDF models trained on ImageNet.
5 Computational Gains
The runtime cost of the MC predictive approximation is $\mathcal{O}(S\cdot C+C^{3})$, and the memory cost is $\mathcal{O}(S\cdot C+C^{2})$, where $S$ is the number of samples. The $C^{2}$ term in memory comes from the logit covariance matrix; the $C^{3}$ term in runtime comes from the need to Cholesky-decompose such covariance matrix before sampling. For example, on a ResNet-50, sampling one thousand logit vectors from the logit-space Gaussian means an approximately 7% overhead on the forward pass. In embedded systems with much smaller models, this overhead is even larger. In contrast, the runtime and memory cost of calculating a predictive through our framework is $\mathcal{O}(C)$, which is negligible in comparison.
Indeed, as noted at the end of Section 2.3, the constraint (Equation 18) ensures that only the mean and the diagonal of the logit covariance are needed in order to calculate the predictive exactly. That is, the logit quantities required to compute a predictive scale as $\mathcal{O}(C)$ instead of $\mathcal{O}(C^{2})$. This provides further opportunities for computational savings in the methods used to obtain the Gaussian distributions.
In HET, for example, the logit covariance is trained with both a diagonal term and a low-rank term. Our framework thus allows one to drop the low-rank term without losing any expressivity in the model.
In SNGP, the off-diagonal terms of the logit covariance are ignored (Liu et al., 2023, p. 47). For unconstrained models, this incurs an approximation error as the inter-class covariance structure can encode vital information for inference. When Equation 18 is satisfied, however, we are guaranteed not to lose expressivity for the predictive computation by discarding the off-diagonal covariance terms, theoretically backing the approximation in SNGP.
6 Experiments
We now investigate our two research questions:
  • Do we have to sacrifice performance for sample-free predictives?
  • What are the effects of changing the learning objective?
To answer the first question, we consider fixed (method, activation) pairs and verify that our methods perform on par with the Monte Carlo sampled predictives in Section 6.1.
We train four classes of models: Softmax ($\varphi=\exp$, trained with vanilla CE), Exp ($\varphi=\exp$, trained with the explicit constraint (Section 4.1)), NormCDF ($\varphi=\Phi$, trained with the implicit constraint, (Section 4.2)) and Sigmoid ($\varphi=\rho$, trained with the implicit constraint, (Section 4.2)).
We consider Heteroscedastic Classifiers (HET) (Collier et al., 2021), Spectral-Normalized Gaussian Processes (SNGP) (Liu et al., 2020), and last-layer Laplace approximation methods (Daxberger et al., 2021) as backbones. The resulting 18 (method, activation, predictive) triplets are evaluated on ImageNet-1k (Deng et al., 2009) and CIFAR-10 (Krizhevsky & Hinton, 2009) on five metrics aligning with practical needs from uncertainty estimates (Mucsányi et al., 2023). Further, we test the Out-of-Distribution (OOD) detection capabilities of the models on balanced mixtures of In-Distribution (ID) and OOD samples. For ImageNet, we treat ImageNet-C (Hendrycks & Dietterich, 2019) samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
For the second question, we evaluate our analytic predictives and moment-matched Dirichlets against softmax models with approximate inference tools – Laplace Bridge (Hobbhahn et al., 2022) and Mean Field (Lu et al., 2021) approximations, as well as MC sampling, in Section 6.2.
To provide a fair comparison, we reimplement each method as simple-to-use wrappers around deterministic backbones. For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following Mucsányi et al. (2024). On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. We search for ideal hyperparameters and checkpoints with a ten-step Bayesian Optimization scheme (Shahriari et al., 2015) in Weights & Biases (Biewald, 2020).
The main paper focuses on ImageNet results to highlight the scalability of our framework and only shows proper scoring results for CIFAR-10. For a complete overview of CIFAR-10 results, refer to Appendix K.
In all plots of the main paper and the Appendix, the error bars are calculated over five independent runs with different seeds and show the minimum and maximum performance.
Table 1: Comparison of ECE results for different predictives using a fixed Laplace backbone.
Method Mean Std
Sigmoid Laplace
Analytic 0.0101 0.0006
MC 1000 0.0153 0.0011
MC 100 0.0153 0.0012
MC 10 0.0165 0.0014
NormCDF Laplace
MC 10 0.0112 0.0005
Analytic 0.0114 0.0007
MC 100 0.0115 0.0006
MC 1000 0.0118 0.0007
6.1 Quality of Sample-Free Predictives
In this section, we investigate our first research question: whether there is a price to pay for sample-free predictives and second-order distributions. To this end, we take the two best-performing methods on the Expected Calibration Error (ECE) metric (Naeini et al., 2015), NormCDF and Sigmoid Laplace (c.f. Figure 6) and evaluate their per-predictive performance. As Table 1 shows, the analytic predictive is always on par with the (unbiased) MC estimates on ImageNet while being cheaper. This empirical observation supports our theoretical claim that when the constraint is satisfied, the analytic predictives are exact. Refer to Appendix L for an evaluation of the moment-matched Dirichlet distributions.
6.2 Effects of Changing the Learning Objective
The previous section shows that for an already trained model that satisfies the constraint (Equation 18), our analytic predictives always perform on par with MC sampling while being more efficient. However, our proposed objectives (Sections 4.1 and 4.2) change the training dynamics of models. Therefore, in this section, we showcase the performance of methods equipped with our analytic (i.e., sample-free) predictives and learning objectives against softmax models. For the latter, we use the best-performing predictive and estimator (excluding ours). See Appendices F and G for an overview of predictives and estimators. We employ our methods with the analytic predictives and second-order Dirichlet distributions to demonstrate their competitive nature while being more efficient than MC sampling.
Figure 4: ImageNet log probability proper scoring results. Analytic predictives with implicit ( ) constraints are either on par with or outperform Softmax ( ) on all methods.
Figure 5: CIFAR-10 log probability proper scoring results. Analytic predictives with implicit ( ) constraints consistently outperform Softmax ( ) on all methods.
We first evaluate how calibrated the models are using the log probability scoring rule (Gneiting & Raftery, 2007) and the Expected Calibration Error (ECE) metric (Naeini et al., 2015). Figure 5 shows that on CIFAR-10, the score of our analytic predictives with implicit constraints (Sigmoid, NormCDF) are consistently better than the corresponding Softmax results for all methods. On the large-scale ImageNet dataset, Sigmoid and NormCDF either outperform or are on par with Softmax.
Figure 6: ImageNet ECE results. Analytic predictives with implicit constraints ( ) outperform Softmax ( ) across all methods. Explicit constraints ( ) work best on HET but do not benefit the other two methods. Note the restricted $y$-limits for readability.
The ECE requires the models’ confidence to match their accuracy. Figure 6 shows that on ImageNet, our analytic predictives with implicit constraints (Sigmoid, NormCDF) have a clear advantage over Softmax across all methods. Explicit constraints benefit SNGP and HET but not the post-hoc Laplace method.
Figure 7: ImageNet log probability proper scoring results for the binary correctness prediction task. Analytic predictives with implicit ( ) constraints consistently outperform Softmax ( ).
Next, we turn to the correctness prediction task: whether the models can predict the correctness of their own predictions. We consider correctness estimators $\tilde{c}(x)\in[0,1]$ for inputs $x\in\mathcal{X}$ derived from the predictives. Framed as a binary prediction task, the goal of these estimators is to predict the probability of the predicted class’ correctness. We then measure the binary log probability score of $\tilde{c}(x)$. Figure 7 shows that our analytic predictives with implicit constraints outperform all Softmax predictives across all methods.
Figure 8: Analytic predictives do not sacrifice accuracy. ImageNet validation accuracies. Analytic predictives with implicit ( ) or explicit ( ) constraints either outperform or are on par with Softmax ( ) across all methods.
Analytic predictives do not sacrifice accuracy.
Figure 8 evidences this claim on ImageNet: our analytic predictives either outperform or are on par with Softmax predictives. The most accurate methods are NormCDF and Sigmoid SNGP. These results are in line with the findings of Wightman et al. (2021), who also recommend training with a per-class binary cross-entropy loss using the sigmoid activation function.
Figure 9: ImageNet OOD detection AUROC results for severity level one. Across all methods, the best-performing predictive is analytic with an implicit constraint ( ).
Finally, we consider another binary prediction task, where a general uncertainty estimator $u(x)\in\mathbb{R}$ (derived from predictives or second-order Dirichlet distributions) is tasked to separate ID and OOD samples from a balanced mixture. As the uncertainty estimator can take on any real value, we measure the Area Under the Receiver Operating Characteristic curve (AUROC), which quantifies the separability of ID and OOD samples w.r.t. the uncertainty estimator. As OOD inputs, we consider corrupted ImageNet-C (Hendrycks & Dietterich, 2019) samples. Figure 9 shows that our analytic predictives with implicit constraints (Sigmoid, NormCDF) outperform Softmax across all methods, even though separating ID and OOD samples does not require a fine-grained representation of uncertainty, unlike the ECE or proper scoring rules. Importantly, these analytic predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5). See Appendix M for the other severity levels.
7 Conclusion and Limitations
We developed a framework that allows to obtain predictives and other quantities of interest from logit space Gaussian distributions analytically. Our experimental results suggest the ubiquitous softmax activation should be replaced by normCDF or sigmoid.
We proposed to approximate the Gaussian pushforwards by Dirichlet distributions which cannot encode correlations between classes. Leveraging a more expressive yet tractable family of distributions on the simplex that can achieve this presents an interesting area of future research.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgments
The authors gratefully acknowledge co-funding by the European Union (ERC, ANUBIS, 101123955). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. BM & PH are supported by the DFG through Project HE 7114/6-1 in SPP2298/2. NDC is supported by the Fonds National de la Recherche, Luxembourg, Project 17917615. PH is a member of the Machine Learning Cluster of Excellence, funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. The authors also gratefully acknowledge the German Federal Ministry of Education and Research (BMBF) through the Tübingen AI Center (FKZ: 01IS18039A); and funds from the Ministry of Science, Research and Arts of the State of Baden-Württemberg.
Appendix A Gaussian Integral Derivations
In this appendix, we derive the closed-form formula for the mean of Gaussian pushforwards through exp (Equation 8) and normCDF (Equation 14), as well as the approximations for pushforwards through sigmoid (Equation 16) and softmax.
A.1 Gaussian Exp Integral
By absorbing the exponential into the Gaussian probability density function, we get
[…]
$=\int_{\mathbb{R}}\exp\left(\mu+\frac{\sigma^{2}}{2}\right)\;\mathcal{N}(\mu+\sigma^{2},\sigma^{2})(dy)$
$=\exp\left(\mu+\frac{\sigma^{2}}{2}\right).$
A.2 Gaussian NormCDF Integral
Here, we derive the classical normCDF Gaussian integration formula (Owen, 1980, Eq. 10,010.8).
For $\lambda>0$, $Z\sim\mathcal{N}(0,1)$ and $Y\sim\mathcal{N}(\mu,\sigma^{2})$ with $Y$ and $Z$ independent,
$\int_{\mathbb{R}}\Phi(\lambda y)\;\mathcal{N}(\mu,\sigma^{2})(dy)$ $=\int_{\mathbb{R}}p(Z\leq y/\lambda)\;\mathcal{N}(\mu,\sigma^{2})(dy)$ (33)
$=p\left(Z\leq\frac{Y}{\lambda}\right)$
$=p\left(\frac{Z/\lambda-Y+\mu}{\sqrt{\lambda^{-2}+\sigma^{2}}}\leq\frac{\mu}{\sqrt{\lambda^{-2}+\sigma^{2}}}\right)$
$=\Phi\left(\frac{\mu}{\sqrt{\lambda^{-2}+\sigma^{2}}}\right)$
where we used $\frac{Z/\lambda-X+\mu}{\sqrt{\lambda^{-2}+\sigma^{2}}}\sim\mathcal{N}(0,1)$ for the last equality. Taking $\lambda=1$, this gives the formula for the exact predictive with normCDF.
A.3 Probit Approximation for Gaussian Sigmoid Integral
Taylor expanding $\rho$ to first order about 00,
$\rho(y)$ $=\frac{1}{2}+\frac{1}{4}y+o(y)$ (34)
$\Phi(y)$ $=\frac{1}{2}+\frac{1}{\sqrt{2\pi}}y+o(y).$
Hence matching $\rho$ and $\Phi$ to first order we get the approximation $\rho(y)\approx\Phi\left(\sqrt{\frac{\pi}{8}}y\right)$. So using Equation 33 we derive the probit approximation (Spiegelhalter & Lauritzen, 1990; MacKay, 1992)
$\int_{\mathbb{R}}\rho(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)\stackrel{{\scriptstyle(1)}}{{\approx}}\int_{\mathbb{R}}\Phi\left(\sqrt{\frac{\pi}{8}}y\right)\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{\frac{8}{\pi}+\sigma^{2}}}\right)\stackrel{{\scriptstyle(2)}}{{\approx}}\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right).$ (35)
Note that the approximation (2) is not strictly needed, as $\Phi$ is computationally tractable. However, adding (2) empirically improves the quality of the overall approximation. This may be due to the fact the thicker tails of $\rho$ in the integrand of the left-hand side are better captured by $\rho$ than $\Phi$ on the right-hand side.
A.4 Mean Field Approximation for Gaussian Softmax Integral
For $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, the mean field approximation to the Gaussian softmax integral (Lu et al., 2021) is obtained as follows
$\mathbb{E}_{\bm{Y}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\operatorname{softmax}_{c}\bm{Y}]$ $=\mathbb{E}\left[\left(2-C+\sum_{c^{\prime}\neq c}\rho(Y_{c}-Y_{c^{\prime}})^{-1}\right)^{-1}\right]$ (36)
$\stackrel{{\scriptstyle(1)}}{{\approx}}\left(2-C+\sum_{c^{\prime}\neq c}\mathbb{E}[\rho(Y_{c}-Y_{c^{\prime}})]^{-1}\right)^{-1}$
$\stackrel{{\scriptstyle(2)}}{{\approx}}\left(2-C+\sum_{c^{\prime}\neq c}\mathbb{E}[\rho(Y_{c}-\mu_{c^{\prime}})]^{-1}\right)^{-1}$
$\stackrel{{\scriptstyle(3)}}{{\approx}}\left(2-C+\sum_{c^{\prime}\neq c}\rho\left(\frac{\mu_{c}-\mu_{c^{\prime}}}{\sqrt{1+\frac{\pi}{8}\sigma_{c}^{2}}}\right)^{-1}\right)^{-1}$
$=\operatorname{softmax}_{c}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)$
i.e.,
$\mathbb{E}_{\bm{Y}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\operatorname{\mathbf{softmax}}\bm{Y}]\approx\operatorname{\mathbf{softmax}}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right).$ (37)
(1) is the mean field approximation, and (3) uses the probit approximation Equation 35. (Lu et al., 2021) provides two other variants of this approximation with other choices of approximation (2).
Appendix B Comparison with the Multinomial Probit Model
In this appendix, we show that a model whose output activation is a composition of an element-wise normCDF activation $\bm{\Phi}$ and a normalisation $\bm{n}$ is distinct from the classical multinomial probit model (Daganzo, 1979).
Given a logit $\bm{y}\in\mathbb{R}^{C}$, the multinomial probit model sets
$Z(\bm{y})=\operatorname*{arg\,max}_{1\leq c\leq C}Y_{c}$ (38)
where $Y_{c}=y_{c}+\epsilon_{c}$ and the $\epsilon_{c}$ are i.i.d. standard Normal. So
$p(c\mid y)=p(y_{c}+\epsilon_{c}>y_{c^{\prime}}+\epsilon_{c^{\prime}}\;\forall\>c^{\prime}\neq c)$ (39)
which is generally not analytically tractable. On the other hand, a model that uses normCDF and normalisation as output activation yields
$p(c\mid y)=\frac{p(y_{c}+\epsilon_{c}>0)}{\sum_{c^{\prime}=1}^{C}p(y_{c^{\prime}}+\epsilon_{c^{\prime}}>0)}=\frac{\Phi(y_{c})}{\sum_{c^{\prime}=1}^{C}\Phi(y_{c}^{\prime})}.$ (40)
In the case $C=2$, the multinomial probit model Equation 39 outputs closed form probabilities. This allows us to construct an explicit counterexample to the equivalence of the two models Equation 39 and Equation 40:
$p(y_{1}+\epsilon_{1}>y_{2}+\epsilon_{2})$ $=p\left(\frac{y_{1}-y_{2}}{2}+\frac{\epsilon_{1}-\epsilon_{2}}{2}>0\right)=\Phi\left(\frac{y_{1}-y_{2}}{2}\right)\neq\frac{\Phi(y_{1})}{\Phi(y_{1})+\Phi(y_{2})}.$ (41)
Appendix C Analysis of the Predictive Approximations’ Quality
In this appendix, we provide theoretical analyses of the quality of various predictive approximations. This complements the empirical analyses taken, for instance, in the synthetic experiment (Figure 1) or in (Daunizeau, 2017).
Due to its information-theoretic interpretation, a natural metric to equip the probability simplex $\Delta^{C-1}$ with is the KL divergence
$\operatorname{D_{KL}}(\bm{p},\bm{q})=\sum_{c=1}^{C}p_{c}(\log p_{c}-\log q_{c})$ (42)
which is well defined if $\bm{p}$, $\bm{q}$ lie in the interior of the simplex ($p_{i},q_{i}\neq 0,1$ for all $i$). So, for a predictive approximation, $\bm{\hat{p}}$, we would like to analyse $\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$, where $\bm{p}:=\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}]$ is the true predictive, $\bm{\mu}$ and $\bm{\Sigma}$ are some logit space mean and covariance and $\bm{a}$ is an output activation (e.g. $\bm{a}=\bm{n}\circ\bm{\varphi}$).
C.1 Analysis of Monte Carlo Approximations
An $N$ sample Monte Carlo estimate is defined as
$\bm{\hat{P}}^{S}:=\frac{1}{S}\sum_{s=1}^{S}\bm{\hat{P}}^{(s)}$ (43)
where the $\bm{\hat{P}}^{(s)}$ are i.i.d. $\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})$. The computational cost of MC integration is $\mathcal{O}(S\cdot C)$. This becomes prohibitive for large $S$ and $C$. Thus, for a fair assessment of the quality of such an estimate in terms of the number of classes, one should consider MC estimates $\bm{\hat{P}}^{\lceil S/C\rceil}$.
We now give an informal theoretical argument for the linear growth of the KL divergence between $\bm{p}$ and $\bm{\hat{P}}^{\lceil S/C\rceil}$ in terms of $C$, under the distributional conditions of the synthetic experiment (Figure 1).
Taylor expanding Equation 42 about $\bm{p}$ to second order we obtain
$\operatorname{D_{KL}}(\bm{p},\bm{q})$ $=\sum_{c=1}^{C}p_{c}(\log p_{c}-\log q_{c})$ (44)
$=\sum_{c=1}^{C}p_{c}\sum_{k=1}^{\infty}\frac{(p_{c}-q_{c})^{k}}{kp_{c}^{k}}$
$=\underbrace{\sum_{c=1}^{C}p_{c}}_{=1}-\underbrace{\sum_{c=1}^{C}q_{c}}_{=1}+\sum_{c=1}^{C}\sum_{k=2}^{\infty}\frac{(p_{c}-q_{c})^{k}}{kp_{c}^{k-1}}$
$\stackrel{{\scriptstyle(1)}}{{\approx}}\sum_{c=1}^{C}\frac{(p_{c}-q_{c})^{2}}{2p_{c}}$
$\stackrel{{\scriptstyle(2)}}{{\approx}}\frac{C}{2}\|\bm{p}-\bm{q}\|_{2}^{2}$
where approximation (1) assumes $\|\bm{p}-\bm{q}\|_{2}$ is small, and (2) assumes $p_{c}\approx 1/C$.
In the synthetic experiment (Figure 1), the logit class-wise means $\mu_{c}$ and variances $\sigma_{c}^{2}$ are sampled in an i.i.d. way. Let $\bm{Y}\sim N(\bm{\mu},\bm{\Sigma})$ with $\bm{\Sigma}:=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, $\bm{Q}=\bm{\varphi}(\bm{Y})$ the unnormalised ‘probabilities’ and $\bm{P}:=\bm{Q}/\sum_{c=1}^{C}Q_{c}$ the probabilities. We have
$\mathrm{Var}[\bm{P}]=\mathbb{E}\left[\frac{\bm{Q}^{2}}{\left(\sum_{c=1}^{C}Q_{c}\right)^{2}}\right]-\mathbb{E}\left[\frac{\bm{Q}}{\sum_{c=1}^{C}Q_{c}}\right]^{2}\approx\frac{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}{\left(\sum_{c=1}^{C}\mathbb{E}[Q_{c}]\right)^{2}}=\frac{\mathrm{Var}[\bm{Q}]}{C^{2}\mathbb{E}[Q_{1}]^{2}}$ (45)
where all operations are taken element-wise. Thus
$\mathbb{E}\left[\|\bm{p}-\bm{P}\|^{2}_{2}\right]=\sum_{c=1}^{C}\mathrm{Var}[P_{c}]\approx\sum_{c=1}^{C}\frac{\mathrm{Var}[Q_{c}]}{C^{2}\mathbb{E}[Q_{1}]^{2}}=\frac{\mathrm{Var}[Q_{1}]}{C\mathbb{E}[Q_{1}]^{2}}.$ (46)
Now the MC samples $\bm{\hat{P}}^{(s)}$ are i.i.d. copies of $\bm{P}$. So we have
$\mathbb{E}\big{[}\|\bm{p}-\bm{\hat{P}}^{\lceil S/C\rceil}\|^{2}_{2}\big{]}=\frac{1}{\lceil S/C\rceil}\mathbb{E}\left[\|\bm{p}-\bm{P}\|^{2}_{2}\right]\approx\frac{\mathrm{Var}[Q_{1}]}{S\mathbb{E}[Q_{1}]^{2}}.$ (47)
Plugging this into Equation 44 we get
$\mathbb{E}[\operatorname{D_{KL}}(\bm{p},\bm{\hat{P}}^{\lceil S/C\rceil})]\approx\frac{\mathrm{Var}[Q_{1}]}{2\mathbb{E}[Q_{1}]^{2}}\cdot\frac{C}{S}$ (48)
which grows linearly with the number of classes, as observed in Figure 1.
C.2 Analysis of the Probit Approximation for Sigmoid Models
Recall the probit approximation to the Gaussian sigmoid integral (Section A.3)
$q:=\int_{\mathbb{R}}\rho(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)=:\hat{q}.$ (49)
In the multiclass setting, for $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, define
$\bm{q}:=\mathbb{E}$ [${}_{\bm{Q}\sim\bm{\rho}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{Q}]$ (50)
$\bm{\hat{q}}:=\bm{\rho}$ $\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)$
$\bm{p}:=\mathbb{E}_{\bm{P}\sim\bm{n}_{*}\bm{\rho}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}$ $[\bm{P}]=\mathbb{E}_{\bm{Q}\sim\bm{\rho}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}\left[\frac{\bm{Q}}{\sum_{c=1}^{C}Q_{c}}\right]$
$\bm{\hat{p}}:=\frac{\bm{\hat{q}}}{\sum_{c=1}^{C}\hat{q}_{c}}$ $=\frac{\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)}{\sum_{c=1}^{C}\rho\left(\frac{\mu_{c}}{\sqrt{1+\frac{\pi}{8}\sigma^{2}_{c}}}\right)}$
We assume that, e.g. through training, the constraint Equation 18 is perfectly satisfied, such that $\bm{p}=\bm{q}/\sum_{c=1}^{C}q_{c}$. Then we obtain
Theorem C.1.
Let $\mathcal{K}\subset\mathbb{R}\times\mathbb{R}_{>0}$ compact. Using the compactness of $\mathcal{K}$ define $\delta(\mathcal{K}):=\max_{(\mu,\sigma^{2})\in\mathcal{K}}|q-\hat{q}|$ and $u(\mathcal{K}):=\min_{(\mu,\sigma^{2})\in\mathcal{K}}q>0$, where $q$ and $\hat{q}$ are as in Equation 49. Then if $(\mu_{c},\sigma_{c}^{2})\in\mathcal{K}$ for all $c$ we have
$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})\leq\log\left(\frac{u+\delta}{u-\delta}\right).$ (51)
Remark C.2.
The bound Equation 51 is of practical value as
  1. 1.
    it is independent of $C$,
  2. 2.
    it tends to 0 as $\delta$ tends to 0.
In other words, given knowledge of the worst case error in the approximation Equation 49 on the compact set $\mathcal{K}$, we can bound the KL divergence in terms of that error independently of the number of classes. Due to the simplicity of our assumptions, the bound remains quite raw and could be strengthened with further distributional assumptions on the means and variances.
Proof.
$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$ $=\sum_{c=1}^{C}p_{c}(\log p_{c}-\log\hat{p}_{c})$ (52)
$=\sum_{c=1}^{C}\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\left(\log\left(\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)-\log\left(\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}}\right)\right)$
$=\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(\frac{\hat{q}_{c}}{q_{c}}\right)+\log\left(\frac{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)\right)$
$\leq\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(\frac{q_{c}-\delta}{q_{c}}\right)+\log\left(\frac{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}+C\delta}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)\right)$
$=\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(1-\frac{\delta}{q_{c}}\right)+\log\left(1+\frac{C\delta}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)\right)$
$\leq\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(1-\frac{\delta}{u}\right)+\log\left(1+\frac{C\delta}{Cu}\right)\right)$
$=\log\left(\frac{u+\delta}{u-\delta}\right).$
C.3 Analysis of the Mean Field Approximation for Softmax Models
Recall the mean field approximation to the Gaussian softmax integral (Section A.4)
$\bm{p}:=\int_{\mathbb{R}^{C}}\operatorname{\mathbf{softmax}}(\bm{x})\;\mathcal{N}(\bm{\mu},\bm{\Sigma})(d\bm{x})\approx\operatorname{\mathbf{softmax}}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)=:\bm{\hat{p}}.$ (53)
This approximation is obtained through three subsequent approximations, that is, (1), (2), and (3) in Equation 36. Approximation (2) is not actually necessary, as it can be replaced by an exact computation, as shown in (Lu et al., 2021). However, this empirically does not appear to improve the quality of the overall approximation. We focus our attention on the main approximation (1), namely
$\mathbb{E}\left[\left(2-C+\sum_{c^{\prime}\neq c}\rho(Y_{c}-Y_{c^{\prime}})^{-1}\right)^{-1}\right]\stackrel{{\scriptstyle(1)}}{{\approx}}\left(2-C+\sum_{c^{\prime}\neq c}\mathbb{E}[\rho(Y_{c}-Y_{c^{\prime}})]^{-1}\right)^{-1}.$ (54)
This is exact if $\rho(Y_{c}-Y_{c^{\prime}})$ is a constant random variable for all $Y_{c}$ and $Y_{c^{\prime}}$. For $C>2$, this would imply that $Y_{c}$ is distributed according to a Dirac delta. However, this does not allow for any second-order distribution over the logits. We thus see that we cannot train for exactness in the mean field approximation as we do for Equation 13. This is a fundamental limitation of this mean field approximation.
Appendix D Moment Matching Beta distributions
As noted in Section 3.2, when $\varphi=\Phi\text{ or }\rho$, we can construct a mapping
$\mathrm{p}\colon\mathcal{G}(\mathbb{R}^{C})\to\mathcal{B}((0,1))^{C}$ (55)
by moment matching. Specifically, the parameters $\bm{\alpha},\bm{\beta}\in\mathbb{R}_{>0}^{C}$ that match the moments of $\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}$ for some $\mathrm{f}\in\mathcal{G}(\mathbb{R}^{C})$ must satisfy
$\mathbb{E}[\bm{Q}]$ $=\frac{\bm{\alpha}}{\bm{\alpha}+\bm{\beta}},$ (56)
$\mathbb{E}[\bm{Q}^{2}]$ $=\frac{\bm{\alpha}(\bm{\alpha}+1)}{(\bm{\alpha}+\bm{\beta})(\bm{\alpha}+\bm{\beta}+1)},$
where all vector operations are element-wise. Multiplying out the denominators on the right-hand side of the equations of Equation 56, we obtain a system of two linear equations with two unknowns (for each $c$), which can be solved uniquely, yielding
$\bm{\alpha}$ $:=\frac{\mathbb{E}[\bm{Q}]-\mathbb{E}[\bm{Q}^{2}]}{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}\mathbb{E}[\bm{Q}],$ (57)
$\bm{\beta}$ $:=\frac{\mathbb{E}[\bm{Q}]-\mathbb{E}[\bm{Q}^{2}]}{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}\left(1-\mathbb{E}[\bm{Q}]\right).$
which give us the parameters of the Beta distributions $\mathrm{p}(\mathrm{f})$.
D.1 Information Geometric Interpretation of the Pushforward through NormCDF
Figure 10: Illustration of the statistical manifolds of the pushforward Gaussian distributions $\mathcal{G}(\mathbb{R})$ through the normCDF and the sigmoid respectively compared to the statistical manifold of Beta distributions $\mathcal{B}((0,1))$. NormCDF, unlike the sigmoid, makes the manifolds intersect at the point $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$.
Here, we argue that the normCDF activation is an ideal choice for approximating Gaussian pushforwards with Beta distributions by interpreting Figure 10. This extends the work from (MacKay, 1998), as it shows how one can make sense of the ‘right’ basis for performing Laplace approximations in the classification setting.
$\Phi_{*}\mathcal{G}(\mathbb{R})$, $\rho_{*}\mathcal{G}(\mathbb{R})$ the space of pushforwards of Gaussian distributions by normCDF and sigmoid respectively, and $\mathcal{B}((0,1))$, the space of Beta distributions, are statistical manifolds naturally equipped with Riemannian metrics, that is their respective Fisher information metrics. We would like to visualise these manifolds. However, two difficulties arise.
The first difficulty is that, while these manifolds all lie in the infinite-dimensional vector space of signed measures on the open unit interval $\mathcal{M}((0,1))$, there is no subspace $\mathbb{V}\subset\mathcal{M}((0,1))$ which is 3-dimensional ($\mathbb{V}\cong\mathbb{R}^{3}$) and contains any two of these statistical manifolds ($\Phi_{*}\mathcal{G}(\mathbb{R}),\mathcal{B}((0,1))\subset\mathbb{V}$ or $\rho_{*}\mathcal{G}(\mathbb{R}),\mathcal{B}((0,1))\subset\mathbb{V}$). We will work around this by building distinct isometric embeddings $\Phi_{*}\mathcal{G}(\mathbb{R})\hookrightarrow\mathbb{V}$, $\rho_{*}\mathcal{G}(\mathbb{R})\hookrightarrow\mathbb{V}$ and $\mathcal{B}((0,1))\hookrightarrow\mathbb{V}$ for some 3-dimensional vector space $\mathbb{V}$. This means that while the shape of the manifold illustrations is meaningful, the positioning of a manifold with respect to another is not, apart from some design choices that we describe below.
The second difficulty is that some–if not all–of these manifolds cannot be embedded isometrically into Euclidean space. As a workaround, we instead embed them into the 3-dimensional Minkowski space $\mathbb{R}^{2,1}$, that is $\mathbb{R}^{3}$ equipped with the pseudo-Riemannian metric $dx_{1}^{2}+dx_{2}^{2}-dx_{3}^{2}$.
The key observation is that $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$ are isometric to $\mathcal{G}(\mathbb{R})$. This is because $\Phi$ and $\rho$ are diffeomorphisms, so in particular sufficient statistics, and the Fisher information metric is invariant under sufficient statistics (Ay et al., 2017, Section 5.1.3). Visually, this means that $\varphi_{*}\mathcal{G}(\mathbb{R})$ has the same shape irrespectively of the diffeomorphism activation function $\varphi$. One can thus observe that, given that $\mathcal{B}((0,1))$ is not isometric to $\mathcal{G}(\mathbb{R})$, there exists no activation $\varphi$ such that $\varphi_{*}\mathcal{G}(\mathbb{R})=\mathcal{B}((0,1))$. To design an activation $\varphi$ that maps Gaussians to Betas, the best one can hope to do is to map one specific Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ to a specific Beta distribution $B(\alpha,\beta)$. This is done with the map $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}$ where $\Phi_{\mu,\sigma^{2}}$ and $F_{\alpha,\beta}$ are the cumulative distribution functions of $\mathcal{N}(\mu,\sigma^{2})$ and $B(\alpha,\beta)$ respectively. Taking $\mu=0$, $\sigma^{2}=1$, $\alpha=\beta=1$ we get $\Phi_{0,1}=\Phi$ and $F_{1,1}=\operatorname{id}_{(0,1)}$, yielding $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}=\Phi$.
Now $\mathcal{G}(\mathbb{R})$, and hence $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$, is isometric to the hyperbolic plane Ay et al. (2017, Example 3.1). We can embed this isometrically into Minkowski space with the classical hyperboloid model of the hyperbolic plane (Reynolds, 1993),
$\mathcal{G}(\mathbb{R})\hookrightarrow\mathbb{R}^{2,1}.$ (58)
For $\mathcal{B}((0,1))$, we use the isometric embedding from Le Brigant et al. (2021, Proposition 2):
$\mathcal{B}((0,1))$ $\hookrightarrow\mathbb{R}^{2,1}$ (59)
$(\alpha,\beta)$ $\mapsto(\eta(\alpha),\eta(\beta),\eta(\alpha+\beta))$
[…]
Finally, we choose our embedding Equation 58 for $\Phi_{*}\mathcal{G}(\mathbb{R})$ such that it intersects the embedding Equation 59 at a point, to highlight that the statistical manifolds $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ intersect at a point in the infinite-dimensional ambient space $\mathcal{M}((0,1))$, while $\rho_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ do not.
Moreover, since $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$, our moment matching approximation Equation 56 is exact when $\mathrm{f}(x)$ is the standard Normal distribution.
Appendix E Comparison with the Energy Model View of Classifiers
Grathwohl et al. (2019) propose to view classifiers of the form Equation 2 as energy-based models by
$p(c\mid x)$ $=\frac{\exp(f_{c}(x))}{\sum_{c^{\prime}=1}^{C}\exp(f_{c^{\prime}}(x))},$ (60)
$p(x,c)$ $=\frac{\exp(f_{c}(x))}{Z},$
$p(x)$ $=\frac{\sum_{c=1}^{C}\exp(f_{c}(x))}{Z}$
where $Z=\int_{\mathcal{X}}\sum_{c=1}^{C}\exp(f_{c}(x))\;dx$. Note the consistency
$p(c\mid x)=\frac{p(x,c)}{p(x)}.$ (61)
Observe that the choice of $p(x,c)$ and $p(x)$ in Equation 60 is not canonical: we could have multiplied $p(x,c)$ and $p(x)$ by $\frac{h(x)}{Z^{\prime}}$ for any non-vanishing function $h$ and appropriate normalisation constant $Z^{\prime}$, and we would still obtain the consistency Equation 61. Grathwohl et al. (2019) propose to use the degree of freedom in the logits to train for Equation 60, turning discriminative models into generative models.
Note that $\exp$ in Equation 60 can be replaced by any positive activation $\varphi$. In our framework, we train for $\sum_{c=1}^{C}\varphi(f_{c}(x))=s(x)$. More specifically, in the present work, we train for $\sum_{c=1}^{C}\varphi(f_{c}(x))=1$ (Section 4), making use of the degree of freedom in the logits to obtain exact predictives in Equation 13. In the formalism Equation 60, this means we are setting $x$ to be uniformly distributed. An interesting avenue of future research would be to combine our framework with that of Grathwohl et al. (2019), by e.g. training for $\sum_{c=1}^{C}\varphi(f_{c}(x))\propto p(x)$, and extending the framework of Grathwohl et al. (2019) to second-order distributions.
Appendix F List of Predictive and Dirichlet Parameters Formulas
F.1 Predictive Formulas
We gather formulas for all predictive estimators $\bm{\hat{p}}$ of the true predictive $\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}]$, $\bm{\Sigma}=\bm{\operatorname{\mathbf{diag}}}(\bm{\sigma}^{2})$, used in our experiments (Section 6).
F.1.1 Softmax
Monte Carlo
One can Monte Carlo estimate the true predictive as follows:
$\bm{\hat{p}}:=\frac{1}{S}\sum_{s=1}^{S}\bm{\hat{p}}^{(s)}$ (62)
where the $\bm{\hat{p}}^{(s)}$ are sampled i.i.d. from $\mathcal{N}(\bm{\mu},\bm{\Sigma})$.
Mean Field (Section A.4)
The Mean Field predictive (Lu et al., 2021) uses the following approximation for the true predictive:
$\bm{\hat{p}}:=\bm{\operatorname{\mathbf{softmax}}}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right).$ (63)
Laplace Bridge
The Laplace Bridge predictive (Hobbhahn et al., 2022) approximates the true predictive as follows:
$\bm{\hat{p}}:=\frac{\frac{1}{\bm{\tilde{\sigma}}^{2}}\left(1-\frac{2}{C}+\frac{e^{\bm{\tilde{\mu}}}}{C^{2}}\sum_{c=1}^{C}e^{-\tilde{\mu}_{c}}\right)}{\sum_{c=1}^{C}\frac{1}{\tilde{\sigma}^{2}_{c}}\left(1-\frac{2}{C}+\frac{e^{\bm{\tilde{\mu}}}}{C^{2}}\sum_{c^{\prime}=1}^{C}e^{-\tilde{\mu}_{c^{\prime}}}\right)}$ (64)
where
$\bm{\tilde{\mu}}^{2}:=\sqrt{\frac{\sqrt{C/2}}{\sum_{c=1}^{C}\sigma^{2}_{c}}}\bm{\mu},\;\bm{\tilde{\sigma}^{2}}:=\frac{\sqrt{C/2}}{\sum_{c=1}^{C}\sigma^{2}_{c}}\bm{\sigma}^{2}.$ (65)
F.1.2 Exp
Our analytic predictive for the exp activation function (Section 2.1) is given by
$\bm{\hat{p}}:=\frac{\bm{\exp}\left(\bm{\mu}+\frac{\bm{\sigma}^{2}}{2}\right)}{\sum_{c=1}^{C}\exp\left(\mu_{c}+\frac{\sigma^{2}_{c}}{2}\right)}.$ (66)
F.1.3 NormCDF
The analytic predictive for the normCDF activation function (Section 2.2) is computed as
$\bm{\hat{p}}:=\frac{\bm{\Phi}\left(\frac{\bm{\mu}}{\sqrt{1+\bm{\sigma}^{2}}}\right)}{\sum_{c=1}^{C}\Phi\left(\frac{\mu_{c}}{\sqrt{1+\sigma^{2}_{c}}}\right)}.$ (67)
F.1.4 Sigmoid
For the sigmoid activation function (Section 2.2), the analytic predictive can be computed as
$\bm{\hat{p}}:=\frac{\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)}{\sum_{c=1}^{C}\rho\left(\frac{\mu_{c}}{\sqrt{1+\frac{\pi}{8}\sigma^{2}_{c}}}\right)}.$ (68)
F.2 Dirichlet Parameters Formulas
We now gather the formulas of the parameters $\bm{\gamma}$ for the Dirichlet approximations to the Gaussian pushforwards.
F.2.1 Softmax
The Laplace Bridge method (Hobbhahn et al., 2022) uses the following Dirichlet parameters:
$\bm{\gamma}:=\frac{1}{\bm{\tilde{\sigma}}^{2}}\left(1-\frac{2}{C}+\frac{e^{\bm{\tilde{\mu}}}}{C^{2}}\sum_{c=1}^{C}e^{-\tilde{\mu}_{c}}\right)$ (69)
where
$\bm{\tilde{\mu}}^{2}:=\sqrt{\frac{\sqrt{C/2}}{\sum_{c=1}^{C}\sigma^{2}_{c}}}\bm{\mu},\;\bm{\tilde{\sigma}^{2}}:=\frac{\sqrt{C/2}}{\sum_{c=1}^{C}\sigma^{2}_{c}}\bm{\sigma}^{2}.$ (70)
F.2.2 Exp
Exp uses the analytic parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
$\bm{\gamma}:=\left(\prod_{c=1}^{C}\frac{a_{c}\cdot\max(\sum_{c^{\prime}=1}^{C}a_{c^{\prime}},1)-b_{c}}{b_{c}-a_{c}^{2}}\right)^{1/C}\frac{\bm{a}}{\sum_{c=1}^{C}a_{c}}$ (71)
where
$\bm{a}=\exp\left(\bm{\mu}+\frac{\bm{\sigma}^{2}}{2}\right),\;\bm{b}=\exp(2\bm{\mu}+2\bm{\sigma}^{2}).$ (72)
F.2.3 NormCDF
NormCDF uses the analytic parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
$\bm{\gamma}:=\left(\prod_{c=1}^{C}\frac{a_{c}\cdot\max(\sum_{c^{\prime}=1}^{C}a_{c^{\prime}},1)-b_{c}}{b_{c}-a_{c}^{2}}\right)^{1/C}\frac{\bm{a}}{\sum_{c=1}^{C}a_{c}}$ (73)
where
$\bm{a}=\bm{\Phi}\left(\frac{\bm{\mu}}{\sqrt{1+\bm{\sigma}^{2}}}\right),\;\bm{b}=\bm{\Phi}\left(\frac{\bm{\mu}}{\sqrt{1+\bm{\sigma}^{2}}}\right)-2\bm{T}\left(\frac{\bm{\mu}}{\sqrt{1+\bm{\sigma}^{2}}},\frac{1}{\sqrt{1+2\bm{\sigma}^{2}}}\right).$ (74)
F.2.4 Sigmoid
Sigmoid uses the analytic parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
$\bm{\gamma}:=\left(\prod_{c=1}^{C}\frac{a_{c}\cdot\max(\sum_{c^{\prime}=1}^{C}a_{c^{\prime}},1)-b_{c}}{b_{c}-a_{c}^{2}}\right)^{1/C}\frac{\bm{a}}{\sum_{c=1}^{C}a_{c}}$ (75)
where
$\bm{a}=\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right),\;\bm{b}=\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)-\frac{1}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)\left(1-\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)\right).$ (76)
Appendix G List of Uncertainty Estimators
In this section, we list the uncertainty estimators used in our experiments (Section 6).
G.1 Predictive
Given a predictive $\bm{p}$, we consider two uncertainty estimators.
Maximum Probability
$\operatorname*{arg\,max}_{c\in\{1,\dots C\}}p_{c}.$ (77)
Entropy
$-\sum_{c=1}^{C}p_{c}\log p_{c}.$ (78)
G.2 Monte Carlo
Given $S$ Monte Carlo samples $\bm{\hat{p}}^{(1)},\dots,\bm{\hat{p}}^{(S)}$ with mean $\bm{\hat{p}}$, one can calculate a predictive as their average and derive the estimators in Section G.1. However, Monte Carlo samples allow one to calculate two additional estimators detailed below.
Expected Entropy
$-\frac{1}{S}\sum_{s=1}^{S}\sum_{c=1}^{C}\hat{p}^{(s)}_{c}\log\hat{p}^{(s)}_{c}.$ (79)
Mutual Information/Jensen-Shannon Divergence
$-\sum_{c=1}^{C}{\hat{p}}_{c}\log{\hat{p}}_{c}+\frac{1}{S}\sum_{s=1}^{S}\sum_{c=1}^{C}\hat{p}^{(s)}_{c}\log\hat{p}^{(s)}_{c}.$ (80)
G.3 Dirichlet
Given a second-order Dirichlet distribution with parameters $\bm{\gamma}$, one can obtain the expected entropy and mutual information estimators without the need for Monte Carlo samples.
Expected Entropy
$-\sum_{c=1}^{C}\frac{\gamma_{c}}{\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}}\left(\psi(\gamma_{c}+1)-\psi\left(\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}+1\right)\right)$ (81)
where $\psi$ is the digamma function.
Mutual Information
$\sum_{c=1}^{C}\frac{\gamma_{c}}{\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}}\left(-\log\gamma_{c}+\log\left(\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}\right)+\psi(\gamma_{c}+1)-\psi\left(\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}+1\right)\right).$ (82)
Appendix H Experimental Setup
This section describes our experimental setup in detail.
We have two main research questions:
  • What are the effects of changing the learning objective?
  • Do we have to sacrifice performance for sample-free predictives?
To answer the first question, we evaluate our analytic predictives (Sigmoid, NormCDF, Exp) and moment-matched Dirichlet distributions against softmax models equipped with approximate inference tools (Laplace Bridge (Hobbhahn et al., 2022), Mean Field (Lu et al., 2021), Monte Carlo sampling). We consider Heteroscedastic Classifiers (HET) (Collier et al., 2021), Spectral-Normalized Gaussian Processes (SNGP) (Liu et al., 2020), and last-layer Laplace approximation methods (Daxberger et al., 2021) as backbones (see Appendix J for details).
The resulting 18 (method, activation, predictive) triplets are evaluated on ImageNet-1k (Deng et al., 2009) and CIFAR-10 (Krizhevsky & Hinton, 2009) on five metrics aligning with practical needs from uncertainty estimates (Mucsányi et al., 2023):
  1. 1.
    Log probability proper scoring rule for the predictive,
  2. 2.
    Expected calibration error of the predictive’s maximum-probability confidence,
  3. 3.
    Binary log probability proper scoring rule for the correctness prediction task,
  4. 4.
    Accuracy of the predictive’s argmax,
  5. 5.
    AUROC for the out-of-distribution (OOD) detection task.
See Appendix I for details.
For ImageNet, we treat ImageNet-C (Hendrycks & Dietterich, 2019) samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
For the second question, we consider fixed (method, activation) pairs and test whether our methods perform on par with the Monte Carlo sampled predictives.
To provide a fair comparison, we reimplement each method as simple-to-use wrappers around deterministic backbones.
For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following Mucsányi et al. (2024). We train with the LAMB optimizer (You et al., 2019) using a batch size of $128$ and gradient accumulation across $16$ batches, resulting in an effective batch size of $2048$, following Tran et al. (2022). We further use a cosine learning rate schedule with a single warmup epoch using a warmup learning rate of $0.0001$. The learning rate is treated as a hyperparameter and selected from the interval $[0.0005,0.05]$ based on the validation performance. The weight decay is selected from the set $\{0.01,0.02\}$. During training, we keep track of the best-performing checkpoint on the validation set and load it before testing. We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme (Shahriari et al., 2015) in Weights & Biases (Biewald, 2020) based on the negative log-likelihood.
On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. The only exceptions are the SNGP models that are trained for 125 epochs (Liu et al., 2020). We train with Momentum SGD using a batch size of $128$ and no gradient accumulation. Similarly to ImageNet, we use a cosine learning rate schedule but with five warmup epochs and warmup learning rate $1\mathrm{e}{-5}$. The learning rate is also treated as a hyperparameter on CIFAR-10. We use the interval $[0.05,1]$ for Sigmoid and NormCDF, and $[0.01,0.15]$ for Softmax and Exp. The optimal learning rates for Sigmoid and NormCDF are generally larger, as the class-wise binary cross-entropies are averaged instead of summed. The weight decay is selected from the interval $[1\mathrm{e}{-6},1\mathrm{e}{-4}]$. Similarly to ImageNet, we use the best-performing checkpoint in the tests and use a ten-step Bayesian Optimization scheme to select performant hyperparameters.
Appendix I Benchmark Metrics
Our experiments use five tasks/metrics:
  1. 1.
    Log probability proper scoring rule for the predictive,
  2. 2.
    Expected calibration error of the predictive’s maximum-probability confidence,
  3. 3.
    Binary log probability proper scoring rule for the correctness prediction task,
  4. 4.
    Accuracy of the predictive’s argmax,
  5. 5.
    AUROC for the out-of-distribution detection task.
Below, we describe these metrics and their respective tasks.
I.1 Log Probability Proper Scoring Rule for the Predictive
First, we briefly discuss proper and strictly proper scoring rules over general probability measures based on (Mucsányi et al., 2023).
Consider a function $S\colon\mathcal{Q}\times\mathcal{Y}\to\mathbb{R}$ where $\mathcal{Q}$ is a family of probability distributions over the space $\mathcal{Y}$, called the label space.
$S$ is called a proper scoring rule if and only if
$\max_{q\in\mathcal{Q}}\mathbb{E}_{Y\sim p}S(q,Y)=\mathbb{E}_{Y\sim p}S(p,Y),$ (83)
i.e., $p$ is one of the maximisers of $S$ in $q$ in expectation. $S$ is further strictly proper if $\operatorname*{arg\,max}_{q\in\mathcal{Q}}\mathbb{E}_{Y\sim p}S(q,Y)=p$ is the unique maximiser of $S$ in $q$ in expectation.
The log probability scoring rule for categorical distributions is defined as
$S(q,c)=\sum_{c^{\prime}=1}^{C}\delta_{c,c^{\prime}}\log q_{c^{\prime}}(x)=\log q_{c}(x),$ (84)
where $c\in\{1,\dotsc,C\}$ is the true class and $\delta$ is the Kronecker delta. $S$ defined this way is a strictly proper scoring rule, i.e. $\mathbb{E}_{Y\sim p}S(q,Y)$ is maximal if and only if
$q(Y=c\mid x)=p(Y=c\mid x)\;\forall c\in\{1,\dotsc,C\}.$ (85)
The score above is equivalent to the negative cross-entropy loss.
I.2 Expected Calibration Error
To set up the required quantities for the Expected Calibration Error (ECE) metric (Naeini et al., 2015), we follow the steps below, based on (Mucsányi et al., 2023).
  1. 1.
    Train a neural network on the training dataset.
  2. 2.
    Create predictions and confidence estimates on the test data.
  3. 3.
    Group the predictions into $M$ bins based on the confidences estimates. Define bin $B_{m}$ to be the set of all indices $n$ of predictions $(\hat{y}_{n},\tilde{c}_{n})$ for which
    $\tilde{c}_{n}\in\left(\frac{m-1}{M},\frac{m}{M}\right].$ (86)
The Expected Calibration Error (ECE) metric (Naeini et al., 2015) is then defined as
$\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\right|$ (87)
where
$\mathrm{acc}(B_{m})$ $=\frac{1}{|B_{m}|}\sum_{n\in B_{m}}1\left(\hat{y}_{n}=c_{n}\right),$ (88)
$\mathrm{conf}(B_{m})$ $=\frac{1}{|B_{m}|}\sum_{n\in B_{m}}\max_{c\in\{1,\dotsc,C\}}f_{c}(x_{n}).$ (89)
Intuitively, the ECE is high when the model’s per-bin confidences match its accuracy on the bin. We use $M=15$ bins in this paper.
I.3 Binary Log Probability Proper Scoring Rule for Correctness Prediction
The correctness prediction task measures the models’ ability to predict the correctness of their own predictions. We consider correctness estimators $\tilde{c}(x)\in[0,1]$ for inputs $x\in\mathcal{X}$ derived from the predictives. Framed as a binary prediction task, the goal of these estimators is to predict the probability of the predicted class’ correctness. In particular, for an (input, target) pair $(x,y)$ with $x\in\mathcal{X},y\in\mathcal{Y}$, we set the correctness target to
$\ell\equiv\ell(x,y)=1\left(\max_{c\in\{1,\dotsc,C\}}h_{c}(x)=y\right).$ (90)
Dropping the dependency on $x\in\mathcal{X}$ for brevity, the log probability score for binary targets $\ell\in\{0,1\}$ and estimators $\tilde{c}\in[0,1]$ is defined as
$S(\tilde{c},\ell)=\begin{cases}\log c&\mathrm{if}\ \ell=1\\ \log(1-\tilde{c})&\mathrm{if}\ \ell=0\end{cases}=\ell\log\tilde{c}+(1-\ell)\log(1-\tilde{c}).$ (91)
One can show that this is indeed a strictly proper scoring rule (Mucsányi et al., 2023).
I.4 Accuracy
For completeness, the accuracy of a predictive $h$ on a dataset $(x_{n},c_{n})_{n=1}^{N}$ is
$\operatorname{acc}(h;(x_{n},c_{n})_{n=1}^{N})=\frac{1}{N}\sum_{n=1}^{N}1\left(\operatorname*{arg\,max}_{c\in\{1,\dotsc,C\}}h_{c}(x_{n})=\tilde{c}_{n}\right).$ (92)
I.5 Area Under the Receiver Operating Characteristic Curve for Out-of-Distribution Detection
Out-of-distribution detection is another binary prediction task where a general uncertainty estimator $u(x)\in\mathbb{R}$ (derived from predictives or second-order Dirichlet distributions) is tasked to separate ID and OOD samples from a balanced mixture. The target OOD indicator variable $o(x)$ is, therefore, binary. As the uncertainty estimator can take on any real value, we measure the Area Under the Receiver Operating Characteristic curve (AUROC), which quantifies the separability of ID and OOD samples w.r.t. the uncertainty estimator.
Given uncertainty estimates $u_{n}\equiv u(x_{n})$ and target binary labels $o_{i}$ on a balanced dataset $(x_{n},o_{n})_{n=1}^{N}$, as well as a threshold $t\in\mathbb{R}$, we predict $1$ (out-of-distribution) when $u_{n}\geq t$ and 00 (in-distribution) when $u_{n}<t$. This lets us define the following index sets:
[…]
The Receiver Operating Characteristic (ROC) curve compares the following quantities:
$\mathrm{TPR}(t)$ $=\frac{|\mathrm{TP}(t)|}{|\mathrm{TP}(t)|+|\mathrm{FN}(t)|}=\frac{|\mathrm{TP}(t)|}{|\mathrm{P}|}$
$\mathrm{FPR}(t)$ $=\frac{|\mathrm{FP}(t)|}{|\mathrm{FP}(t)|+|\mathrm{TN}(t)|}=\frac{|\mathrm{FP}(t)|}{|\mathrm{N}|}.$
Here, FPR tells us how many of the actual negative samples in the dataset are recalled (predicted positive) at threshold $t$.
One can draw a curve of $(\mathrm{FPR}(t),\mathrm{TPR}(t))$ for all $t$ from $\infty$- ∞ to +$\infty$+ ∞. This is the ROC curve. The area under this curve quantifies how well the uncertainty estimator $u(x)$ can separate in-distribution and out-of-distribution inputs.
Appendix J Benchmarked Methods
This section describes our benchmarked methods and provides further implementation details.
J.1 Spectral Normalized Gaussian Process
Spectral normalized Gaussian processes (SNGP) (Liu et al., 2020) use spectral normalization of the parameter tensors for distance-awareness and a last-layer Gaussian process approximated by Fourier features to capture uncertainty. For an input $x\in\mathcal{X}$ and number of classes $C$, they predict a $C$-variate Gaussian distribution
$\mathcal{N}\left(\bm{B}\bm{\phi}(x),\bm{\phi}(x)^{\top}\left(\bm{\Psi}^{\top}\bm{\Psi}+\bm{I}\right)^{-1}\bm{\phi}(x)\bm{I}_{C}\right),$ (93)
in logit space.
  • $\bm{B}\in\mathbb{R}^{C\times D}$ is a learned parameter matrix that maps pre-logits to logits.
  • $\bm{\phi}(x)=\bm{\cos}\left(\bm{W}\bm{f}^{L-1}(x)+\bm{b}\right)\in\mathbb{R}^{D}$ is a random pre-logit embedding of the input $x\in\mathcal{X}$. $\bm{f}^{L-1}(x)$ denotes the pre-logit embedding. $\bm{W}$ is a fixed semi-orthogonal random matrix, and $\bm{b}$ is also a fixed random vector but sampled from $\text{Uniform}(0,2\pi)$.
  • $\bm{\Psi}^{\top}\bm{\Psi}$ is the (unnormalised) empirical covariance matrix of the pre-logits of the training set. This is calculated via accumulating the mini-batch estimates during the last epoch. [^2] As we use a cosine learning rate decay in all experiments, the model makes negligible changes in its pre-logit feature space in the last epoch. Thus, the empirical covariance matrix is approximately consistent.
The method applies spectral normalization to the hidden weights in each layer using a power iteration scheme with a single iteration per batch to obtain the largest singular value. Liu et al. (2020) claim this helps with input distance awareness.
J.2 Heteroscedastic Classifier
Heteroscedastic classifiers (HET) (Collier et al., 2021) construct a Gaussian distribution in the logit space to model per-input uncertainties:
$\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x)),$ (94)
where $\bm{f}(x)\in\mathbb{R}^{D}$ is the logit mean for input $x\in\mathcal{X}$ and
$\bm{\Sigma}(x)=\bm{V}(x)^{\top}\bm{V}(x)+\bm{\operatorname{\mathbf{diag}}}(\bm{d}(x))$ (95)
is a (positive definite) covariance matrix. Both the low-rank term $\bm{V}(x)$ and the diagonal term $\bm{d}(x)$ are calculated as a linear function of the pre-logit layer’s output.
To learn the per-input covariance matrices from the training set, one has to construct a predictive estimate from $\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x))$ using any of the methods in Appendix F. This predictive estimate is then trained using a standard cross-entropy (NLL) loss.
HET uses a temperature parameter to scale the logits before calculating the BMA. This is chosen using a validation set.
In Section 2.3, we show that when our constraint is satisfied, the off-diagonal terms of the covariance matrix do not matter for the predictive. This means that, in our framework, one can discard the low-rank term $\bm{V}(x)$ and only model the diagonal term $\bm{d}(x)$ without a decrease in expressivity. To keep comparisons fair and use the same backbone with the same number of parameters, we also only model $\bm{d}(x)$ for softmax-based predictives.
J.3 Laplace Approximation
The Laplace approximation (Daxberger et al., 2021) approximates the posterior $p(\bm{\theta}\mid\mathcal{D})$ over the network parameters $\bm{\theta}$ for a Gaussian prior $p(\bm{\theta})$ and likelihood defined by the network architecture by a Gaussian. In its simplest form, it uses the maximum a posteriori (MAP) weights $\bm{\theta}_{\text{MAP}}\in\mathbb{R}^{P}$ as the mean and the inverse Hessian of the regularized loss over the training set ~$\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})=\mathcal{L}(\bm{\theta};\mathcal{D})+\lambda\|\bm{\theta}\|_{2}^{2}$ evaluated at the MAP as the covariance matrix:
$\mathcal{N}\left(\bm{\theta}_{\text{MAP}},\left(\left.\frac{\partial^{2}\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})}{\partial\theta_{i}\partial\theta_{j}}\right|_{\bm{\theta}_{\text{MAP}}}\right)^{-1}\right)=\mathcal{N}\left(\bm{\theta}_{\text{MAP}},\left(\left.\frac{\partial^{2}\mathcal{L}(\bm{\theta};\mathcal{D})}{\partial\theta_{i}\partial\theta_{j}}\right|_{\bm{\theta}_{\text{MAP}}}+\lambda\bm{I}_{P}\right)^{-1}\right).$ (96)
This is a locally optimal post-hoc Gaussian approximation of the true posterior $p(\bm{\theta}\mid\mathcal{D})$ based on a second-order Taylor approximation. For details, see (Tatzel et al., 2025).
For deep neural networks, the Hessian matrix is often replaced with the Generalized Gauss-Newton (GGN) matrix. The GGN is guaranteed to be positive semidefinite even for suboptimal weights and has efficient approximation schemes, such as Kronecker-Factored Approximate Curvature (Martens & Grosse, 2015) or low-rank approximations.
Denoting our curvature estimate of choice by $\bm{G}$, the logit-space Gaussian is obtained by pushing forward the weight-space Gaussian measure through the linearised model around $\bm{\theta}_{\text{MAP}}$. For an input $x\in\mathcal{X}$, this results in
$\mathcal{N}\left(\left(\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\right)\bm{\theta}_{\text{MAP}},\left(\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\right)\left(\bm{G}+\lambda\bm{I}_{P}\right)^{-1}\left(\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\right)^{\top}\right),$ (97)
where $\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\in\mathbb{R}^{C\times P}$ is the model Jacobian matrix.
We use a last-layer KFAC Laplace variant in our experiments and use the full training set for calculating the GGN instead of a mini-batch based on recent works on the bias in mini-batch estimates (Tatzel et al., 2025).
Appendix K CIFAR-10 Experiments
This appendix section repeats the experiments presented in the main paper on the CIFAR-10 dataset. For a detailed description of the experimental setup, refer to Appendix H. Appendix I describes the used tasks and metrics.
As stated in the main paper, our two research questions are:
  • What are the effects of changing the learning objective? (Section K.2)
  • Do we have to sacrifice performance for sample-free predictives? (Section K.1)
K.1 Quality of Sample-Free Predictives
Table 2: Comparison of ECE results for different predictives using a fixed Laplace backbone.
Method Mean Std
Softmax Laplace
MC 100 0.0096 0.0013
MC 1000 0.0102 0.0016
MC 10 0.0120 0.0013
Mean Field 0.0121 0.0029
Laplace Bridge Predictive 0.5933 0.0105
NormCDF Laplace
Analytic 0.0074 0.0012
MC 1000 0.0092 0.0022
MC 100 0.0095 0.0015
MC 10 0.0100 0.0020
Similarly to the main paper, in this section, we investigate our first research question: whether there is a price to pay for sample-free predictives. Table 2 showcases the two best-performing (activation, method) pairs on the ECE metric and the CIFAR-10 dataset: Softmax and NormCDF Laplace. Mean Field (MF) is a strong alternative for sample-free predictives, but it has no guarantees and can fall behind MC sampling (see also Figure 1). When the constraint is satisfied, the analytic NormCDF predictive is exact. Empirically, our analytic predictives always perform on par with MC sampling.
K.2 Effects of Changing the Learning Objective
As in the main paper, in this section, we use the best-performing predictive and estimator (see Appendix F) for softmax models and employ our methods with the analytic predictives.
K.2.1 Calibration and Proper Scoring
We first evaluate calibration using the log probability scoring rule (Gneiting & Raftery, 2007) and the Expected Calibration Error (ECE) metric (Naeini et al., 2015). Figure 5 shows that on CIFAR-10, the score of our analytic predictives with implicit constraints (Sigmoid, NormCDF) are consistently better than the corresponding softmax results for all methods.
Figure 11: CIFAR-10 ECE results. Analytic predictives with implicit ( ) or explicit ( ) constraints outperform Softmax ( ) on HET and SNGP. Laplace tunes its hyperparameters based on the ECE metric – NormCDF Laplace is the overall best method. Note the restricted $y$-limits for readability.
Figure 11 shows that on CIFAR-10, our analytic predictives have a clear advantage on HET and SNGP. Laplace is a post-hoc method that tunes its hyperparameters on the ECE metric, hence its enhanced performance. Our NormCDF predictive is on par with Softmax.
K.2.2 Correctness Prediction
Figure 12: CIFAR-10 log probability proper scoring results for the binary correctness prediction task. Analytic predictives with implicit ( ) constraints consistently outperform Softmax ( ) on all methods.
Figure 12 shows that on the correctness prediction task, our analytic predictives with implicit constraints outperform all Softmax predictives across all methods, as measured by the log probability proper scoring rule.
K.2.3 Accuracy
Figure 13: Analytic predictives do not sacrifice accuracy. CIFAR-10 accuracies. Analytic predictives with implicit ( ) or explicit ( ) constraints either outperform or are on par with Softmax ( ) across all methods.
Analytic predictives do not sacrifice accuracy. Figure 13 evidences this claim on CIFAR-10: our analytic predictives either outperform or are on par with Softmax predictives. The most accurate method is Sigmoid HET. These results support the findings of Wightman et al. (2021) that showcase desirable training dynamics of the class-wise cross-entropy loss.
K.2.4 Out-of-Distribution Detection
Figure 14: CIFAR-10C OOD detection AUROC results for severity level one. Across all methods, the best-performing predictive is analytic with an implicit constraint ( ).
Finally, we consider the OOD detection task on a balanced mixture of ID (CIFAR-10) and OOD inputs. As OOD inputs, we consider corrupted CIFAR-10C samples. We use the AUROC metric to evaluate the methods’ performance. As shown in Figure 14, the best-performing method is Sigmoid SNGP, an analytic method. Generally, Softmax performs on par with our analytic predictives. Intuitively, separating ID and OOD samples does not require a fine-grained representation of uncertainty, unlike the ECE or proper scoring rules. Nevertheless, the analytic predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5).
Appendix L ImageNet Dirichlet Evaluation
This appendix section evaluates our moment-matched Dirichlet distributions on the OOD detection task (Appendix I). Table 3 shows the per-estimator performances of the three best-performing (method, activation) pairs on ImageNet: Sigmoid Laplace, NormCDF HET, and NormCDF Laplace. In all three cases, the best estimator is derived from the moment-matched Dirichlet distributions.
Table 3: Comparison of OOD AUROC results for different uncertainty estimators on the three best-performing (method, activation) pairs: Sigmoid Laplace, NormCDF HET, and NormCDF Laplace. The best estimator is derived from the moment-matched Dirichlet distributions in all three cases.
Method Mean Std
Sigmoid Laplace
Dirichlet Mutual Information 0.6388 0.0005
Analytic Predictive Entropy 0.6353 0.0003
MC 1000 Predictive Entropy 0.6323 0.0001
MC 100 Predictive Entropy 0.6323 0.0001
MC 1000 Expected Entropy 0.6321 0.0001
NormCDF HET
Dirichlet Expected Entropy 0.6353 0.0021
MC 1000 Expected Entropy 0.6281 0.0020
MC 100 Expected Entropy 0.6281 0.0020
MC 10 Expected Entropy 0.6281 0.0020
Analytic Predictive Entropy 0.6277 0.0021
NormCDF Laplace
Dirichlet Expected Entropy 0.6321 0.0009
MC 100 Predictive Entropy 0.6296 0.0012
MC 1000 Predictive Entropy 0.6296 0.0012
Closed-Form Predictive Entropy 0.6296 0.0012
MC 10 Predictive Entropy 0.6294 0.0012
Appendix M Further Out-of-Distribution Detection Results
Figure 15 shows OOD detection results across all ImageNet-C severity levels. Our analytic predictives with implicit constraints consistently outperform Softmax.
(a) OOD detection AUROC with severity level one.
(b) OOD detection AUROC with severity level two.
(c) OOD detection AUROC with severity level three.
(d) OOD detection AUROC with severity level four.
(e) OOD detection AUROC with severity level five.
Figure 15: The OOD detection performance of all methods increases steadily as we increase the severity of the perturbed half of the mixed dataset on the ImageNet validation dataset. Our analytic predictives consistently outperform Softmax.

### v2

@@
 Such outputs only capture aleatoric uncertainty.
-To capture epistemic uncertainty, approximate Gaussian inference methods have been proposed, which output Gaussian distributions over the logit space.
+To capture epistemic uncertainty, approximate Gaussian inference methods have been proposed.
+We develop a common formalism to describe such methods, which we view as outputting Gaussian distributions over the logit space.
 Predictives are then obtained as the expectations of the Gaussian distributions pushed forward through the softmax.
 However, such softmax Gaussian integrals cannot be solved analytically, and Monte Carlo (MC) approximations can be costly and noisy.
-We propose a simple change in the learning objective which allows the exact computation of predictives and enjoys improved training dynamics, with no runtime or memory overhead.
-This framework is compatible with a family of output activation functions that includes the softmax, as well as element-wise normCDF and sigmoid.
-Moreover, it allows for approximating the Gaussian pushforwards with Dirichlet distributions by analytic moment matching.
-We evaluate our approach combined with several approximate Gaussian inference methods (Laplace, HET, SNGP) on large- and small-scale datasets (ImageNet, CIFAR-10), demonstrating improved uncertainty quantification capabilities compared to softmax MC sampling.
-Code is available at github.com/bmucsanyi/probit.
-Deep learning, Uncertainty quantification, Probabilistic modeling, Bayesian inference
-[…]
-[ linewidth=1pt, roundcorner=3pt, innerleftmargin=4pt, innerrightmargin=4pt, innertopmargin=-5pt, innerbottommargin=4pt, leftmargin=0pt, rightmargin=0pt, skipabove=2pt, skipbelow=0pt, ]eqbox
+We propose to replace the softmax activation by element-wise normCDF or sigmoid, which allows for the accurate sampling-free approximation of predictives.
+This also enables the approximation of the Gaussian pushforwards by Dirichlet distributions with moment matching.
+This approach entirely eliminates the runtime and memory overhead associated with MC sampling.
+We evaluate it combined with several approximate Gaussian inference methods (Laplace, HET, SNGP) on large- and small-scale datasets (ImageNet, CIFAR-100, CIFAR-10), demonstrating improved uncertainty quantification capabilities compared to softmax MC sampling.
 1 Introduction
-Figure 1: Mean KL divergence between the ‘true’ predictive (approximated with a $10000$ sample MC approximation) and different predictive approximations on a synthetic data set.
-This synthetic dataset consists of 100 logit means $\bm{\mu}$ and standard deviations $\bm{\sigma}$, whose components are i.i.d.
-uniformly distributed.
-For a fair comparison we scale the logit dataset for the different activations to match the activations to first order, so that $\mu_{c}\in[-1,1]$, $\sigma_{c}\in[0,1]$ for sigmoid, $\mu_{c}\in[-1/2-\log 2,1/2-\log 2]$, $\sigma_{c}\in[0,1/2]$ for softmax, $\mu_{c}\in[-\sqrt{\pi/8},\sqrt{\pi/8}]$, $\sigma_{c}\in[0,\pi/8]$ for normCDF.
-The MC approximations are capped at a fixed computational budget of $10000$ class samples, i.e.
-$\lceil 10000/C\rceil$ samples.
-We see that, even without enforcing the constraint Equation 18, our approximations are the only ones that do not become worse with the number of classes.
-See Appendix C for theoretical analyses of the quality of the different approximations.
 Given an input space $\mathcal{X}$, $C$ classes and training data $(x_{n},c_{n})_{n=1}^{N}\subset\mathcal{X}\times\{1,\dots,C\}$, the goal of probabilistic classification is to learn a function $\bm{h}\colon\mathcal{X}\to\Delta^{C-1}$, where
@@
 In this case, $\mathbb{R}^{C}$ is called the logit space, and, for $x\in\mathcal{X}$, $\bm{f}(x)$ is a logit.
-With the usual assumption of i.i.d. data, the likelihood under the model $\bm{h}$ is given by ($\delta$ is the Kronecker delta)
-$p((c_{n})_{n=1}^{N}\mid(x_{n})_{n=1}^{N})$ $=\prod_{n=1}^{N}p(c_{n}\mid x_{n})$ (3)
-$=\prod_{n=1}^{N}\prod_{c=1}^{C}h_{c}(x_{n})^{\delta_{c_{n},c}}.$
-This yields a natural loss function for classification problems, the cross-entropy (CE) loss
-$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ $=-\log(p((c_{n})_{n=1}^{N}\mid(x_{n})_{n=1}^{N}))$ (4)
-$=-\sum_{n=1}^{N}\sum_{c=1}^{C}\delta_{c_{n},c}\log(h_{c}(x_{n})).$
+With the usual assumption of i.i.d. data, the likelihood under the model $\bm{h}$ is given by
+$p((c_{n})_{n=1}^{N}\mid(x_{n})_{n=1}^{N})$ $=\prod_{n=1}^{N}p(c_{n}\mid x_{n})=\prod_{n=1}^{N}\prod_{c=1}^{C}h_{c}(x_{n})^{\delta_{c_{n},c}},$ (3)
+where $\delta$ is the Kronecker delta. This yields a natural loss function for classification problems, the cross-entropy (CE) loss
+$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ $=-\log(p((c_{n})_{n=1}^{N}\mid(x_{n})_{n=1}^{N}))=-\sum_{n=1}^{N}\sum_{c=1}^{C}\delta_{c_{n},c}\log(h_{c}(x_{n})).$ (4)
 For typical models $\bm{h}$ trained through such approximate maximum likelihood estimation, an output probability vector $\bm{h}(x)\in\Delta^{C-1}$ should be interpreted as an estimate of $p_{\text{gen}}(\bm{y}\mid x)$, the probability under the generative model.
 Notably, the output probability does not take into account the uncertainty of the model due to the finite nature of the data.
-In the uncertainty quantification formalism, such models can only estimate aleatoric uncertainty and disregard epistemic uncertainty (Hüllermeier & Waegeman, 2021).
-The probabilistic way to capture epistemic uncertainty is to require the model to output a second-order distribution (as in “a distribution over probability distributions”).
-That is, for each input $x\in\mathcal{X}$, the model should output a probability measure over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$, as opposed to a point estimate in the simplex $\bm{h}(x)\in\Delta^{C-1}$ (Sensoy et al., 2018; Charpentier et al., 2020).
+In the uncertainty quantification formalism, such models can only estimate aleatoric uncertainty and disregard epistemic uncertainty [15].
+Figure 1: Mean KL divergence between the ‘true’ predictive (approximated with a $10000$ sample MC approximation) and different predictive approximations on a synthetic data set.
+This synthetic dataset consists of 100 logit means $\bm{\mu}$ and standard deviations $\bm{\sigma}$, whose components are i.i.d.
+uniformly distributed.
+For a fair comparison we scale the logit dataset for the different activations to match the activations to first order, so that $\mu_{c}\in[-1,1]$, $\sigma_{c}\in[0,1]$ for sigmoid, $\mu_{c}\in[-1/2-\log 2,1/2-\log 2]$, $\sigma_{c}\in[0,1/2]$ for softmax, $\mu_{c}\in[-\sqrt{\pi/8},\sqrt{\pi/8}]$, $\sigma_{c}\in[0,\pi/8]$ for normCDF.
+The MC approximations are capped at a fixed computational budget of $10000$ class samples, i.e.
+$\lceil 10000/C\rceil$ samples (see also Section C.1).
+The probabilistic way to capture epistemic uncertainty is to require the model to output a second-order distribution (that is ‘a distribution over probability distributions’).
+Namely, for each input $x\in\mathcal{X}$, the model should output a probability measure over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$, as opposed to a point estimate in the simplex $\bm{h}(x)\in\Delta^{C-1}$ [32, 4].
 A number of methods for such distributional uncertainty quantification in classification rely on approximate Gaussian inference in logit space. That is, $\mathrm{h}$ is written as a composition
 $\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\operatorname{\mathbf{softmax}}_{*}}\mathcal{P}(\Delta^{C-1})$ (5)
-where $f$ is learned from the training data, $\mathcal{G}(\mathbb{R}^{C})$ is the set of Gaussian measures on the logit space $\mathbb{R}^{C}$, and $\operatorname{\mathbf{softmax}}_{*}$ pushes forward the Gaussian probability measures through the softmax.
-For example, Heteroscedastic Classifiers (HET) (Collier et al., 2021) learn the output means and covariances of $\mathrm{f}$ explicitly with a neural network.
-In linearised Laplace approximations (Daxberger et al., 2021), only the means are learned explicitly with a neural network, while the covariances are obtained post-hoc from that neural network by approximating its parameter posterior distribution with a Gaussian, and then pushing it forward to logit space by linearising the network in the parameters.
-Last-layer Laplace approximations (Kristiadi et al., 2020) are a variant of linearised Laplace approximations where only the posterior over the last-layer parameters is Gaussian-approximated, requiring no linearisation of the network.
-Spectral-Normalized Gaussian processes (SNGP) (Liu et al., 2020) are neural networks with last layers that are approximately Gaussian processes, through random feature expansions and Laplace approximations.
-To obtain predictive probabilities from a distributional classifier as in Equation 5, we marginalise it out by the measure-theoretic change of variables, using $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$:
-$\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]$ $=\int_{\Delta^{C-1}}p\;\mathrm{h}(x)(dp)$ (6)
-$=\int_{\mathbb{R}^{C}}\operatorname{\mathbf{softmax}}(y)\;\mathrm{f}(x)(dy).$
-In addition to predictive probabilities, the distributional framework allows for the acquisition of other quantities of interest, such as variances, entropy, and information-theoretic decompositions of the predictive uncertainty into aleatoric and epistemic parts (Depeweg et al., 2018; Mukhoti et al., 2023; Wimmer et al., 2023).
+where $\mathrm{f}$ is learned from the training data, $\mathcal{G}(\mathbb{R}^{C})$ is the set of Gaussian measures on the logit space $\mathbb{R}^{C}$, and $\operatorname{\mathbf{softmax}}_{*}$ pushes forward the Gaussian probability measures through the softmax.
+For example, Heteroscedastic Classifiers (HET) [5] learn the output means and covariances of $\mathrm{f}$ explicitly with a neural network.
+In linearised Laplace approximations [8], only the means are learned explicitly with a neural network, while the covariances are obtained post-hoc from that neural network by approximating its parameter posterior distribution with a Gaussian, and then pushing it forward to logit space by linearising the network in the parameters.
+Last-layer Laplace approximations [16] are a variant of linearised Laplace approximations where only the posterior over the last-layer parameters is Gaussian-approximated, requiring no linearisation of the network.
+Spectral-Normalized Gaussian processes (SNGP) [20] are neural networks with last layers that are approximately Gaussian processes, through random feature expansions and Laplace approximations.
+To obtain predictive probabilities from a distributional classifier as in Eq.
+5, we marginalise it out by the measure-theoretic change of variables, using $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$:
+$\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]$ $=\int_{\Delta^{C-1}}p\;\mathrm{h}(x)(dp)=\int_{\mathbb{R}^{C}}\operatorname{\mathbf{softmax}}(y)\;\mathrm{f}(x)(dy).$ (6)
+In addition to predictive probabilities, the distributional framework allows for the acquisition of other quantities of interest, such as variances, entropy, and information-theoretic decompositions of the predictive uncertainty into aleatoric and epistemic parts [10, 28, 39].
 These are usually obtained by Monte Carlo (MC) sampling from the Gaussian measure $\mathrm{f}(x)$.
-Indeed, while the probability density of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ can be obtained in closed form using a change of variables formula (Atchison & Shen, 1980), this does not seem of practical use.
-Even predictives cannot be computed analytically.
+Indeed, while the probability density of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ can be obtained in closed form using a change of variables formula [1], this does not seem of practical use—even predictives cannot be computed analytically.
 On the other hand, given a variance, the runtime and memory cost required for an MC approximation to achieve an estimator of said variance grows linearly with the number of classes.
 Crucially, this computational cost is not limited to training time, but also to inference time, when the computational budget may be much more limited.
-This limits the application of such Gaussian inference methods in classification tasks with large number of classes, such as those common in computer vision and natural language processing.
+This limits the application of such Gaussian inference methods in classification with a large number of classes, such as those common in computer vision and language modeling.
 1.1 Proposed Recipe
-We can summarise our framework as a 4-step recipe:
-  1. 1.
-    Choose an approximate Gaussian inference method: a method that outputs logit covariances.
-  2. 2.
-    Choose an output activation for which the corresponding Gaussian integral can be solved (Section 2).
-  3. 3.
-    Train with an objective that enforces a certain normalisation constraint (Section 4).
-  4. 4.
-    At inference time, use our analytic formulas (see Section 2 for the predictives, Section 3 for other quantities of interest).
-2 Gaussian Inference with Analytic Predictives
+We summarise our framework as a 3-step recipe:
+  1. (i)
+    Choose an approximate Gaussian inference method that outputs logit means and covariances.
+  2. (ii)
+    Choose an output activation that yields a closed-form Gaussian integral (Section 2).
+  3. (iii)
+    At inference time, use our closed-form formulas (see Section 2 for the predictives, Section 3 for other quantities of interest).
+2 Gaussian Inference with Closed-Form Predictives
 2.1 Motivation from Softmax Models
@@
 The softmax function is the composition $\bm{\operatorname{\mathbf{softmax}}}=\bm{n}\circ\bm{\exp}$, where $\bm{n}$ is the normalisation function given by $\bm{n}(\bm{q}):=\bm{q}/(q_{1}+\dots+q_{C})$ and $\bm{\exp}$ is applied element-wise.
-Thus, $\mathrm{h}$ can be written as the composition (c.f.
-Equation 5)
-$\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\bm{\exp}_{*}}\mathcal{P}(\mathbb{R}_{>0}^{C})\xrightarrow{\bm{n}_{*}}\mathcal{P}(\Delta^{C-1}).$ (7)
-Now it is easy to see that (Section A.1)
+Thus, $\mathrm{h}$ can be written as the composition (see Eq.
+5)
+$\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\bm{\exp}_{*}}\mathcal{P}((0,\infty)^{C})\xrightarrow{\bm{n}_{*}}\mathcal{P}(\Delta^{C-1}).$ (7)
+Now it is straightforward to see that (Section A.1)
 $\int_{\mathbb{R}}\exp(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)=\exp\left(\mu+\frac{\sigma^{2}}{2}\right)$ (8)
 and thus
-$\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}[\bm{Q}]$ $=\int_{\mathbb{R}^{C}}\bm{\exp}(\bm{y})\;\mathrm{f}(x)(d\bm{y})$ (9)
-$=\bm{\exp}\left(\bm{\mu}(x)+\frac{\bm{\sigma}^{2}(x)}{2}\right)$
+$\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}[\bm{Q}]$ $=\int_{\mathbb{R}^{C}}\bm{\exp}(\bm{y})\;\mathrm{f}(x)(d\bm{y})=\bm{\exp}\left(\bm{\mu}(x)+\frac{\bm{\sigma}^{2}(x)}{2}\right)$ (9)
 where $\bm{\mu}(x)$ and $\bm{\sigma}^{2}(x)$ are, respectively, the mean and the diagonal of the covariance of $\mathrm{f}(x)$, and the vector division is element-wise.
-Note that no approximations are made so far.
-We can use this to approximate the predictive of the model in closed form:
-$\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]$ $=\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}\left[\frac{\bm{Q}}{\sum_{c=1}^{C}Q_{c}}\right]$ (10)
-$\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}[\bm{Q}]}{\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}\left[\sum_{c=1}^{C}Q_{c}\right]}$
-$=\frac{\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}[\bm{Q}]}{\sum_{c=1}^{C}\mathbb{E}_{Q_{c}\sim\exp_{*}\mathrm{f}_{c}(x)}\left[Q_{c}\right]}$
+Note that no approximations have been made so far.
+One can use this to approximate the model’s predictive in closed form:
+$\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]=\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}\left[\frac{\bm{Q}}{\sum_{c=1}^{C}Q_{c}}\right]$ $\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}[\bm{Q}]}{\mathbb{E}_{\bm{Q}\sim\bm{\exp}_{*}\mathrm{f}(x)}\left[\sum_{c=1}^{C}Q_{c}\right]}$ (10)
 $=\frac{\bm{\exp}\left(\bm{\mu}(x)+\frac{\bm{\sigma}^{2}(x)}{2}\right)}{\sum_{c=1}^{C}\exp\left(\mu_{c}(x)+\frac{\sigma^{2}_{c}(x)}{2}\right)}.$
 2.2 The General Framework
-In fact, we can apply the previous argumentation to a general family of output activation functions: suppose we can write the output activation as $\bm{n}\circ\bm{\varphi}$, where $\bm{\varphi}$ is the element-wise application of some activation function $\varphi\colon\mathbb{R}\to\mathbb{R}_{>0}$.
-So we have (c.f.
-Equation 7)
-$\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\bm{\varphi}_{*}}\mathcal{P}(\mathbb{R}_{>0}^{C})\xrightarrow{\bm{n}_{*}}\mathcal{P}(\Delta^{C-1}).$ (11)
-If we can solve the one-dimensional integrals (c.f. Equation 8)
-$\int_{\mathbb{R}}\varphi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)$ (12)
-then we can compute the approximate predictive (c.f. Equation 10)
+The previous argumentation can be applied to a general family of output activation functions: suppose the output activation is of the form $\bm{n}\circ\bm{\varphi}$, where $\bm{\varphi}$ is the element-wise application of some activation function $\varphi\colon\mathbb{R}\to(0,\infty)$.
+Then we have (see Eq.
+7)
+$\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\bm{\varphi}_{*}}\mathcal{P}((0,\infty)^{C})\xrightarrow{\bm{n}_{*}}\mathcal{P}(\Delta^{C-1}).$ (11)
+Solving the one-dimensional integrals $\int_{\mathbb{R}}\varphi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)$ (see Eq. 8) allows us to obtain the approximate predictive (see Eq. 10)
 $\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{Q}]}{\sum_{c=1}^{C}\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}\left[Q_{c}\right]}.$ (12)
-For instance, taking $\varphi=\Phi$, the standard Normal cumulative distribution function (normCDF) [^1] The resulting model is distinct from a multinomial probit model, see Appendix B., we have by a classical result (Section A.2),
-$\int_{\mathbb{R}}\Phi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{1+\sigma^{2}}}\right)$ (14)
+For instance, taking $\varphi=\Phi$, the standard Normal cumulative distribution function (normCDF) [^1] The resulting model is distinct from a multinomial probit model, see Appendix B., we have by a classical result (Section A.2)
+$\int_{\mathbb{R}}\Phi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{1+\sigma^{2}}}\right).$ (13)
 and thus
 $\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]\approx\frac{\bm{\Phi}\left(\frac{\bm{\mu}(x)}{\sqrt{1+\bm{\sigma}^{2}(x)}}\right)}{\sum_{c=1}^{C}\Phi\left(\frac{\mu_{c}(x)}{\sqrt{1+\sigma^{2}_{c}(x)}}\right)}.$ (14)
-Taking instead $\varphi=\rho$, the logistic sigmoid, we can approximate Equation 12 with the probit approximation (Section A.3)
+Taking instead $\varphi=\rho$, the logistic sigmoid, we approximate Eq. 13 using the probit approximation (Section A.3),
 $\int_{\mathbb{R}}\rho(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)$ (15)
@@
 $\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]\approx\frac{\bm{\rho}\left(\frac{\bm{\mu}(x)}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}(x)}}\right)}{\sum_{c=1}^{C}\rho\left(\frac{\mu_{c}(x)}{\sqrt{1+\frac{\pi}{8}\sigma^{2}_{c}(x)}}\right)}.$ (16)
-2.3 Making the Approximations Exact
-At this point, the approximate Equation 13 may seem ad-hoc.
-However, note that Equation 13 is exact when $\sum_{c=1}^{C}Q_{c}$ is a constant random variable, as can be seen from the derivation in Equation 10.
-Thus, we would like to enforce a constraint in the model of the form
-$\sum_{c=1}^{C}Q_{c}$ $=s(x)$ $\text{where }\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)$ (18)
-where $s\colon\mathcal{X}\to\mathbb{R}$ is a deterministic function. In other words, the law of $\sum_{c=1}^{C}Q_{c}$ is a Dirac measure for all $x$.
-In neural networks, constraints can be learned through appropriate choices of learning objectives during training, ensuring that they are satisfied within the training data distribution.
-In Section 4, we describe how this can be achieved depending on $\bm{\varphi}$ and the method that is used to obtain the mapping to Gaussian distributions $\mathrm{f}$.
-Given the constraint of Equation 18 is satisfied, Equations 10 and 15 ($\varphi=\exp\text{ or }\Phi$) both become exact.
-For Equation 17 ($\varphi=\rho$), the only error incurred is the one from the probit approximation (Equation 16) per each class.
-In the synthetic experiment (Figure 1), we see that this induces no significant effect on the quality of the predictive.
-We justify this observation with the following result:
+2.3 Theoretical Analysis of the Approximation’s Quality
+In Fig.
+1, we see empirically that the quality of our predictive approximations outperforms existing predictive approximations on synthetic data.
+We now explore the quality of our predictive approximations theoretically.
+We conduct the analysis for the case $\varphi=\rho$.
+Let
+$q=q(\mu,\sigma^{2}):=\mathbb{E}_{Q\sim\rho_{*}\mathcal{N}(\mu,\sigma^{2})}[Q]\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)=:\hat{q}(\mu,\sigma^{2})=\hat{q},$ (17)
+where $\hat{q}$ is the probit approximation (Section A.3).
+Moreover we write $q_{c}:=q(\mu_{c},\sigma_{c}^{2})$, $\hat{q}_{c}:=\hat{q}(\mu_{c},\sigma_{c}^{2})$ where $(\mu_{c},\sigma_{c}^{2})$ is the logit mean-variance pair for class $c$, $\bm{p}$ for the true predictive and $\bm{\hat{p}}$ for our approximate predictive, i.e.
+$\bm{p}:=\mathbb{E}_{\bm{P}\sim(\bm{n}\circ\bm{\rho})_{*}\mathcal{N}(\bm{\mu},\bm{\operatorname{\mathbf{diag}}}(\bm{\sigma}^{2}))}[\bm{P}],\quad\bm{\hat{p}}:=\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}}.$ (18)
+To analyse the quality of the approximation, we thus want to bound $\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$.
+Clearly, when all the $Q_{c}$ are deterministic (i.e.
+constant random variables), where $\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)$, we have $\bm{p}=\bm{\hat{p}}$.
+Alternatively, from the derivation Eq.
+10, we see that in the limiting case Eq.
+12 is exact when $\sum_{c=1}^{C}Q_{c}$ is deterministic.
+This is a weaker assumption than requiring all $Q_{c}$ to be deterministic individually.
+In fact, we can derive a rate of decay of $\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$ as a function of $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)$:
 Theorem 2.1.
-Let $\varphi=\rho$.
-Suppose that the logit means and variances $(\mu_{c},\sigma^{2}_{c})$ lie in some compact set $\mathcal{K}\subset\mathbb{R}\times\mathbb{R}_{>0}$ for each class $c$.
-Then, assuming the constraint Equation 18 is perfectly satisfied, the KL divergence between the true and approximate predictive Equation 17 is bounded by some constant $M(\mathcal{K})$ only depending on $\mathcal{K}$, not on $C$.
-Section C.2 provides a more precise statement, a formula for $M(\mathcal{K})$, and a proof. Importantly, this result states that the approximation error does not increase with the number of classes.
-Finally, observe that, when enforcing Equation 18, we obtain exact predictives with Equation 13, which only uses the diagonal of $\mathrm{f}(x)$’s covariance.
-Enforcing this constraint thus allows one to ignore non-diagonal terms in the logit covariances.
-In Section 5, we discuss how this can be leveraged for computational gains.
+Suppose the means and variances $(\mu_{c},\sigma_{c}^{2})$ lie in some compact set $\mathcal{K}\subset\mathbb{R}\times[0,\infty)$ for each class $c$. Then
+$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})=M(\mathcal{K})+O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$ (19)
+as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, for some constant $M(\mathcal{K})\in[0,\infty]$ which only depends on $\mathcal{K}$, not on $C$, such that $M(\mathcal{K})<\infty$ when $\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$, and $M(\mathcal{K})=0$ when $q=\hat{q}$ for all $(\mu,\sigma^{2})\in\mathcal{K}$.
+Section C.2 provides a formula for $M(\mathcal{K})$ and a proof.
+We also conduct an empirical investigation of whether the assumptions made in the theorem hold in practice, namely whether the logit means and variances lie in a compact set, and whether $M(\mathcal{K})<\infty$.
+The $O$ term in Eq.
+19 is due to the splitting of the expectation into the numerator and denominator in the derivation Eq.
+10, while $M(\mathcal{K})$ results from the error due to the probit approximation.
+Importantly, Theorem 2.1 states that the latter does not increase with the number of classes.
 3 Distributional Approximations of the Gaussian Pushforwards
@@
 As previously discussed, such a distribution allows, for instance, closed-form decompositions of aleatoric and epistemic uncertainties.
-The formulation of Equation 11 does not yet allow this. However, we will now see that we can obtain good Dirichlet approximations by constructing a tractable approximation of Equation 11 of the form
+The formulation of Eq. 11 does not yet allow this. However, we will now see that one can obtain good Dirichlet approximations by constructing a tractable approximation of Eq. 11 of the form
 $\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\quad\mathrm{a}\quad}\mathcal{D}(\Delta^{C-1}),$ (20)
 where $\mathcal{D}(\Delta^{C-1})$ is the space of Dirichlet distributions on $\Delta^{C-1}$.
-The family of Dirichlets presents an ideal choice of distributions over the simplex as, for such distributions, all quantities of interest can be computed analytically (c.f.
-Appendix G).
+The family of Dirichlets presents an ideal choice of distributions over the simplex as, for such distributions, all quantities of interest are closed-form (see Appendix F).
 The map $\mathrm{a}$ is constructed by moment matching: we match the first two moments of the Dirichlet distribution to the moments of the Gaussian pushforward $\bm{\varphi}_{*}\mathrm{f}(x)$.
-Conveniently, just as the first moments described in Section 2, we can obtain second moments for $\exp$
+Conveniently, like the first moments described in Section 2, we can obtain second moments for $\exp$
 $\int_{\mathbb{R}}\exp(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)=\exp(2\mu+2\sigma^{2})$ (21)
-and for $\Phi$ (Owen (1980, Eq. 20,010.4))
-$\int_{\mathbb{R}}\Phi(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)$ (21)
-$\quad=\Phi\left(\frac{\mu}{\sqrt{1+\sigma^{2}}}\right)-2T\left(\frac{\mu}{\sqrt{1+\sigma^{2}}},\frac{1}{\sqrt{1+2\sigma^{2}}}\right),$
-where $T$ is Owen’s T function. For $\rho$, we have the following approximation (Daunizeau (2017, Equation 23))
-$\int_{\mathbb{R}}\rho(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)$ (22)
-$-\frac{\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\left(1-\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\right)}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}.$
+and for $\Phi$ (Owen [30, Eq. 20,010.4])
+$\int_{\mathbb{R}}\Phi(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{1+\sigma^{2}}}\right)-2T\left(\frac{\mu}{\sqrt{1+\sigma^{2}}},\frac{1}{\sqrt{1+2\sigma^{2}}}\right),$ (22)
+where $T$ is Owen’s T function. For $\rho$, we have the following approximation (Daunizeau [7, Equation 23])
+$\int_{\mathbb{R}}\rho(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)-\frac{\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\left(1-\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\right)}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}.$ (23)
 This enables us to compute the second moment of the pushforward distribution via
-$\mathbb{E}_{\bm{P}\sim\bm{n}_{*}\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{P}^{2}]$ $=\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}\left[\frac{\bm{Q}^{2}}{\left(\sum_{c=1}^{C}Q_{c}\right)^{2}}\right]$ (23)
-$\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{Q}^{2}]}{\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}\left[\sum_{c=1}^{C}Q_{c}\right]^{2}}.$
-As in Equation 13, the approximation in Equation 23 is exact when $\sum_{c=1}^{C}Q_{c}$ is a constant random variable, as under the constraint of Equation 18.
-If we were to directly attempt to infer the parameters of the Dirichlet $\mathrm{h}(x)$ from Equations 13 and 23, we would obtain an overparametrised system of equations with $2C$ equations and $C$ unknowns.
-Thus, we instead use a classical Dirichlet method of moments (Minka, 2000, Eqs.
-19 & 23), which reduces these equations, as we will now describe.
+$\mathbb{E}_{\bm{P}\sim\bm{n}_{*}\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{P}^{2}]$ $=\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}\left[\frac{\bm{Q}^{2}}{\left(\sum_{c=1}^{C}Q_{c}\right)^{2}}\right]\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{Q}^{2}]}{\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}\left[\sum_{c=1}^{C}Q_{c}\right]^{2}}.$ (24)
+As in Eq. 12, the approximation in Eq. 24 is exact when $\sum_{c=1}^{C}Q_{c}$ is a constant random variable.
+If we were to directly attempt to infer the parameters of the Dirichlet $\mathrm{h}(x)$ from Eqs.
+12 and 24, we would obtain an overparametrised system of equations with $2C$ equations and $C$ unknowns.
+Thus, we instead use a classical Dirichlet method of moments [25, Eqs.
+19 & 23], which reduces these equations, as we will now describe.
 The sum of the Dirichlet parameters $\sum_{c=1}^{C}\gamma_{c}$ may be estimated by
@@
 for any $1\leq c\leq C$.
-So, a natural estimate of $\sum_{c=1}^{C}\gamma_{c}$ that uses all $P_{c}$ is the geometric mean of Equation 24.
-Since the mean of the Dirichlet is $\bm{\gamma}/\sum_{c=1}^{C}\gamma_{c}$, we obtain using Equation 24 an expression for the Dirichlet parameters
+So, a natural estimate of $\sum_{c=1}^{C}\gamma_{c}$ that uses all $P_{c}$ is the geometric mean of Eq.
+25.
+Since the mean of the Dirichlet is $\bm{\gamma}/\sum_{c=1}^{C}\gamma_{c}$, we obtain using Eq.
+25 an expression for the Dirichlet parameters
 $\left(\prod_{c=1}^{C}\frac{\mathbb{E}[P_{c}]-\mathbb{E}[P^{2}_{c}]}{\mathbb{E}[P^{2}_{c}]-\mathbb{E}[P_{c}]^{2}}\right)^{1/C}\mathbb{E}[\bm{P}].$ (26)
-Hence, using Equations 13 and 23, we obtain the computationally feasible Dirichlet parameters
+Hence, using Eqs. 12 and 24, we obtain the computationally feasible Dirichlet parameters
 $\bm{\gamma}:=\left(\prod_{c=1}^{C}\frac{\mathbb{E}[Q_{c}]\cdot S-\mathbb{E}[Q^{2}_{c}]}{\mathbb{E}[Q^{2}_{c}]-\mathbb{E}[Q_{c}]^{2}}\right)^{1/C}\left(\frac{\mathbb{E}[\bm{Q}]}{\sum_{c=1}^{C}\mathbb{E}[Q_{c}]}\right)$ (27)
 for $\mathrm{h}(x)$, where $S=\max(\sum_{c=1}^{C}\mathbb{E}[Q_{c}],1)$.
-Taking the maximum with $1$ in the expression for $S$ is an additional approximation to ensure $\bm{\gamma}\geq\bm{0}$, and is justified when taking $s(x)=1$ for all $x$ in Equation 18.
-Note that performing such moment matching to obtain the Dirichlet parameters in Equation 26, the mean of this Dirichlet matches precisely the approximate predictive Equation 13.
-That is, the approximate predictive of the exact distributional model (Equation 11) matches the exact predictive of the approximate distributional model (Equation 19).
+Taking the maximum with $1$ in the expression for $S$ is an additional approximation to ensure $\bm{\gamma}>\bm{0}$.
+Note that by performing such moment matching to obtain the Dirichlet parameters in Eq.
+27, the mean of this Dirichlet matches precisely the approximate predictive Eq.
+12.
+That is, the approximate predictive of the exact distributional model (Eq.
+11) matches the exact predictive of the approximate distributional model (Eq.
+20).
 3.2 Intermediate Beta Matching
@@
 Moreover, when $\mathrm{f}(x)$ is the standard Normal $\mathcal{N}(0,1)$ (on the left in blue), the approximation for the pushforward through normCDF is exact (Section D.1).
-When $\varphi=\Phi\text{ or }\rho$, noting that their range is $(0,1)$, we can decompose the approximate distributional model (Equation 19) further:
+When $\varphi=\Phi\text{ or }\rho$, noting that their range is $(0,1)$, we can decompose the approximate distributional model (Eq. 20) further:
 $\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\mathrm{p}}\mathcal{B}((0,1))^{C}\xrightarrow{\mathrm{n}}\mathcal{D}(\Delta^{C-1}),$ (28)
@@
 In Section D.1, we give an information-geometric motivation for this, which shows that the choice of activation $\Phi$ is ideal for approximating the Gaussian pushforwards with Beta distributions.
-The quality of such Beta approximations is encouraging for that of the output Dirichlet approximations: observe that taking $s(x)=1$ for all $x$ in Equation 18, $\bm{n}$ becomes the identity (as it normalises by 1), and thus all $\mathrm{n}$ does is combining the $C$ Beta approximations into one Dirichlet.
-4 Learning the Constraint
-As discussed in Sections 2.3 and 3, the quality of our approximations relies on the enforcement of the constraint $\sum_{c=1}^{C}Q_{c}=s(x)$ for some deterministic function $s\colon\mathcal{X}\to\mathbb{R}$ (Equation 18).
-In this work, we simply choose $s(x)=1$ for all $x$.
-We thus make use of the degree of freedom in the logits to obtain exact predictives in Equation 13.
-In Appendix E, we make the connection with the energy-based model view of classifiers (Grathwohl et al., 2019), which instead makes use of this degree of freedom to turn discriminative models into generative models.
-Future work could combine both frameworks by learning $s(x)\propto p_{\operatorname{gen}}(x)$.
-4.1 Explicit Constraint
-Up to this point, we have treated the model $\mathrm{f}$ as a black box.
-However, to discuss how to impose Equation 18, we need to think about the specifics of how such models are trained and how Gaussian distributions are then obtained.
-In linearised Laplace approximations and SNGP, a neural network $\bm{f}\colon\mathcal{X}\to\mathbb{R}^{C}$ is trained for the mean of $\mathrm{f}$, while the covariance is obtained post-hoc.
-Therefore, to enforce Equation 18 on the training data distribution, we can train for $\sum_{c=1}^{C}\varphi(f_{c}(x_{n}))=1$ for all $1\leq n\leq N$.
-This can be achieved, for example, by adding an appropriate regulariser to the loss. Namely,
-$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ (28)
-=~$\quad=\tilde{\mathcal{L}}((x_{n},c_{n})_{n=1}^{N})+\lambda\sum_{n=1}^{N}\left(\sum_{c=1}^{C}\varphi(f_{c}(x_{n}))-1\right)^{2}$
-for some tuned regularisation constant $\lambda$, where $\tilde{\mathcal{L}}$ is the original unregularised loss.
-In contrast to the previously mentioned methods, HET directly trains not only for the mean $\bm{f}$ but also for the covariance matrix $\bm{\Sigma}$ of $\mathrm{f}$ through the predictives.
-Thus, in this case, we can regularise with the means of the Gaussian pushforwards through $\varphi$:
-$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ =~$=\tilde{\mathcal{L}}((x_{n},c_{n})_{n=1}^{N})$ (29)
-$+\lambda\sum_{n=1}^{N}\left(\sum_{c=1}^{C}\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}[Q_{c}]-1\right)^{2}.$
-In Section 6, we will see that training with such a learning objective does not always outperform vanilla CE training.
-For $\varphi=\exp$, for instance, we observe that the regulariser can take on extreme values, dominating the data term, slowing down training.
-We shall now see that, when $\varphi$ has range $(0,1)$, we can enforce the constraint with the learning objective in a more natural, implicit way.
-4.2 Implicit Constraint
-In the case of $\varphi=\Phi\text{ or }\rho$, $\sum_{c=1}^{C}\varphi(f_{c}(x_{n}))=1$ can be achieved more advantageously: since the range of $\Phi$ and $\rho$ is $(0,1)$, under the condition $\sum_{c=1}^{C}\varphi(f_{c}(x_{n}))=1$, we can view $\varphi(f_{c}(x_{n}))$ as estimates of $p(c\mid x_{n})$.
-This motivates the use of class-wise cross-entropy losses (class-wise CE).
-For linearised Laplace approximations and SNGP, which solely train a mean $\bm{f}$, the class-wise CE is given by
-$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ $=-\sum_{n=1}^{N}\sum_{c=1}^{C}\big{(}\delta_{c_{n},c}\log(\varphi(f_{c}(x)))$ (30)
-$\quad+(1-\delta_{c_{n},c})\log(1-\varphi(f_{c}(x))\big{)}.$
-The labels being one-hot vectors and the binary cross-entropy loss being a (negative) strictly proper scoring rule (c.f.
-Appendix I) together ensure that Equation 30 enforces the soft constraint $\sum_{c=1}^{C}\varphi\circ\mu_{c}=1$ on the training data, and hence in the training data distribution (Figure 3).
-We note that the loss Equation 30 is only used for training and not for inference or evaluation.
-Indeed, methods such as Laplace approximations and SNGP – which rely on negative log-likelihood losses to obtain an approximate Gaussian distribution – can then use the actual log-likelihood loss of the model, that is, the cross-entropy loss of $\bm{n}\circ\bm{\varphi}\circ\bm{f}$.
-Similarly, for HET, we use the class-wise CE with the pushforward means
-$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})$ $=-\sum_{n=1}^{N}\sum_{c=1}^{C}\big{(}\delta_{c_{n},c}\log(\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}[Q_{c}])$ (31)
-$+(1-\delta_{c_{n},c})\log(1-\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}[Q_{c}])\big{)}.$
-Enforcing the constraint in this implicit way is powerful since, in contrast to the regularised CE objective (Equation 28), the class-wise CE (Equation 30) fully aligns with the task.
-In Section 6 we shall see that training on class-wise CE generally outperforms softmax CE training.
-Figure 3: Evolution of the constraint on validation data for normCDF models trained on ImageNet.
+The quality of such Beta approximations is encouraging for that of the output Dirichlet approximations.
+Analysing this theoretically presents an interesting direction for future research.
+4 Training
+With the choices $\varphi=\Phi$ or $\rho$, we could still train the underlying model with a cross-entropy (CE) loss, replacing the softmax by the activation $\bm{a}=\bm{n}\circ\bm{\Phi}$ or $\bm{n}\circ\bm{\rho}$.
+However, we can instead leverage the fact that $\Phi$ and $\rho$ have values in $(0,1)$ to train using the binary cross-entropy (BCE) loss.
+The BCE loss is a valid choice of loss as it is a (negative) strictly proper scoring rule (see Appendix H).
+Moreover, it has been observed to provide advantageous training dynamics compared to the CE loss [37, 19].
+The choice of $\varphi=\Phi$ or $\rho$ not only fits our framework for closed-form predictives, but allows one to leverage such advantageous training dynamics (see experimental results in Section 6).
+For example, in linearised Laplace and SNGP, a neural network $\bm{f}\colon\mathcal{X}\to\mathbb{R}^{C}$ is trained for the mean of $\mathrm{f}$, while the covariance is obtained post-hoc.
+Thus, for these methods, we can train the mean function $\bm{f}$ with the BCE loss
+$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})=-\sum_{n=1}^{N}\sum_{c=1}^{C}\big{(}\delta_{c_{n},c}\log(\varphi(f_{c}(x)))+(1-\delta_{c_{n},c})\log(1-\varphi(f_{c}(x))\big{)}.$ (29)
 5 Computational Gains
 The runtime cost of the MC predictive approximation is $\mathcal{O}(S\cdot C+C^{3})$, and the memory cost is $\mathcal{O}(S\cdot C+C^{2})$, where $S$ is the number of samples.
-The $C^{2}$ term in memory comes from the logit covariance matrix; the $C^{3}$ term in runtime comes from the need to Cholesky-decompose such covariance matrix before sampling.
+The $C^{2}$ term in memory comes from the logit covariance matrix; the $C^{3}$ term in runtime comes from the need to Cholesky-decompose such a covariance matrix before sampling.
 For example, on a ResNet-50, sampling one thousand logit vectors from the logit-space Gaussian means an approximately 7% overhead on the forward pass.
 In embedded systems with much smaller models, this overhead is even larger.
-In contrast, the runtime and memory cost of calculating a predictive through our framework is $\mathcal{O}(C)$, which is negligible in comparison.
-Indeed, as noted at the end of Section 2.3, the constraint (Equation 18) ensures that only the mean and the diagonal of the logit covariance are needed in order to calculate the predictive exactly.
+In contrast, the runtime and memory cost of calculating our predictives is $\mathcal{O}(C)$, which is negligible in comparison.
+Indeed, the approximation Eq.
+12 uses only the mean and the diagonal of the logit covariance to calculate the predictive.
 That is, the logit quantities required to compute a predictive scale as $\mathcal{O}(C)$ instead of $\mathcal{O}(C^{2})$.
-This provides further opportunities for computational savings in the methods used to obtain the Gaussian distributions.
-In HET, for example, the logit covariance is trained with both a diagonal term and a low-rank term.
-Our framework thus allows one to drop the low-rank term without losing any expressivity in the model.
-In SNGP, the off-diagonal terms of the logit covariance are ignored (Liu et al., 2023, p.
-47).
-For unconstrained models, this incurs an approximation error as the inter-class covariance structure can encode vital information for inference.
-When Equation 18 is satisfied, however, we are guaranteed not to lose expressivity for the predictive computation by discarding the off-diagonal covariance terms, theoretically backing the approximation in SNGP.
+This can provide further opportunities for computational savings in the methods used to obtain the Gaussian distributions.
+In HET, for example, the logit covariance is trained with both a diagonal term and a low-rank term. Our framework thus allows one to drop the low-rank term (Section I.2).
 6 Experiments
 We now investigate our two research questions:
-  • Do we have to sacrifice performance for sample-free predictives?
-  • What are the effects of changing the learning objective?
-To answer the first question, we consider fixed (method, activation) pairs and verify that our methods perform on par with the Monte Carlo sampled predictives in Section 6.1.
-We train four classes of models: Softmax ($\varphi=\exp$, trained with vanilla CE), Exp ($\varphi=\exp$, trained with the explicit constraint (Section 4.1)), NormCDF ($\varphi=\Phi$, trained with the implicit constraint, (Section 4.2)) and Sigmoid ($\varphi=\rho$, trained with the implicit constraint, (Section 4.2)).
-We consider Heteroscedastic Classifiers (HET) (Collier et al., 2021), Spectral-Normalized Gaussian Processes (SNGP) (Liu et al., 2020), and last-layer Laplace approximation methods (Daxberger et al., 2021) as backbones.
-The resulting 18 (method, activation, predictive) triplets are evaluated on ImageNet-1k (Deng et al., 2009) and CIFAR-10 (Krizhevsky & Hinton, 2009) on five metrics aligning with practical needs from uncertainty estimates (Mucsányi et al., 2023).
+  1. (i)
+    Do we have to sacrifice performance for sample-free predictives? (Section 6.1)
+  2. (ii)
+    What are the effects of changing the learning objective? (Section 6.2)
+To answer the first question, we consider fixed (method, activation) pairs and verify that our methods perform on par with the Monte Carlo sampled predictives.
+In the main paper, we report results on Softmax ($\varphi=\exp$, trained with vanilla CE), NormCDF ($\varphi=\Phi$) and Sigmoid ($\varphi=\rho$) models.
+In Section K.1, we investigate the performance of Softmax models with closed-form predictives.
+Table 1: Comparison of ECE results for different predictives using a fixed Laplace backbone.
+Method Mean $\pm$ 2 std
+Sigmoid Laplace
+Closed-Form 0.0101 $\pm$ 0.0006
+MC 1000 0.0153 $\pm$ 0.0011
+MC 100 0.0153 $\pm$ 0.0012
+MC 10 0.0165 $\pm$ 0.0014
+NormCDF Laplace
+MC 10 0.0112 $\pm$ 0.0005
+Closed-Form 0.0114 $\pm$ 0.0007
+MC 100 0.0115 $\pm$ 0.0006
+MC 1000 0.0118 $\pm$ 0.0007
+We consider Heteroscedastic Classifiers (HET) [5], Spectral-Normalized Gaussian Processes (SNGP) [20], and last-layer Laplace-Approximated Networks [8] as backbones supplying logit-space Gaussian distributions.
+The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [9], CIFAR-10, and CIFAR-100 [17] on five metrics aligning with practical needs from uncertainty estimates [26].
 Further, we test the Out-of-Distribution (OOD) detection capabilities of the models on balanced mixtures of In-Distribution (ID) and OOD samples.
-For ImageNet, we treat ImageNet-C (Hendrycks & Dietterich, 2019) samples with 15 corruption types and 5 severity levels as OOD samples.
-For CIFAR-10, we use the CIFAR-10C corruptions.
-For the second question, we evaluate our analytic predictives and moment-matched Dirichlets against softmax models with approximate inference tools – Laplace Bridge (Hobbhahn et al., 2022) and Mean Field (Lu et al., 2021) approximations, as well as MC sampling, in Section 6.2.
+For ImageNet, we treat ImageNet-C [13] samples with 15 corruption types and 5 severity levels as OOD samples.
+For CIFAR-10 and CIFAR-100, we use the CIFAR-10C corruptions.
+For the second question, we evaluate our closed-form predictives and moment-matched Dirichlets against softmax models with the Laplace Bridge [14] and Mean Field [21] approximations, as well as MC sampling, in Section 6.2.
 To provide a fair comparison, we reimplement each method as simple-to-use wrappers around deterministic backbones.
-For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following Mucsányi et al.
-(2024).
+For ImageNet evaluation, we use a ResNet-50 backbone from the timm library [37] pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following Mucsányi et al.
+[27].
+For Vision Transformer [11] experiments, refer to Section K.2.
 On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs.
-We search for ideal hyperparameters and checkpoints with a ten-step Bayesian Optimization scheme (Shahriari et al., 2015) in Weights & Biases (Biewald, 2020).
+The CIFAR-100 experiments train WideResNet-28-5 models for 200 epochs.
+We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [33] in Weights & Biases [3].
 The main paper focuses on ImageNet results to highlight the scalability of our framework and only shows proper scoring results for CIFAR-10.
-For a complete overview of CIFAR-10 results, refer to Appendix K.
-In all plots of the main paper and the Appendix, the error bars are calculated over five independent runs with different seeds and show the minimum and maximum performance.
-Table 1: Comparison of ECE results for different predictives using a fixed Laplace backbone.
-Method Mean Std
+For a complete overview of CIFAR-10 and CIFAR-100 results, refer to Appendix J and Section K.3, respectively.
+In all plots of the main paper and the Appendix, the error bars are calculated over five independent runs with different seeds and show the minimum and maximum metric value.
+Table 2: Comparison of OOD AUROC results on ImageNet using severity level one for different uncertainty estimators on the three best-performing (method, activation) pairs: Sigmoid Laplace, NormCDF HET, and NormCDF Laplace.
+The best estimator is derived from the moment-matched Dirichlet distributions in all three cases.
+MC estimates use 1000 samples.
+Method Mean $\pm$ 2 std
 Sigmoid Laplace
-Analytic 0.0101 0.0006
-MC 1000 0.0153 0.0011
-MC 100 0.0153 0.0012
-MC 10 0.0165 0.0014
+Dirichlet Mutual Information 0.6388 $\pm$ 0.0005
+Closed-form Predictive Entropy 0.6353 $\pm$ 0.0003
+Monte Carlo Expected Entropy 0.6321 $\pm$ 0.0001
+NormCDF HET
+Dirichlet Expected Entropy 0.6353 $\pm$ 0.0021
+Monte Carlo Expected Entropy 0.6281 $\pm$ 0.0020
+Closed-form Predictive Entropy 0.6277 $\pm$ 0.0021
 NormCDF Laplace
-MC 10 0.0112 0.0005
-Analytic 0.0114 0.0007
-MC 100 0.0115 0.0006
-MC 1000 0.0118 0.0007
+Dirichlet Expected Entropy 0.6321 $\pm$ 0.0009
+Monte Carlo Predictive Entropy 0.6296 $\pm$ 0.0012
+Closed-Form Predictive Entropy 0.6296 $\pm$ 0.0012
 6.1 Quality of Sample-Free Predictives
-In this section, we investigate our first research question: whether there is a price to pay for sample-free predictives and second-order distributions.
-To this end, we take the two best-performing methods on the Expected Calibration Error (ECE) metric (Naeini et al., 2015), NormCDF and Sigmoid Laplace (c.f.
-Figure 6) and evaluate their per-predictive performance.
-As Table 1 shows, the analytic predictive is always on par with the (unbiased) MC estimates on ImageNet while being cheaper.
-This empirical observation supports our theoretical claim that when the constraint is satisfied, the analytic predictives are exact.
-Refer to Appendix L for an evaluation of the moment-matched Dirichlet distributions.
+In this section, we investigate whether there is a price to pay for sample-free predictives and second-order distributions.
+To this end, we take the two best-performing methods on the Expected Calibration Error (ECE) metric [29]—NormCDF and Sigmoid Laplace (see Fig.
+4(a))—and evaluate their per-predictive performance.
+As Table 1 shows, the closed-form predictive is always on par with the (unbiased) MC estimates on ImageNet while being cheaper.
+Section K.4 reports the KL divergence of different approximations to the ‘true’ predictive (estimated using 10,000 MC samples) on ImageNet.
+Table 2 shows that the moment-matched second-order Dirichlet distributions consistently outperform both MC and sample-free predictives on the OOD detection task, showcasing the practical utility of second-order distributional knowledge.
+(a) Closed-form predictives ( ) are either on par with or outperform Softmax ( ) across all methods.
+(b) Closed-form predictives ( ) consistently outperform Softmax ( ) across all methods.
+Figure 3: ImageNet (left) and CIFAR-10 (right) test negative log likelihood results.
 6.2 Effects of Changing the Learning Objective
-The previous section shows that for an already trained model that satisfies the constraint (Equation 18), our analytic predictives always perform on par with MC sampling while being more efficient.
-However, our proposed objectives (Sections 4.1 and 4.2) change the training dynamics of models.
-Therefore, in this section, we showcase the performance of methods equipped with our analytic (i.e., sample-free) predictives and learning objectives against softmax models.
-For the latter, we use the best-performing predictive and estimator (excluding ours).
-See Appendices F and G for an overview of predictives and estimators.
-We employ our methods with the analytic predictives and second-order Dirichlet distributions to demonstrate their competitive nature while being more efficient than MC sampling.
-Figure 4: ImageNet log probability proper scoring results. Analytic predictives with implicit ( ) constraints are either on par with or outperform Softmax ( ) on all methods.
-Figure 5: CIFAR-10 log probability proper scoring results. Analytic predictives with implicit ( ) constraints consistently outperform Softmax ( ) on all methods.
-We first evaluate how calibrated the models are using the log probability scoring rule (Gneiting & Raftery, 2007) and the Expected Calibration Error (ECE) metric (Naeini et al., 2015).
-Figure 5 shows that on CIFAR-10, the score of our analytic predictives with implicit constraints (Sigmoid, NormCDF) are consistently better than the corresponding Softmax results for all methods.
-On the large-scale ImageNet dataset, Sigmoid and NormCDF either outperform or are on par with Softmax.
-Figure 6: ImageNet ECE results.
-Analytic predictives with implicit constraints ( ) outperform Softmax ( ) across all methods.
-Explicit constraints ( ) work best on HET but do not benefit the other two methods.
-Note the restricted $y$-limits for readability.
-The ECE requires the models’ confidence to match their accuracy.
-Figure 6 shows that on ImageNet, our analytic predictives with implicit constraints (Sigmoid, NormCDF) have a clear advantage over Softmax across all methods.
-Explicit constraints benefit SNGP and HET but not the post-hoc Laplace method.
-Figure 7: ImageNet log probability proper scoring results for the binary correctness prediction task. Analytic predictives with implicit ( ) constraints consistently outperform Softmax ( ).
+The previous section shows that for models trained with Sigmoid or NormCDF, our closed-form predictives always perform on par with MC sampling while being more efficient.
+However, our proposed objective (see Eq.
+29) changes the training dynamics and, subsequently, the performance of models.
+In this section, we showcase the performance of methods equipped with our closed-form predictives and learning objectives against Softmax models.
+For the latter, we use the best-performing predictive and estimator (see Appendices E and F for an overview of predictive approximations and estimators).
+We employ our methods with the closed-form predictives and second-order Dirichlet distributions to demonstrate their competitive nature while being more efficient than MC sampling.
+We first evaluate how calibrated the models are using the test NLL [12] and Expected Calibration Error (ECE) [29] metrics.
+Fig.
+3(b) shows that on CIFAR-10, the score of our closed-form predictives (Sigmoid, NormCDF) is consistently better than the corresponding Softmax results for all methods.
+On the large-scale ImageNet dataset, Sigmoid and NormCDF predictives either outperform or are on par with Softmax (see Fig.
+3(a)).
+(a) Closed-form predictives ( ) are more calibrated than Softmax ( ) across all methods. Note the restricted $y$-limits for readability.
+(b) Our closed-form predictives ( ) consistently outperform Softmax ( ) on predicting their own correctness.
+Figure 4: ImageNet ECE (left) and log probability proper scoring results for the binary correctness prediction task (right).
+The ECE metric requires the models’ confidence (i.e., maximum class probability) to match their accuracy.
+Fig.
+4(a) shows that, on ImageNet, our closed-form predictives have a clear advantage over Softmax across all methods.
 Next, we turn to the correctness prediction task: whether the models can predict the correctness of their own predictions.
-We consider correctness estimators $\tilde{c}(x)\in[0,1]$ for inputs $x\in\mathcal{X}$ derived from the predictives.
-Framed as a binary prediction task, the goal of these estimators is to predict the probability of the predicted class’ correctness.
-We then measure the binary log probability score of $\tilde{c}(x)$.
-Figure 7 shows that our analytic predictives with implicit constraints outperform all Softmax predictives across all methods.
-Figure 8: Analytic predictives do not sacrifice accuracy.
-ImageNet validation accuracies.
-Analytic predictives with implicit ( ) or explicit ( ) constraints either outperform or are on par with Softmax ( ) across all methods.
-Analytic predictives do not sacrifice accuracy.
-Figure 8 evidences this claim on ImageNet: our analytic predictives either outperform or are on par with Softmax predictives.
-The most accurate methods are NormCDF and Sigmoid SNGP.
+We consider correctness estimators $t(x):=\max_{1\leq c\leq C}\mathbb{E}_{\mathbf{P}\sim\mathrm{h}(x)}\left[P_{c}\right]\in[0,1]$ for inputs $x\in\mathcal{X}$ derived from the predictive mean $\mathbb{E}_{\mathbf{P}\sim\mathrm{h}(x)}\left[\mathbf{P}\right]$.
+Framed as a binary prediction task, the goal of these estimators is to predict the probability of the predicted class’ correctness, defined as $v(x)=[\hat{y}(x)=y]\in\{0,1\}$ with $\hat{y}(x):=\arg\max_{1\leq c\leq C}\mathbb{E}_{\mathbf{P}\sim\mathrm{h}(x)}\left[P_{c}\right]\in\{1,\dots,C\}$.
+We then measure the binary log probability score of $\tilde{t}(x)$, given by $v(x)\log t(x)+(1-v(x))\log(1-t(x))$.
+Fig.
+4(b) shows that our closed-form predictives outperform all Softmax predictives across all methods.
+(a) Closed-form predictives do not sacrifice accuracy. Closed-form predictives ( ) either outperform or are on par with Softmax ( ) across all methods.
+(b) Closed-form predictives yield superior OOD detection performance. Across all methods, the best-performing predictive is closed-form ( ).
+Figure 5: ImageNet validation accuracy (left) and OOD detection AUROC results for ImageNet-C transforms of severity level one (right).
+Closed-form predictives do not sacrifice accuracy.
+Fig.
+5(a) evidences this claim on ImageNet: our closed-form predictives either outperform or are on par with Softmax predictives.
 These results are in line with the findings of Wightman et al.
-(2021), who also recommend training with a per-class binary cross-entropy loss using the sigmoid activation function.
-Figure 9: ImageNet OOD detection AUROC results for severity level one. Across all methods, the best-performing predictive is analytic with an implicit constraint ( ).
-Finally, we consider another binary prediction task, where a general uncertainty estimator $u(x)\in\mathbb{R}$ (derived from predictives or second-order Dirichlet distributions) is tasked to separate ID and OOD samples from a balanced mixture.
+[38], who also recommend training with a per-class binary cross-entropy loss using the sigmoid activation function.
+Section K.5 shows that the BCE loss already yields performance gains in the model backbones (without the logit-space Gaussians).
+Finally, we consider another binary prediction task, where a general uncertainty estimator $u(x)\in\mathbb{R}$ (derived from predictives or second-order Dirichlet distributions) is tasked to separate ID and OOD samples from a balanced mixture thereof.
 As the uncertainty estimator can take on any real value, we measure the Area Under the Receiver Operating Characteristic curve (AUROC), which quantifies the separability of ID and OOD samples w.r.t.
 the uncertainty estimator.
-As OOD inputs, we consider corrupted ImageNet-C (Hendrycks & Dietterich, 2019) samples.
-Figure 9 shows that our analytic predictives with implicit constraints (Sigmoid, NormCDF) outperform Softmax across all methods, even though separating ID and OOD samples does not require a fine-grained representation of uncertainty, unlike the ECE or proper scoring rules.
-Importantly, these analytic predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5).
-See Appendix M for the other severity levels.
+As OOD inputs, we consider corrupted ImageNet-C [13] samples.
+Fig.
+5(b) shows that our closed-form predictives outperform Softmax across all methods.
+Importantly, these closed-form predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5).
+See Section K.6 for the other severity levels.
 7 Conclusion and Limitations
-We developed a framework that allows to obtain predictives and other quantities of interest from logit space Gaussian distributions analytically.
-Our experimental results suggest the ubiquitous softmax activation should be replaced by normCDF or sigmoid.
-We proposed to approximate the Gaussian pushforwards by Dirichlet distributions which cannot encode correlations between classes.
-Leveraging a more expressive yet tractable family of distributions on the simplex that can achieve this presents an interesting area of future research.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning.
-There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
+We developed a framework that allows for obtaining predictives and other quantities of interest from logit space Gaussian distributions in closed form.
+Our experimental results suggest that the ubiquitous softmax activation should be replaced by normCDF or sigmoid for uncertainty quantification tasks.
+A limitation is that our approximate predictives and approximate Dirichlet distributions do not encode correlations between classes.
+While our theoretical analysis suggests strong bounds on the approximations, leveraging more expressive yet tractable second-order distributions on the simplex that can model correlations presents an interesting area of future research.
 Acknowledgments
@@
 Appendix A Gaussian Integral Derivations
-In this appendix, we derive the closed-form formula for the mean of Gaussian pushforwards through exp (Equation 8) and normCDF (Equation 14), as well as the approximations for pushforwards through sigmoid (Equation 16) and softmax.
+In this appendix, we derive the closed-form formula for the mean of Gaussian pushforwards through exp (Eq.
+8) and normCDF (Eq.
+13), as well as the approximations for pushforwards through sigmoid (Eq.
+15) and softmax.
 A.1 Gaussian Exp Integral
@@
 A.2 Gaussian NormCDF Integral
-Here, we derive the classical normCDF Gaussian integration formula (Owen, 1980, Eq. 10,010.8).
+Here, we derive the classical normCDF Gaussian integration formula [30, Eq. 10,010.8].
 For $\lambda>0$, $Z\sim\mathcal{N}(0,1)$ and $Y\sim\mathcal{N}(\mu,\sigma^{2})$ with $Y$ and $Z$ independent,
@@
 Taylor expanding $\rho$ to first order about 00,
-$\rho(y)$ $=\frac{1}{2}+\frac{1}{4}y+o(y)$ (34)
+$\rho(y)$ $=\frac{1}{2}+\frac{1}{4}y+o(y),$ (32)
 $\Phi(y)$ $=\frac{1}{2}+\frac{1}{\sqrt{2\pi}}y+o(y).$
+as $y\to 0$.
 Hence matching $\rho$ and $\Phi$ to first order we get the approximation $\rho(y)\approx\Phi\left(\sqrt{\frac{\pi}{8}}y\right)$.
-So using Equation 33 we derive the probit approximation (Spiegelhalter & Lauritzen, 1990; MacKay, 1992)
+So using Eq.
+31 we derive the probit approximation [34, 23]
 $\int_{\mathbb{R}}\rho(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)\stackrel{{\scriptstyle(1)}}{{\approx}}\int_{\mathbb{R}}\Phi\left(\sqrt{\frac{\pi}{8}}y\right)\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{\frac{8}{\pi}+\sigma^{2}}}\right)\stackrel{{\scriptstyle(2)}}{{\approx}}\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right).$ (33)
@@
 A.4 Mean Field Approximation for Gaussian Softmax Integral
-For $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, the mean field approximation to the Gaussian softmax integral (Lu et al., 2021) is obtained as follows
+For $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, the mean field approximation to the Gaussian softmax integral [21] is obtained as follows
 $\mathbb{E}_{\bm{Y}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\operatorname{softmax}_{c}\bm{Y}]$ $=\mathbb{E}\left[\left(2-C+\sum_{c^{\prime}\neq c}\rho(Y_{c}-Y_{c^{\prime}})^{-1}\right)^{-1}\right]$ (34)
@@
 $\mathbb{E}_{\bm{Y}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\operatorname{\mathbf{softmax}}\bm{Y}]\approx\operatorname{\mathbf{softmax}}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right).$ (35)
-(1) is the mean field approximation, and (3) uses the probit approximation Equation 35. (Lu et al., 2021) provides two other variants of this approximation with other choices of approximation (2).
+(1) is the mean field approximation, and (3) uses the probit approximation Eq. 33. [21] provides two other variants of this approximation with other choices of approximation (2).
 Appendix B Comparison with the Multinomial Probit Model
-In this appendix, we show that a model whose output activation is a composition of an element-wise normCDF activation $\bm{\Phi}$ and a normalisation $\bm{n}$ is distinct from the classical multinomial probit model (Daganzo, 1979).
+In this appendix, we show that a model whose output activation is a composition of an element-wise normCDF activation $\bm{\Phi}$ and a normalisation $\bm{n}$ is distinct from the classical multinomial probit model [6].
 Given a logit $\bm{y}\in\mathbb{R}^{C}$, the multinomial probit model sets
@@
 $p(c\mid y)=\frac{p(y_{c}+\epsilon_{c}>0)}{\sum_{c^{\prime}=1}^{C}p(y_{c^{\prime}}+\epsilon_{c^{\prime}}>0)}=\frac{\Phi(y_{c})}{\sum_{c^{\prime}=1}^{C}\Phi(y_{c}^{\prime})}.$ (38)
-In the case $C=2$, the multinomial probit model Equation 39 outputs closed form probabilities.
-This allows us to construct an explicit counterexample to the equivalence of the two models Equation 39 and Equation 40:
+In the case $C=2$, the multinomial probit model Eq. 37 outputs closed-form probabilities. This allows us to construct an explicit counterexample to the equivalence of the two models Eq. 37 and Eq. 38:
 $p(y_{1}+\epsilon_{1}>y_{2}+\epsilon_{2})$ $=p\left(\frac{y_{1}-y_{2}}{2}+\frac{\epsilon_{1}-\epsilon_{2}}{2}>0\right)=\Phi\left(\frac{y_{1}-y_{2}}{2}\right)\neq\frac{\Phi(y_{1})}{\Phi(y_{1})+\Phi(y_{2})}.$ (39)
-Appendix C Analysis of the Predictive Approximations’ Quality
-In this appendix, we provide theoretical analyses of the quality of various predictive approximations.
-This complements the empirical analyses taken, for instance, in the synthetic experiment (Figure 1) or in (Daunizeau, 2017).
-Due to its information-theoretic interpretation, a natural metric to equip the probability simplex $\Delta^{C-1}$ with is the KL divergence
+Appendix C Theoretical Analyses
+In this appendix, we provide (formal and informal) theoretical analyses of the quality of various predictive approximations.
+This complements the empirical analyses, for instance in the synthetic experiment (Fig.
+1) or in [7].
+Due to its information-theoretic interpretation, a natural divergence to equip the probability simplex $\Delta^{C-1}$ with is the Kullback-Leibler (KL) divergence
 $\operatorname{D_{KL}}(\bm{p},\bm{q})=\sum_{c=1}^{C}p_{c}(\log p_{c}-\log q_{c})$ (40)
@@
 $\bm{a}=\bm{n}\circ\bm{\varphi}$).
-C.1 Analysis of Monte Carlo Approximations
+C.1 Informal Analysis of Monte Carlo Approximations
 An $N$ sample Monte Carlo estimate is defined as
@@
 Thus, for a fair assessment of the quality of such an estimate in terms of the number of classes, one should consider MC estimates $\bm{\hat{P}}^{\lceil S/C\rceil}$.
-We now give an informal theoretical argument for the linear growth of the KL divergence between $\bm{p}$ and $\bm{\hat{P}}^{\lceil S/C\rceil}$ in terms of $C$, under the distributional conditions of the synthetic experiment (Figure 1).
-Taylor expanding Equation 42 about $\bm{p}$ to second order we obtain
+We now give an informal theoretical argument for the linear growth of the KL divergence between $\bm{p}$ and $\bm{\hat{P}}^{\lceil S/C\rceil}$ in terms of $C$, under the distributional conditions of the synthetic experiment (Fig.
+1).
+Taylor expanding Eq. 40 about $\bm{p}$ to second order we obtain
 $\operatorname{D_{KL}}(\bm{p},\bm{q})$ $=\sum_{c=1}^{C}p_{c}(\log p_{c}-\log q_{c})$ (42)
-$=\sum_{c=1}^{C}p_{c}\sum_{k=1}^{\infty}\frac{(p_{c}-q_{c})^{k}}{kp_{c}^{k}}$
-$=\underbrace{\sum_{c=1}^{C}p_{c}}_{=1}-\underbrace{\sum_{c=1}^{C}q_{c}}_{=1}+\sum_{c=1}^{C}\sum_{k=2}^{\infty}\frac{(p_{c}-q_{c})^{k}}{kp_{c}^{k-1}}$
-$\stackrel{{\scriptstyle(1)}}{{\approx}}\sum_{c=1}^{C}\frac{(p_{c}-q_{c})^{2}}{2p_{c}}$
+$\stackrel{{\scriptstyle(1)}}{{\approx}}\sum_{c=1}^{C}p_{c}\left(1-\frac{p_{c}}{q_{c}}+\frac{(p_{c}-q_{c})^{2}}{2p_{c}^{2}}\right)$
+$=\underbrace{\sum_{c=1}^{C}p_{c}}_{=1}-\underbrace{\sum_{c=1}^{C}q_{c}}_{=1}+\sum_{c=1}^{C}\frac{(p_{c}-q_{c})^{2}}{2p_{c}}$
+$=\sum_{c=1}^{C}\frac{(p_{c}-q_{c})^{2}}{2p_{c}}$
 $\stackrel{{\scriptstyle(2)}}{{\approx}}\frac{C}{2}\|\bm{p}-\bm{q}\|_{2}^{2}$
 where approximation (1) assumes $\|\bm{p}-\bm{q}\|_{2}$ is small, and (2) assumes $p_{c}\approx 1/C$.
-In the synthetic experiment (Figure 1), the logit class-wise means $\mu_{c}$ and variances $\sigma_{c}^{2}$ are sampled in an i.i.d.
+In the synthetic experiment (Fig.
+1), the logit class-wise means $\mu_{c}$ and variances $\sigma_{c}^{2}$ are sampled in an i.i.d.
 way.
-Let $\bm{Y}\sim N(\bm{\mu},\bm{\Sigma})$ with $\bm{\Sigma}:=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, $\bm{Q}=\bm{\varphi}(\bm{Y})$ the unnormalised ‘probabilities’ and $\bm{P}:=\bm{Q}/\sum_{c=1}^{C}Q_{c}$ the probabilities.
+Let $\bm{Q}\sim\mathcal{N}(\bm{\mu},\operatorname{\mathbf{diag}}(\bm{\sigma}^{2}))$ the unnormalised ‘probabilities’ and $\bm{P}:=\bm{Q}/\sum_{c=1}^{C}Q_{c}$ the probabilities.
 We have
@@
 $\mathbb{E}\big{[}\|\bm{p}-\bm{\hat{P}}^{\lceil S/C\rceil}\|^{2}_{2}\big{]}=\frac{1}{\lceil S/C\rceil}\mathbb{E}\left[\|\bm{p}-\bm{P}\|^{2}_{2}\right]\approx\frac{\mathrm{Var}[Q_{1}]}{S\mathbb{E}[Q_{1}]^{2}}.$ (45)
-Plugging this into Equation 44 we get
+Plugging this into Eq. 42 we get
 $\mathbb{E}[\operatorname{D_{KL}}(\bm{p},\bm{\hat{P}}^{\lceil S/C\rceil})]\approx\frac{\mathrm{Var}[Q_{1}]}{2\mathbb{E}[Q_{1}]^{2}}\cdot\frac{C}{S}$ (46)
-which grows linearly with the number of classes, as observed in Figure 1.
-C.2 Analysis of the Probit Approximation for Sigmoid Models
-Recall the probit approximation to the Gaussian sigmoid integral (Section A.3)
-$q:=\int_{\mathbb{R}}\rho(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)=:\hat{q}.$ (49)
-In the multiclass setting, for $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, define
-$\bm{q}:=\mathbb{E}$ [${}_{\bm{Q}\sim\bm{\rho}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{Q}]$ (50)
-$\bm{\hat{q}}:=\bm{\rho}$ $\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)$
-$\bm{p}:=\mathbb{E}_{\bm{P}\sim\bm{n}_{*}\bm{\rho}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}$ $[\bm{P}]=\mathbb{E}_{\bm{Q}\sim\bm{\rho}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}\left[\frac{\bm{Q}}{\sum_{c=1}^{C}Q_{c}}\right]$
-$\bm{\hat{p}}:=\frac{\bm{\hat{q}}}{\sum_{c=1}^{C}\hat{q}_{c}}$ $=\frac{\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)}{\sum_{c=1}^{C}\rho\left(\frac{\mu_{c}}{\sqrt{1+\frac{\pi}{8}\sigma^{2}_{c}}}\right)}$
-We assume that, e.g. through training, the constraint Equation 18 is perfectly satisfied, such that $\bm{p}=\bm{q}/\sum_{c=1}^{C}q_{c}$. Then we obtain
+which grows linearly with the number of classes, as observed in Fig. 1.
+C.2 Analysis of the Closed-Form Approximations (Theorem 2.1)
+In this section we prove Theorem 2.1 and discuss whether its underlying assumptions are fulfilled in practice. We start with a lemma.
+Lemma C.1.
+Let $(X_{n})_{n\geq 1},(Y_{n})_{n\geq 1}$ be sequences of $(0,M)$-valued random variables for some $M>0$.
+Suppose $\sup_{n}\mathbb{E}[Y_{n}^{-8}]<\infty$ and that $\mathrm{Var}(Y_{n})\to 0$ as $n\to\infty$.
+Then
+$\left|\mathbb{E}\left[\frac{X_{n}}{Y_{n}}\right]-\frac{\mathbb{E}[X_{n}]}{\mathbb{E}[Y_{n}]}\right|=O(\mathrm{Var}(Y_{n})^{1/2})$ (47)
+as $n\to\infty$.
+Proof.
+Let $X,Y$ be $(0,M)$-valued random variables. We take a bivariate Taylor expansion of $\frac{X}{Y}$ for the random variables $X$ and $Y$ around $\mathbb{E}[X]$ and $\mathbb{E}[Y]$ respectively:
+$\frac{X}{Y}=$ $\frac{\mathbb{E}[X]}{\mathbb{E}[Y]}+\frac{1}{\mathbb{E}[Y]}(X-\mathbb{E}[X])-\frac{\mathbb{E}[X]}{\mathbb{E}[Y]^{2}}(Y-\mathbb{E}[Y])$ (48)
+$-\frac{1}{\xi(Y)^{2}}(X-\mathbb{E}[X])(Y-\mathbb{E}[Y])+\frac{\eta(X)}{\xi(Y)^{3}}(Y-\mathbb{E}[Y])^{2}$
+where $\eta(X)\in[\mathbb{E}[X],X]\text{ or }[X,\mathbb{E}[X]]$ and $\xi(Y)\in[\mathbb{E}[Y],Y]\text{ or }[Y,\mathbb{E}[Y]]$, using the Lagrange form of the remainder.
+So by the Cauchy-Schwarz inequality,
+$\left|\mathbb{E}\left[\frac{X}{Y}\right]-\frac{\mathbb{E}[X]}{\mathbb{E}[Y]}\right|$ $=\left|-\mathbb{E}\left[\frac{1}{\xi(Y)^{2}}(X-\mathbb{E}[X])(Y-\mathbb{E}[Y])\right]+\mathbb{E}\left[\frac{\eta(X)}{\xi(Y)^{3}}(Y-\mathbb{E}[Y])^{2}\right]\right|$ (49)
+$\leq\mathbb{E}\left[\frac{1}{\xi(Y)^{8}}\right]^{1/4}\mathbb{E}[(X-\mathbb{E}[X])^{4}]^{1/4}\mathbb{E}[(Y-\mathbb{E}[Y])^{2}]^{1/2}$
+$\quad+\mathbb{E}\left[\frac{\eta(X)^{2}}{\xi(Y)^{6}}\right]^{1/2}\mathbb{E}[(Y-\mathbb{E}[Y])^{4}]^{1/2}.$
+Note that
+$\eta(X_{n})$ $\leq M,$ (50)
+$\sup_{n}\mathbb{E}\left[\frac{1}{\xi(Y_{n})^{k}}\right]$ $\leq\sup_{n}\mathbb{E}\left[\frac{1}{\min(Y_{n}^{k},\mathbb{E}[Y_{n}]^{k})}\right]\leq\sup_{n}\mathbb{E}\left[\frac{1}{Y_{n}^{k}}+\frac{1}{\mathbb{E}[Y_{n}]^{k}}\right]<\infty\text{ for }k\in\{6,8\},$
+$\mathbb{E}[(X_{n}-\mathbb{E}[X_{n}])^{4}]$ $\leq M^{4},$
+$\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{4}]$ $\leq M^{2}\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{2}].$
+So from Eq. 49 we obtain
+$\left|\mathbb{E}\left[\frac{X_{n}}{Y_{n}}\right]-\frac{\mathbb{E}[X_{n}]}{\mathbb{E}[Y_{n}]}\right|=O(\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{2}]^{1/2})=O(\mathrm{Var}(Y_{n})^{1/2})$ (51)
+as $n\to\infty$. ∎
+Now recall our notation
+$q=q(\mu,\sigma^{2}):=\mathbb{E}_{Q\sim\rho_{*}\mathcal{N}(\mu,\sigma^{2})}[Q]\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)=:\hat{q}(\mu,\sigma^{2})=\hat{q}$ (52)
+where $\hat{q}$ is the probit approximation (Section A.3).
+Moreover we write $q_{c}:=q(\mu_{c},\sigma_{c}^{2})$, $\hat{q}_{c}:=\hat{q}(\mu_{c},\sigma_{c}^{2})$, $\bm{p}$ for the true predictive, $\bm{\hat{p}}$ for our approximate predictive, and in addition $\bm{\tilde{p}}$ for the approximate predictive using the exact one dimensional integrals, i.e.
+$\bm{p}:=\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}],\quad\bm{\hat{p}}:=\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}},\quad\bm{\tilde{p}}:=\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}.$ (53)
+We now restate Theorem 2.1 with an explicit expression for $M(\mathcal{K})$:
 Theorem C.2.
-Let $\mathcal{K}\subset\mathbb{R}\times\mathbb{R}_{>0}$ compact.
-Using the compactness of $\mathcal{K}$ define $\delta(\mathcal{K}):=\max_{(\mu,\sigma^{2})\in\mathcal{K}}|q-\hat{q}|$ and $u(\mathcal{K}):=\min_{(\mu,\sigma^{2})\in\mathcal{K}}q>0$, where $q$ and $\hat{q}$ are as in Equation 49.
-Then if $(\mu_{c},\sigma_{c}^{2})\in\mathcal{K}$ for all $c$ we have
-$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})\leq\log\left(\frac{u+\delta}{u-\delta}\right).$ (51)
-Remark C.2.
-The bound Equation 51 is of practical value as
+Suppose the means and variances $(\mu_{c},\sigma_{c}^{2})$ lie in some compact set $\mathcal{K}\subset\mathbb{R}\times[0,\infty)$ for each class $c$. Using the compactness of $\mathcal{K}$, define
+  • $\delta(\mathcal{K}):=\sup_{(\mu,\sigma^{2})\in\mathcal{K}}(\hat{q}-q)$,
+  • $u(\mathcal{K}):=\inf_{(\mu,\sigma^{2})\in\mathcal{K}}q>0$,
+  • $\Delta(\mathcal{K}):=\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}$.
+If $\Delta(\mathcal{K})>1$ then
+$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})\leq\log\left(\frac{1+\delta/u}{1-\Delta}\right)+O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$ (54)
+as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$
+Proof.
+We have
+$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})=\sum_{c=1}^{C}p_{c}(\log p_{c}-\log\hat{p}_{c})=\underbrace{\sum_{c=1}^{C}p_{c}(\log p_{c}-\log\tilde{p}_{c})}_{(1)}+\underbrace{\sum_{c=1}^{C}p_{c}(\log\tilde{p}_{c}-\log\hat{p}_{c})}_{(2)}.$ (55)
+Assuming $\Delta<1$, we can bound (2):
+$(2)$ $=\sum_{c=1}^{C}\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\left(\log\left(\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)-\log\left(\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}}\right)\right)$ (56)
+$=\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(\frac{\hat{q}_{c}}{q_{c}}\right)+\log\left(\frac{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)\right)$
+$=\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(1-\frac{q_{c}-\hat{q}_{c}}{q_{c}}\right)+\log\left(1+\frac{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}-\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)\right)$
+$\leq\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(1-\Delta\right)+\log\left(1+\frac{C\delta}{Cu}\right)\right)$
+$=\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\log\left(\frac{1+\delta/u}{1-\Delta}\right)$
+$=\log\left(\frac{1+\delta/u}{1-\Delta}\right).$
+Now for (1), first note that $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$ where [^2] In the case $\varphi=\Phi$ and $Q_{c}\sim\Phi_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$, due to the fast decay of the tail of $\Phi$ we may have $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]=\infty$, making the proof strategy fail in that case.
+$Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$.
+By compactness of $\mathcal{K}$, we have in fact $\sup_{(\bm{\mu},\bm{\sigma})\in\mathcal{K}^{C}}\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$.
+Thus we can apply Lemma C.1 with $(X_{n})_{n\geq 1}$ a sequence of $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$ and $(Y_{n})_{n\geq 1}$ a sequence of $\sum_{c=1}^{C}Q_{c}$, to get
+$|p_{c}-\tilde{p}_{c}|=O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)^{1/2}\right)$ (57)
+as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$. Thus Taylor expanding each term of (1) around $p_{c}$ we get
+$(1)$ $=\sum_{c=1}^{C}p_{c}\left(\frac{p_{c}-\tilde{p}_{c}}{p_{c}}+\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega(\tilde{p}_{c})^{2}}\right)$ (58)
+$=\underbrace{\sum_{c=1}^{C}p_{c}}_{=1}-\underbrace{\sum_{c=1}^{C}\tilde{p}_{c}}_{=1}+\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega(\tilde{p}_{c})^{2}}$
+$=\frac{O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)^{1/2}\right)^{2}}{2\omega(\tilde{p}_{c})^{2}}$
+$=O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$
+as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, where $\omega_{c}(\tilde{p}_{c})\in[p_{c},\tilde{p}_{c}]\text{ or }[\tilde{p}_{c},p_{c}]$.
+Figure C.1: Scatter plots of logit mean and variance pairs $(\mu_{c},\sigma_{c}^{2})$ on real models.
+We see that these seem to remain constrained to a compact set, with no much difference in the support as we increase the number of classes by an order of magnitude (from CIFAR-10 to CIFAR-100).
+A key assumption in Theorem C.2 is the logit means and variances being restricted to a compact set.
+In Fig.
+C.1 we see empirically that this is approximately the case for HET and SNGP models, independently of the number of classes.
+Theorem C.2 is of practical value as $M(\mathcal{K}):=\log\left(\frac{1+\delta/u}{1-\Delta}\right)$ is well-behaved:
   1. 1.
-    it is independent of $C$,
+    $M(\mathcal{K})$ is independent of $C$,
   2. 2.
-    it tends to 0 as $\delta$ tends to 0.
-In other words, given knowledge of the worst case error in the approximation Equation 49 on the compact set $\mathcal{K}$, we can bound the KL divergence in terms of that error independently of the number of classes.
+    $M(\mathcal{K})\to 0$ as $\delta\to 0$.
+In other words, given knowledge of the worst case error in the approximation Eq.
+15 on the compact set $\mathcal{K}$, we can bound the KL divergence in terms of that error independently of the number of classes.
 Due to the simplicity of our assumptions, the bound remains quite raw and could be strengthened with further distributional assumptions on the means and variances.
-Proof.
-$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$ $=\sum_{c=1}^{C}p_{c}(\log p_{c}-\log\hat{p}_{c})$ (52)
-$=\sum_{c=1}^{C}\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\left(\log\left(\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)-\log\left(\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}}\right)\right)$
-$=\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(\frac{\hat{q}_{c}}{q_{c}}\right)+\log\left(\frac{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)\right)$
-$\leq\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(\frac{q_{c}-\delta}{q_{c}}\right)+\log\left(\frac{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}+C\delta}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)\right)$
-$=\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(1-\frac{\delta}{q_{c}}\right)+\log\left(1+\frac{C\delta}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)\right)$
-$\leq\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(1-\frac{\delta}{u}\right)+\log\left(1+\frac{C\delta}{Cu}\right)\right)$
-$=\log\left(\frac{u+\delta}{u-\delta}\right).$
-C.3 Analysis of the Mean Field Approximation for Softmax Models
-Recall the mean field approximation to the Gaussian softmax integral (Section A.4)
-$\bm{p}:=\int_{\mathbb{R}^{C}}\operatorname{\mathbf{softmax}}(\bm{x})\;\mathcal{N}(\bm{\mu},\bm{\Sigma})(d\bm{x})\approx\operatorname{\mathbf{softmax}}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)=:\bm{\hat{p}}.$ (53)
-This approximation is obtained through three subsequent approximations, that is, (1), (2), and (3) in Equation 36.
-Approximation (2) is not actually necessary, as it can be replaced by an exact computation, as shown in (Lu et al., 2021).
-However, this empirically does not appear to improve the quality of the overall approximation.
-We focus our attention on the main approximation (1), namely
-$\mathbb{E}\left[\left(2-C+\sum_{c^{\prime}\neq c}\rho(Y_{c}-Y_{c^{\prime}})^{-1}\right)^{-1}\right]\stackrel{{\scriptstyle(1)}}{{\approx}}\left(2-C+\sum_{c^{\prime}\neq c}\mathbb{E}[\rho(Y_{c}-Y_{c^{\prime}})]^{-1}\right)^{-1}.$ (54)
-This is exact if $\rho(Y_{c}-Y_{c^{\prime}})$ is a constant random variable for all $Y_{c}$ and $Y_{c^{\prime}}$.
-For $C>2$, this would imply that $Y_{c}$ is distributed according to a Dirac delta.
-However, this does not allow for any second-order distribution over the logits.
-We thus see that we cannot train for exactness in the mean field approximation as we do for Equation 13.
-This is a fundamental limitation of this mean field approximation.
+Finally, to obtain a meaningful bound in Theorem C.2, we needed to assume $\Delta(\mathcal{K}):=\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$.
+In Fig.
+C.2 we see that this can be assumed to hold on the compact sets on which logit means and variances tend to live (see also Fig.
+C.1).
+Figure C.2: Plot of the ‘true’ predictive $q$ (approximated with a $50000$ sample MC approximation) versus the probit approximation $\hat{q}$, as well as $\frac{q-\hat{q}}{q}$.
+We see that $\frac{q-\hat{q}}{q}$ tends to be negative, making the assumption $\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$ from Theorem C.2 easily fulfilled.
 Appendix D Moment Matching Beta Distributions
@@
 by moment matching.
-Specifically, the parameters $\bm{\alpha},\bm{\beta}\in\mathbb{R}_{>0}^{C}$ that match the moments of $\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}$ for some $\mathrm{f}\in\mathcal{G}(\mathbb{R}^{C})$ must satisfy
+Specifically, the parameters $\bm{\alpha},\bm{\beta}\in(0,\infty)^{C}$ that match the moments of $\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}$ for some $\mathrm{f}\in\mathcal{G}(\mathbb{R}^{C})$ must satisfy
 $\mathbb{E}[\bm{Q}]$ $=\frac{\bm{\alpha}}{\bm{\alpha}+\bm{\beta}},$ (60)
@@
 where all vector operations are element-wise.
-Multiplying out the denominators on the right-hand side of the equations of Equation 56, we obtain a system of two linear equations with two unknowns (for each $c$), which can be solved uniquely, yielding
+Multiplying out the denominators on the right-hand side of the equations of Eq.
+60, we obtain a system of two linear equations with two unknowns (for each $c$), which can be solved uniquely, yielding
 $\bm{\alpha}$ $:=\frac{\mathbb{E}[\bm{Q}]-\mathbb{E}[\bm{Q}^{2}]}{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}\mathbb{E}[\bm{Q}],$ (61)
@@
 D.1 Information Geometric Interpretation of the Pushforward through NormCDF
-Figure 10: Illustration of the statistical manifolds of the pushforward Gaussian distributions $\mathcal{G}(\mathbb{R})$ through the normCDF and the sigmoid respectively compared to the statistical manifold of Beta distributions $\mathcal{B}((0,1))$.
+Figure D.1: Illustration of the statistical manifolds of the pushforward Gaussian distributions $\mathcal{G}(\mathbb{R})$ through the normCDF and the sigmoid respectively compared to the statistical manifold of Beta distributions $\mathcal{B}((0,1))$.
 NormCDF, unlike the sigmoid, makes the manifolds intersect at the point $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$.
-Here, we argue that the normCDF activation is an ideal choice for approximating Gaussian pushforwards with Beta distributions by interpreting Figure 10.
-This extends the work from (MacKay, 1998), as it shows how one can make sense of the ‘right’ basis for performing Laplace approximations in the classification setting.
+Here, we argue that the normCDF activation is an ideal choice for approximating Gaussian pushforwards with Beta distributions by interpreting Fig.
+D.1.
+This extends the work from [22], as it shows how one can make sense of the ‘right’ basis for performing Laplace approximations in the classification setting.
 $\Phi_{*}\mathcal{G}(\mathbb{R})$, $\rho_{*}\mathcal{G}(\mathbb{R})$ the space of pushforwards of Gaussian distributions by normCDF and sigmoid respectively, and $\mathcal{B}((0,1))$, the space of Beta distributions, are statistical manifolds naturally equipped with Riemannian metrics, that is their respective Fisher information metrics.
@@
 The key observation is that $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$ are isometric to $\mathcal{G}(\mathbb{R})$.
-This is because $\Phi$ and $\rho$ are diffeomorphisms, so in particular sufficient statistics, and the Fisher information metric is invariant under sufficient statistics (Ay et al., 2017, Section 5.1.3).
+This is because $\Phi$ and $\rho$ are diffeomorphisms, so in particular sufficient statistics, and the Fisher information metric is invariant under sufficient statistics [2, Section 5.1.4].
 Visually, this means that $\varphi_{*}\mathcal{G}(\mathbb{R})$ has the same shape irrespectively of the diffeomorphism activation function $\varphi$.
@@
 Now $\mathcal{G}(\mathbb{R})$, and hence $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$, is isometric to the hyperbolic plane Ay et al.
-(2017, Example 3.1).
-We can embed this isometrically into Minkowski space with the classical hyperboloid model of the hyperbolic plane (Reynolds, 1993),
+[2, Example 3.1].
+We can embed this isometrically into Minkowski space with the classical hyperboloid model of the hyperbolic plane [31],
 $\mathcal{G}(\mathbb{R})\hookrightarrow\mathbb{R}^{2,1}.$ (62)
-For $\mathcal{B}((0,1))$, we use the isometric embedding from Le Brigant et al. (2021, Proposition 2):
+For $\mathcal{B}((0,1))$, we use the isometric embedding from Le Brigant et al. [18, Proposition 2]:
 $\mathcal{B}((0,1))$ $\hookrightarrow\mathbb{R}^{2,1}$ (63)
@@
 […]
-Finally, we choose our embedding Equation 58 for $\Phi_{*}\mathcal{G}(\mathbb{R})$ such that it intersects the embedding Equation 59 at a point, to highlight that the statistical manifolds $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ intersect at a point in the infinite-dimensional ambient space $\mathcal{M}((0,1))$, while $\rho_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ do not.
-Moreover, since $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$, our moment matching approximation Equation 56 is exact when $\mathrm{f}(x)$ is the standard Normal distribution.
-Appendix E Comparison with the Energy Model View of Classifiers
-Grathwohl et al. (2019) propose to view classifiers of the form Equation 2 as energy-based models by
-$p(c\mid x)$ $=\frac{\exp(f_{c}(x))}{\sum_{c^{\prime}=1}^{C}\exp(f_{c^{\prime}}(x))},$ (60)
-$p(x,c)$ $=\frac{\exp(f_{c}(x))}{Z},$
-$p(x)$ $=\frac{\sum_{c=1}^{C}\exp(f_{c}(x))}{Z}$
-where $Z=\int_{\mathcal{X}}\sum_{c=1}^{C}\exp(f_{c}(x))\;dx$. Note the consistency
-$p(c\mid x)=\frac{p(x,c)}{p(x)}.$ (61)
-Observe that the choice of $p(x,c)$ and $p(x)$ in Equation 60 is not canonical: we could have multiplied $p(x,c)$ and $p(x)$ by $\frac{h(x)}{Z^{\prime}}$ for any non-vanishing function $h$ and appropriate normalisation constant $Z^{\prime}$, and we would still obtain the consistency Equation 61.
-Grathwohl et al.
-(2019) propose to use the degree of freedom in the logits to train for Equation 60, turning discriminative models into generative models.
-Note that $\exp$ in Equation 60 can be replaced by any positive activation $\varphi$.
-In our framework, we train for $\sum_{c=1}^{C}\varphi(f_{c}(x))=s(x)$.
-More specifically, in the present work, we train for $\sum_{c=1}^{C}\varphi(f_{c}(x))=1$ (Section 4), making use of the degree of freedom in the logits to obtain exact predictives in Equation 13.
-In the formalism Equation 60, this means we are setting $x$ to be uniformly distributed.
-An interesting avenue of future research would be to combine our framework with that of Grathwohl et al.
-(2019), by e.g.
-training for $\sum_{c=1}^{C}\varphi(f_{c}(x))\propto p(x)$, and extending the framework of Grathwohl et al.
-(2019) to second-order distributions.
-Appendix F List of Predictive and Dirichlet Parameters Formulas
-F.1 Predictive Formulas
+Finally, we choose our embedding Eq.
+62 for $\Phi_{*}\mathcal{G}(\mathbb{R})$ such that it intersects the embedding Eq.
+63 at a point, to highlight that the statistical manifolds $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ intersect at a point in the infinite-dimensional ambient space $\mathcal{M}((0,1))$, while $\rho_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ do not.
+Moreover, since $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$, our moment matching approximation Eq. 60 is exact when $\mathrm{f}(x)$ is the standard Normal distribution.
+Appendix E List of Predictive and Dirichlet Parameters Formulas
+E.1 Predictive Formulas
 We gather formulas for all predictive estimators $\bm{\hat{p}}$ of the true predictive $\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}]$, $\bm{\Sigma}=\bm{\operatorname{\mathbf{diag}}}(\bm{\sigma}^{2})$, used in our experiments (Section 6).
-F.1.1 Softmax
+E.1.1 Softmax
 Monte Carlo
@@
 Mean Field (Section A.4)
-The Mean Field predictive (Lu et al., 2021) uses the following approximation for the true predictive:
+The Mean Field predictive [21] uses the following approximation for the true predictive:
 $\bm{\hat{p}}:=\bm{\operatorname{\mathbf{softmax}}}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right).$ (65)
 Laplace Bridge
-The Laplace Bridge predictive (Hobbhahn et al., 2022) approximates the true predictive as follows:
+The Laplace Bridge predictive [14] approximates the true predictive as follows:
 $\bm{\hat{p}}:=\frac{\frac{1}{\bm{\tilde{\sigma}}^{2}}\left(1-\frac{2}{C}+\frac{e^{\bm{\tilde{\mu}}}}{C^{2}}\sum_{c=1}^{C}e^{-\tilde{\mu}_{c}}\right)}{\sum_{c=1}^{C}\frac{1}{\tilde{\sigma}^{2}_{c}}\left(1-\frac{2}{C}+\frac{e^{\bm{\tilde{\mu}}}}{C^{2}}\sum_{c^{\prime}=1}^{C}e^{-\tilde{\mu}_{c^{\prime}}}\right)}$ (66)
@@
 $\bm{\tilde{\mu}}^{2}:=\sqrt{\frac{\sqrt{C/2}}{\sum_{c=1}^{C}\sigma^{2}_{c}}}\bm{\mu},\;\bm{\tilde{\sigma}^{2}}:=\frac{\sqrt{C/2}}{\sum_{c=1}^{C}\sigma^{2}_{c}}\bm{\sigma}^{2}.$ (67)
-F.1.2 Exp
-Our analytic predictive for the exp activation function (Section 2.1) is given by
+E.1.2 Exp
+Our closed-form predictive for the exp activation function (Section 2.1) is given by
 $\bm{\hat{p}}:=\frac{\bm{\exp}\left(\bm{\mu}+\frac{\bm{\sigma}^{2}}{2}\right)}{\sum_{c=1}^{C}\exp\left(\mu_{c}+\frac{\sigma^{2}_{c}}{2}\right)}.$ (68)
-F.1.3 NormCDF
-The analytic predictive for the normCDF activation function (Section 2.2) is computed as
+E.1.3 NormCDF
+The closed-form predictive for the normCDF activation function (Section 2.2) is computed as
 $\bm{\hat{p}}:=\frac{\bm{\Phi}\left(\frac{\bm{\mu}}{\sqrt{1+\bm{\sigma}^{2}}}\right)}{\sum_{c=1}^{C}\Phi\left(\frac{\mu_{c}}{\sqrt{1+\sigma^{2}_{c}}}\right)}.$ (69)
-F.1.4 Sigmoid
-For the sigmoid activation function (Section 2.2), the analytic predictive can be computed as
+E.1.4 Sigmoid
+For the sigmoid activation function (Section 2.2), the closed-form predictive can be computed as
 $\bm{\hat{p}}:=\frac{\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)}{\sum_{c=1}^{C}\rho\left(\frac{\mu_{c}}{\sqrt{1+\frac{\pi}{8}\sigma^{2}_{c}}}\right)}.$ (70)
-F.2 Dirichlet Parameters Formulas
+E.2 Dirichlet Parameters Formulas
 We now gather the formulas of the parameters $\bm{\gamma}$ for the Dirichlet approximations to the Gaussian pushforwards.
-F.2.1 Softmax
-The Laplace Bridge method (Hobbhahn et al., 2022) uses the following Dirichlet parameters:
+E.2.1 Softmax
+The Laplace Bridge method [14] uses the following Dirichlet parameters:
 $\bm{\gamma}:=\frac{1}{\bm{\tilde{\sigma}}^{2}}\left(1-\frac{2}{C}+\frac{e^{\bm{\tilde{\mu}}}}{C^{2}}\sum_{c=1}^{C}e^{-\tilde{\mu}_{c}}\right)$ (71)
@@
 $\bm{\tilde{\mu}}^{2}:=\sqrt{\frac{\sqrt{C/2}}{\sum_{c=1}^{C}\sigma^{2}_{c}}}\bm{\mu},\;\bm{\tilde{\sigma}^{2}}:=\frac{\sqrt{C/2}}{\sum_{c=1}^{C}\sigma^{2}_{c}}\bm{\sigma}^{2}.$ (72)
-F.2.2 Exp
-Exp uses the analytic parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
+E.2.2 Exp
+Exp uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
 $\bm{\gamma}:=\left(\prod_{c=1}^{C}\frac{a_{c}\cdot\max(\sum_{c^{\prime}=1}^{C}a_{c^{\prime}},1)-b_{c}}{b_{c}-a_{c}^{2}}\right)^{1/C}\frac{\bm{a}}{\sum_{c=1}^{C}a_{c}}$ (73)
@@
 $\bm{a}=\exp\left(\bm{\mu}+\frac{\bm{\sigma}^{2}}{2}\right),\;\bm{b}=\exp(2\bm{\mu}+2\bm{\sigma}^{2}).$ (74)
-F.2.3 NormCDF
-NormCDF uses the analytic parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
+E.2.3 NormCDF
+NormCDF uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
 $\bm{\gamma}:=\left(\prod_{c=1}^{C}\frac{a_{c}\cdot\max(\sum_{c^{\prime}=1}^{C}a_{c^{\prime}},1)-b_{c}}{b_{c}-a_{c}^{2}}\right)^{1/C}\frac{\bm{a}}{\sum_{c=1}^{C}a_{c}}$ (75)
@@
 $\bm{a}=\bm{\Phi}\left(\frac{\bm{\mu}}{\sqrt{1+\bm{\sigma}^{2}}}\right),\;\bm{b}=\bm{\Phi}\left(\frac{\bm{\mu}}{\sqrt{1+\bm{\sigma}^{2}}}\right)-2\bm{T}\left(\frac{\bm{\mu}}{\sqrt{1+\bm{\sigma}^{2}}},\frac{1}{\sqrt{1+2\bm{\sigma}^{2}}}\right).$ (76)
-F.2.4 Sigmoid
-Sigmoid uses the analytic parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
+E.2.4 Sigmoid
+Sigmoid uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
 $\bm{\gamma}:=\left(\prod_{c=1}^{C}\frac{a_{c}\cdot\max(\sum_{c^{\prime}=1}^{C}a_{c^{\prime}},1)-b_{c}}{b_{c}-a_{c}^{2}}\right)^{1/C}\frac{\bm{a}}{\sum_{c=1}^{C}a_{c}}$ (77)
 where
-$\bm{a}=\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right),\;\bm{b}=\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)-\frac{1}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)\left(1-\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)\right).$ (76)
-Appendix G List of Uncertainty Estimators
+$\bm{a}$ $=\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right),$ (78)
+$\bm{b}$ $=\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)-\frac{1}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)\left(1-\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)\right).$
+Appendix F List of Uncertainty Estimators
 In this section, we list the uncertainty estimators used in our experiments (Section 6).
-G.1 Predictive
+F.1 Predictive
 Given a predictive $\bm{p}$, we consider two uncertainty estimators.
@@
 $-\sum_{c=1}^{C}p_{c}\log p_{c}.$ (80)
-G.2 Monte Carlo
-Given $S$ Monte Carlo samples $\bm{\hat{p}}^{(1)},\dots,\bm{\hat{p}}^{(S)}$ with mean $\bm{\hat{p}}$, one can calculate a predictive as their average and derive the estimators in Section G.1.
+F.2 Monte Carlo
+Given $S$ Monte Carlo samples $\bm{\hat{p}}^{(1)},\dots,\bm{\hat{p}}^{(S)}$ with mean $\bm{\hat{p}}$, one can calculate a predictive as their average and derive the estimators in Section F.1.
 However, Monte Carlo samples allow one to calculate two additional estimators detailed below.
@@
 $-\sum_{c=1}^{C}{\hat{p}}_{c}\log{\hat{p}}_{c}+\frac{1}{S}\sum_{s=1}^{S}\sum_{c=1}^{C}\hat{p}^{(s)}_{c}\log\hat{p}^{(s)}_{c}.$ (82)
-G.3 Dirichlet
+F.3 Dirichlet
 Given a second-order Dirichlet distribution with parameters $\bm{\gamma}$, one can obtain the expected entropy and mutual information estimators without the need for Monte Carlo samples.
@@
 $\sum_{c=1}^{C}\frac{\gamma_{c}}{\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}}\left(-\log\gamma_{c}+\log\left(\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}\right)+\psi(\gamma_{c}+1)-\psi\left(\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}+1\right)\right).$ (84)
-Appendix H Experimental Setup
+Appendix G Experimental Setup
 This section describes our experimental setup in detail.
@@
   • Do we have to sacrifice performance for sample-free predictives?
-To answer the first question, we evaluate our analytic predictives (Sigmoid, NormCDF, Exp) and moment-matched Dirichlet distributions against softmax models equipped with approximate inference tools (Laplace Bridge (Hobbhahn et al., 2022), Mean Field (Lu et al., 2021), Monte Carlo sampling).
-We consider Heteroscedastic Classifiers (HET) (Collier et al., 2021), Spectral-Normalized Gaussian Processes (SNGP) (Liu et al., 2020), and last-layer Laplace approximation methods (Daxberger et al., 2021) as backbones (see Appendix J for details).
-The resulting 18 (method, activation, predictive) triplets are evaluated on ImageNet-1k (Deng et al., 2009) and CIFAR-10 (Krizhevsky & Hinton, 2009) on five metrics aligning with practical needs from uncertainty estimates (Mucsányi et al., 2023):
+To answer the first question, we evaluate our closed-form predictives (Sigmoid, NormCDF, Exp) and moment-matched Dirichlet distributions against softmax models equipped with approximate inference tools (Laplace Bridge [14], Mean Field [21], Monte Carlo sampling).
+We consider Heteroscedastic Classifiers (HET) [5], Spectral-Normalized Gaussian Processes (SNGP) [20], and last-layer Laplace approximation methods [8] as backbones (see Appendix I for details).
+The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [9] and CIFAR-10 [17] on five metrics aligning with practical needs from uncertainty estimates [26]:
   1. 1.
@@
     AUROC for the out-of-distribution (OOD) detection task.
-See Appendix I for details.
-For ImageNet, we treat ImageNet-C (Hendrycks & Dietterich, 2019) samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
+See Appendix H for details.
+For ImageNet, we treat ImageNet-C [13] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
 For the second question, we consider fixed (method, activation) pairs and test whether our methods perform on par with the Monte Carlo sampled predictives.
@@
 For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following Mucsányi et al.
-(2024).
-We train with the LAMB optimizer (You et al., 2019) using a batch size of $128$ and gradient accumulation across $16$ batches, resulting in an effective batch size of $2048$, following Tran et al.
-(2022).
+[27].
+We train with the LAMB optimiser [40] using a batch size of $128$ and gradient accumulation across $16$ batches, resulting in an effective batch size of $2048$, following Tran et al.
+[36].
 We further use a cosine learning rate schedule with a single warmup epoch using a warmup learning rate of $0.0001$.
@@
 During training, we keep track of the best-performing checkpoint on the validation set and load it before testing.
-We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme (Shahriari et al., 2015) in Weights & Biases (Biewald, 2020) based on the negative log-likelihood.
+We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [33] in Weights & Biases [3] based on the negative log-likelihood.
 On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs.
-The only exceptions are the SNGP models that are trained for 125 epochs (Liu et al., 2020).
+The only exceptions are the SNGP models that are trained for 125 epochs [20].
 We train with Momentum SGD using a batch size of $128$ and no gradient accumulation.
@@
 Similarly to ImageNet, we use the best-performing checkpoint in the tests and use a ten-step Bayesian Optimization scheme to select performant hyperparameters.
-Appendix I Benchmark Metrics
+The hyperparameter optimization, training, and evaluation of the methods used in this paper took $0.8$ GPU years on NVIDIA RTX 2080Ti GPUs in a university compute cluster.
+The individual runs required no more than $50$ GB of RAM and $3$ days of runtime.
+Appendix H Benchmark Metrics
 Our experiments use five tasks/metrics:
@@
 Below, we describe these metrics and their respective tasks.
-I.1 Log Probability Proper Scoring Rule for the Predictive
-First, we briefly discuss proper and strictly proper scoring rules over general probability measures based on (Mucsányi et al., 2023).
+H.1 Log Probability Proper Scoring Rule for the Predictive
+First, we briefly discuss proper and strictly proper scoring rules over general probability measures based on [26].
 Consider a function $S\colon\mathcal{Q}\times\mathcal{Y}\to\mathbb{R}$ where $\mathcal{Q}$ is a family of probability distributions over the space $\mathcal{Y}$, called the label space.
@@
 The score above is equivalent to the negative cross-entropy loss.
-I.2 Expected Calibration Error
-To set up the required quantities for the Expected Calibration Error (ECE) metric (Naeini et al., 2015), we follow the steps below, based on (Mucsányi et al., 2023).
+H.2 Expected Calibration Error
+To set up the required quantities for the Expected Calibration Error (ECE) metric [29], we follow the steps below, based on [26].
   1. 1.
@@
     $\tilde{c}_{n}\in\left(\frac{m-1}{M},\frac{m}{M}\right].$ (88)
-The Expected Calibration Error (ECE) metric (Naeini et al., 2015) is then defined as
+The Expected Calibration Error (ECE) metric [29] is then defined as
 $\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\right|$ (89)
@@
 $\mathrm{acc}(B_{m})$ $=\frac{1}{|B_{m}|}\sum_{n\in B_{m}}1\left(\hat{y}_{n}=c_{n}\right),$ (90)
-$\mathrm{conf}(B_{m})$ $=\frac{1}{|B_{m}|}\sum_{n\in B_{m}}\max_{c\in\{1,\dotsc,C\}}f_{c}(x_{n}).$ (89)
+$\mathrm{conf}(B_{m})$ $=\frac{1}{|B_{m}|}\sum_{n\in B_{m}}\max_{1\leq c\leq C}f_{c}(x_{n}).$ (91)
 Intuitively, the ECE is high when the model’s per-bin confidences match its accuracy on the bin. We use $M=15$ bins in this paper.
-I.3 Binary Log Probability Proper Scoring Rule for Correctness Prediction
+H.3 Binary Log Probability Proper Scoring Rule for Correctness Prediction
 The correctness prediction task measures the models’ ability to predict the correctness of their own predictions.
@@
 In particular, for an (input, target) pair $(x,y)$ with $x\in\mathcal{X},y\in\mathcal{Y}$, we set the correctness target to
-$\ell\equiv\ell(x,y)=1\left(\max_{c\in\{1,\dotsc,C\}}h_{c}(x)=y\right).$ (90)
+$\ell\equiv\ell(x,y)=1\left(\max_{1\leq c\leq C}h_{c}(x)=y\right).$ (92)
 Dropping the dependency on $x\in\mathcal{X}$ for brevity, the log probability score for binary targets $\ell\in\{0,1\}$ and estimators $\tilde{c}\in[0,1]$ is defined as
 $S(\tilde{c},\ell)=\begin{cases}\log c&\mathrm{if}\ \ell=1\\ \log(1-\tilde{c})&\mathrm{if}\ \ell=0\end{cases}=\ell\log\tilde{c}+(1-\ell)\log(1-\tilde{c}).$ (93)
-One can show that this is indeed a strictly proper scoring rule (Mucsányi et al., 2023).
-I.4 Accuracy
+One can show that this is indeed a strictly proper scoring rule [26].
+H.4 Accuracy
 For completeness, the accuracy of a predictive $h$ on a dataset $(x_{n},c_{n})_{n=1}^{N}$ is
 $\operatorname{acc}(h;(x_{n},c_{n})_{n=1}^{N})=\frac{1}{N}\sum_{n=1}^{N}1\left(\operatorname*{arg\,max}_{c\in\{1,\dotsc,C\}}h_{c}(x_{n})=\tilde{c}_{n}\right).$ (94)
-I.5 Area Under the Receiver Operating Characteristic Curve for Out-of-Distribution Detection
+H.5 Area Under the Receiver Operating Characteristic Curve for Out-of-Distribution Detection
 Out-of-distribution detection is another binary prediction task where a general uncertainty estimator $u(x)\in\mathbb{R}$ (derived from predictives or second-order Dirichlet distributions) is tasked to separate ID and OOD samples from a balanced mixture.
@@
 The Receiver Operating Characteristic (ROC) curve compares the following quantities:
-$\mathrm{TPR}(t)$ $=\frac{|\mathrm{TP}(t)|}{|\mathrm{TP}(t)|+|\mathrm{FN}(t)|}=\frac{|\mathrm{TP}(t)|}{|\mathrm{P}|}$
+$\mathrm{TPR}(t)$ $=\frac{|\mathrm{TP}(t)|}{|\mathrm{TP}(t)|+|\mathrm{FN}(t)|}=\frac{|\mathrm{TP}(t)|}{|\mathrm{P}|}$ (96)
 $\mathrm{FPR}(t)$ $=\frac{|\mathrm{FP}(t)|}{|\mathrm{FP}(t)|+|\mathrm{TN}(t)|}=\frac{|\mathrm{FP}(t)|}{|\mathrm{N}|}.$
@@
 The area under this curve quantifies how well the uncertainty estimator $u(x)$ can separate in-distribution and out-of-distribution inputs.
-Appendix J Benchmarked Methods
+Appendix I Benchmarked Methods
 This section describes our benchmarked methods and provides further implementation details.
-J.1 Spectral Normalized Gaussian Process
-Spectral normalized Gaussian processes (SNGP) (Liu et al., 2020) use spectral normalization of the parameter tensors for distance-awareness and a last-layer Gaussian process approximated by Fourier features to capture uncertainty.
+I.1 Spectral Normalized Gaussian Process
+Spectral normalized Gaussian processes (SNGP) [20] use spectral normalization of the parameter tensors for distance-awareness and a last-layer Gaussian process approximated by Fourier features to capture uncertainty.
 For an input $x\in\mathcal{X}$ and number of classes $C$, they predict a $C$-variate Gaussian distribution
-$\mathcal{N}\left(\bm{B}\bm{\phi}(x),\bm{\phi}(x)^{\top}\left(\bm{\Psi}^{\top}\bm{\Psi}+\bm{I}\right)^{-1}\bm{\phi}(x)\bm{I}_{C}\right),$ (93)
+$\mathcal{N}\left(\bm{B}\bm{\phi}(x),\bm{\phi}(x)^{\top}\left(\bm{\Psi}^{\top}\bm{\Psi}+\bm{I}\right)^{-1}\bm{\phi}(x)\bm{I}_{C}\right)$ (97)
 in logit space.
@@
   • $\bm{\Psi}^{\top}\bm{\Psi}$ is the (unnormalised) empirical covariance matrix of the pre-logits of the training set.
-This is calculated via accumulating the mini-batch estimates during the last epoch.
+This is calculated by accumulating the mini-batch estimates during the last epoch.
 [^3] As we use a cosine learning rate decay in all experiments, the model makes negligible changes in its pre-logit feature space in the last epoch.
@@
 Liu et al.
-(2020) claim this helps with input distance awareness.
-J.2 Heteroscedastic Classifier
-Heteroscedastic classifiers (HET) (Collier et al., 2021) construct a Gaussian distribution in the logit space to model per-input uncertainties:
+[20] claim this helps with input distance awareness.
+I.2 Heteroscedastic Classifier
+Heteroscedastic classifiers (HET) [5] construct a Gaussian distribution in the logit space to model per-input uncertainties:
 $\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x)),$ (98)
@@
 is a (positive definite) covariance matrix. Both the low-rank term $\bm{V}(x)$ and the diagonal term $\bm{d}(x)$ are calculated as a linear function of the pre-logit layer’s output.
-To learn the per-input covariance matrices from the training set, one has to construct a predictive estimate from $\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x))$ using any of the methods in Appendix F.
+To learn the per-input covariance matrices from the training set, one has to construct a predictive estimate from $\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x))$ using any of the methods in Appendix E.
 This predictive estimate is then trained using a standard cross-entropy (NLL) loss.
 HET uses a temperature parameter to scale the logits before calculating the BMA. This is chosen using a validation set.
-In Section 2.3, we show that when our constraint is satisfied, the off-diagonal terms of the covariance matrix do not matter for the predictive.
+The off-diagonal terms of the covariance matrix do not affect the approximate predictive (Eq.
+12).
 This means that, in our framework, one can discard the low-rank term $\bm{V}(x)$ and only model the diagonal term $\bm{d}(x)$ without a decrease in expressivity.
 To keep comparisons fair and use the same backbone with the same number of parameters, we also only model $\bm{d}(x)$ for softmax-based predictives.
-J.3 Laplace Approximation
-The Laplace approximation (Daxberger et al., 2021) approximates the posterior $p(\bm{\theta}\mid\mathcal{D})$ over the network parameters $\bm{\theta}$ for a Gaussian prior $p(\bm{\theta})$ and likelihood defined by the network architecture by a Gaussian.
-In its simplest form, it uses the maximum a posteriori (MAP) weights $\bm{\theta}_{\text{MAP}}\in\mathbb{R}^{P}$ as the mean and the inverse Hessian of the regularized loss over the training set ~$\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})=\mathcal{L}(\bm{\theta};\mathcal{D})+\lambda\|\bm{\theta}\|_{2}^{2}$ evaluated at the MAP as the covariance matrix:
+I.3 Laplace Approximation
+The Laplace approximation [8] approximates the posterior $p(\bm{\theta}\mid\mathcal{D})$ over the network parameters $\bm{\theta}$ for a Gaussian prior $p(\bm{\theta})$ and likelihood defined by the network architecture by a Gaussian.
+In its simplest form, it uses the maximum a posteriori (MAP) weights $\bm{\theta}_{\text{MAP}}\in\mathbb{R}^{P}$ as the mean and the inverse Hessian of the regularised loss over the training set ~$\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})=\mathcal{L}(\bm{\theta};\mathcal{D})+\lambda\|\bm{\theta}\|_{2}^{2}$ evaluated at the MAP as the covariance matrix:
 $\mathcal{N}\left(\bm{\theta}_{\text{MAP}},\left(\left.\frac{\partial^{2}\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})}{\partial\theta_{i}\partial\theta_{j}}\right|_{\bm{\theta}_{\text{MAP}}}\right)^{-1}\right)=\mathcal{N}\left(\bm{\theta}_{\text{MAP}},\left(\left.\frac{\partial^{2}\mathcal{L}(\bm{\theta};\mathcal{D})}{\partial\theta_{i}\partial\theta_{j}}\right|_{\bm{\theta}_{\text{MAP}}}+\lambda\bm{I}_{P}\right)^{-1}\right).$ (100)
-This is a locally optimal post-hoc Gaussian approximation of the true posterior $p(\bm{\theta}\mid\mathcal{D})$ based on a second-order Taylor approximation. For details, see (Tatzel et al., 2025).
+This is a locally optimal post-hoc Gaussian approximation of the true posterior $p(\bm{\theta}\mid\mathcal{D})$ based on a second-order Taylor approximation. For details, see [35].
 For deep neural networks, the Hessian matrix is often replaced with the Generalized Gauss-Newton (GGN) matrix.
-The GGN is guaranteed to be positive semidefinite even for suboptimal weights and has efficient approximation schemes, such as Kronecker-Factored Approximate Curvature (Martens & Grosse, 2015) or low-rank approximations.
+The GGN is guaranteed to be positive semidefinite even for suboptimal weights and has efficient approximation schemes, such as Kronecker-Factored Approximate Curvature [24] or low-rank approximations.
 Denoting our curvature estimate of choice by $\bm{G}$, the logit-space Gaussian is obtained by pushing forward the weight-space Gaussian measure through the linearised model around $\bm{\theta}_{\text{MAP}}$.
 For an input $x\in\mathcal{X}$, this results in
-$\mathcal{N}\left(\left(\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\right)\bm{\theta}_{\text{MAP}},\left(\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\right)\left(\bm{G}+\lambda\bm{I}_{P}\right)^{-1}\left(\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\right)^{\top}\right),$ (97)
+$\mathcal{N}\left(\bm{f}(x,\bm{\theta}_{\text{MAP}}),\left(\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\right)\left(\bm{G}+\lambda\bm{I}_{P}\right)^{-1}\left(\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\right)^{\top}\right),$ (101)
 where $\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\in\mathbb{R}^{C\times P}$ is the model Jacobian matrix.
-We use a last-layer KFAC Laplace variant in our experiments and use the full training set for calculating the GGN instead of a mini-batch based on recent works on the bias in mini-batch estimates (Tatzel et al., 2025).
-Appendix K CIFAR-10 Experiments
+We use a last-layer KFAC Laplace variant in our experiments and use the full training set for calculating the GGN instead of a mini-batch based on recent works on the bias in mini-batch estimates [35].
+Appendix J CIFAR-10 Experiments
 This appendix section repeats the experiments presented in the main paper on the CIFAR-10 dataset.
-For a detailed description of the experimental setup, refer to Appendix H.
-Appendix I describes the used tasks and metrics.
+For a detailed description of the experimental setup, refer to Appendix G.
+Appendix H describes the used tasks and metrics.
 As stated in the main paper, our two research questions are:
-  • What are the effects of changing the learning objective? (Section K.2)
-  • Do we have to sacrifice performance for sample-free predictives? (Section K.1)
-K.1 Quality of Sample-Free Predictives
-Table 2: Comparison of ECE results for different predictives using a fixed Laplace backbone.
-Method Mean Std
+  • What are the effects of changing the learning objective? (Section J.2)
+  • Do we have to sacrifice performance for sample-free predictives? (Section J.1)
+J.1 Quality of Sample-Free Predictives
+Table J.1: Comparison of ECE results for different predictives using a fixed Laplace backbone.
+Method Mean $\pm$ 2 std
 Softmax Laplace
-MC 100 0.0096 0.0013
-MC 1000 0.0102 0.0016
-MC 10 0.0120 0.0013
-Mean Field 0.0121 0.0029
-Laplace Bridge Predictive 0.5933 0.0105
+MC 100 0.0096 $\pm$ 0.0013
+MC 1000 0.0102 $\pm$ 0.0016
+MC 10 0.0120 $\pm$ 0.0013
+Mean Field 0.0121 $\pm$ 0.0029
+Laplace Bridge Predictive 0.5933 $\pm$ 0.0105
 NormCDF Laplace
-Analytic 0.0074 0.0012
-MC 1000 0.0092 0.0022
-MC 100 0.0095 0.0015
-MC 10 0.0100 0.0020
+Closed-form 0.0074 $\pm$ 0.0012
+MC 1000 0.0092 $\pm$ 0.0022
+MC 100 0.0095 $\pm$ 0.0015
+MC 10 $0$$\pm$ 0.0020
 Similarly to the main paper, in this section, we investigate our first research question: whether there is a price to pay for sample-free predictives.
-Table 2 showcases the two best-performing (activation, method) pairs on the ECE metric and the CIFAR-10 dataset: Softmax and NormCDF Laplace.
-Mean Field (MF) is a strong alternative for sample-free predictives, but it has no guarantees and can fall behind MC sampling (see also Figure 1).
-When the constraint is satisfied, the analytic NormCDF predictive is exact.
-Empirically, our analytic predictives always perform on par with MC sampling.
-K.2 Effects of Changing the Learning Objective
-As in the main paper, in this section, we use the best-performing predictive and estimator (see Appendix F) for softmax models and employ our methods with the analytic predictives.
-K.2.1 Calibration and Proper Scoring
-We first evaluate calibration using the log probability scoring rule (Gneiting & Raftery, 2007) and the Expected Calibration Error (ECE) metric (Naeini et al., 2015).
-Figure 5 shows that on CIFAR-10, the score of our analytic predictives with implicit constraints (Sigmoid, NormCDF) are consistently better than the corresponding softmax results for all methods.
-Figure 11: CIFAR-10 ECE results.
-Analytic predictives with implicit ( ) or explicit ( ) constraints outperform Softmax ( ) on HET and SNGP.
+Table J.1 showcases the two best-performing (activation, method) pairs on the ECE metric and the CIFAR-10 dataset: Softmax and NormCDF Laplace.
+Mean Field (MF) is a strong alternative for sample-free predictives, but it has no guarantees and can fall behind MC sampling (see also Fig.
+1).
+Empirically, our closed-form predictives always perform on par with MC sampling.
+J.2 Effects of Changing the Learning Objective
+As in the main paper, in this section, we use the best-performing predictive and estimator (see Appendix E) for softmax models and employ our methods with the closed-form predictives.
+J.2.1 Calibration and Proper Scoring
+We first evaluate calibration using the log probability scoring rule [12] and the Expected Calibration Error (ECE) metric [29].
+Fig.
+3(b) shows that on CIFAR-10, the score of our closed-form predictives (Sigmoid, NormCDF) are consistently better than the corresponding softmax results for all methods.
+Figure J.1: CIFAR-10 ECE results.
+Our closed-form predictives ( ) outperform Softmax ( ) on HET and SNGP.
 Laplace tunes its hyperparameters based on the ECE metric – NormCDF Laplace is the overall best method.
 Note the restricted $y$-limits for readability.
-Figure 11 shows that on CIFAR-10, our analytic predictives have a clear advantage on HET and SNGP.
+Fig.
+J.1 shows that on CIFAR-10, our closed-form predictives have a clear advantage on HET and SNGP.
 Laplace is a post-hoc method that tunes its hyperparameters on the ECE metric, hence its enhanced performance.
 Our NormCDF predictive is on par with Softmax.
-K.2.2 Correctness Prediction
-Figure 12: CIFAR-10 log probability proper scoring results for the binary correctness prediction task.
-Analytic predictives with implicit ( ) constraints consistently outperform Softmax ( ) on all methods.
-Figure 12 shows that on the correctness prediction task, our analytic predictives with implicit constraints outperform all Softmax predictives across all methods, as measured by the log probability proper scoring rule.
-K.2.3 Accuracy
-Figure 13: Analytic predictives do not sacrifice accuracy.
-CIFAR-10 accuracies.
-Analytic predictives with implicit ( ) or explicit ( ) constraints either outperform or are on par with Softmax ( ) across all methods.
-Analytic predictives do not sacrifice accuracy.
-Figure 13 evidences this claim on CIFAR-10: our analytic predictives either outperform or are on par with Softmax predictives.
+J.2.2 Correctness Prediction
+Figure J.2: CIFAR-10 log probability proper scoring results for the binary correctness prediction task. Our closed-form predictives ( ) consistently outperform Softmax ( ) on all methods.
+Fig. J.2 shows that on the correctness prediction task, our closed-form predictives outperform all Softmax predictives across all methods, as measured by the log probability proper scoring rule.
+J.2.3 Accuracy
+Figure J.3: Closed-form predictives do not sacrifice accuracy. CIFAR-10 accuracies. Our closed-form predictives ( ) either outperform or are on par with Softmax ( ) across all methods.
+Closed-form predictives do not sacrifice accuracy.
+Fig.
+J.3 evidences this claim on CIFAR-10: our closed-form predictives either outperform or are on par with Softmax predictives.
 The most accurate method is Sigmoid HET.
 These results support the findings of Wightman et al.
-(2021) that showcase desirable training dynamics of the class-wise cross-entropy loss.
-K.2.4 Out-of-Distribution Detection
-Figure 14: CIFAR-10C OOD detection AUROC results for severity level one. Across all methods, the best-performing predictive is analytic with an implicit constraint ( ).
+[38] that showcase desirable training dynamics of the class-wise cross-entropy loss.
+J.2.4 Out-of-Distribution Detection
+Figure J.4: CIFAR-10C OOD detection AUROC results for severity level one. Across all methods, the best-performing predictive is closed-form ( ).
 Finally, we consider the OOD detection task on a balanced mixture of ID (CIFAR-10) and OOD inputs.
@@
 We use the AUROC metric to evaluate the methods’ performance.
-As shown in Figure 14, the best-performing method is Sigmoid SNGP, an analytic method.
-Generally, Softmax performs on par with our analytic predictives.
+As shown in Fig.
+J.4, the best-performing method is Sigmoid SNGP, a closed-form method.
+Generally, Softmax performs on par with our closed-form predictives.
 Intuitively, separating ID and OOD samples does not require a fine-grained representation of uncertainty, unlike the ECE or proper scoring rules.
-Nevertheless, the analytic predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5).
-Appendix L ImageNet Dirichlet Evaluation
-This appendix section evaluates our moment-matched Dirichlet distributions on the OOD detection task (Appendix I).
-Table 3 shows the per-estimator performances of the three best-performing (method, activation) pairs on ImageNet: Sigmoid Laplace, NormCDF HET, and NormCDF Laplace.
-In all three cases, the best estimator is derived from the moment-matched Dirichlet distributions.
-Table 3: Comparison of OOD AUROC results for different uncertainty estimators on the three best-performing (method, activation) pairs: Sigmoid Laplace, NormCDF HET, and NormCDF Laplace.
-The best estimator is derived from the moment-matched Dirichlet distributions in all three cases.
-Method Mean Std
-Sigmoid Laplace
-Dirichlet Mutual Information 0.6388 0.0005
-Analytic Predictive Entropy 0.6353 0.0003
-MC 1000 Predictive Entropy 0.6323 0.0001
-MC 100 Predictive Entropy 0.6323 0.0001
-MC 1000 Expected Entropy 0.6321 0.0001
-NormCDF HET
-Dirichlet Expected Entropy 0.6353 0.0021
-MC 1000 Expected Entropy 0.6281 0.0020
-MC 100 Expected Entropy 0.6281 0.0020
-MC 10 Expected Entropy 0.6281 0.0020
-Analytic Predictive Entropy 0.6277 0.0021
-NormCDF Laplace
-Dirichlet Expected Entropy 0.6321 0.0009
-MC 100 Predictive Entropy 0.6296 0.0012
-MC 1000 Predictive Entropy 0.6296 0.0012
-Closed-Form Predictive Entropy 0.6296 0.0012
-MC 10 Predictive Entropy 0.6294 0.0012
-Appendix M Further Out-of-Distribution Detection Results
-Figure 15 shows OOD detection results across all ImageNet-C severity levels. Our analytic predictives with implicit constraints consistently outperform Softmax.
-(a) OOD detection AUROC with severity level one.
-(b) OOD detection AUROC with severity level two.
-(c) OOD detection AUROC with severity level three.
-(d) OOD detection AUROC with severity level four.
-(e) OOD detection AUROC with severity level five.
-Figure 15: The OOD detection performance of all methods increases steadily as we increase the severity of the perturbed half of the mixed dataset on the ImageNet validation dataset.
-Our analytic predictives consistently outperform Softmax.
+Nevertheless, the closed-form predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5).
+Appendix K Additional Results
+K.1 Closed-form Softmax Predictives
+One can directly apply Eq.
+10 to a neural network trained with the softmax cross-entropy loss in Eq.
+4 to obtain closed-form predictives.
+However, we empirically found this approach to decrease performance, as the denominator of Eq.
+10 increased to the order of billions during training.
+Considering a multivariate normal random variable $\bm{x}\sim\mathcal{N}(\bm{\mu},\Sigma)$ representing a logit distribution,
+$\mathrm{Var}\left(\sum_{i=1}^{C}e^{x_{i}}\right)\geq\mathrm{Var}(e^{x_{j}})=e^{2\mu_{j}+\Sigma_{jj}}\underbrace{(e^{\Sigma_{jj}}-1)}_{>0},$ (102)
+i.e., when $e^{\mu_{j}}$ is large for some $j\in\{1,\dots,C\}$, the variance of the denominator necessarily explodes, violating our assumption.
+To mitigate this issue, one may optimize the regularised cross-entropy loss
+$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})=\tilde{\mathcal{L}}((x_{n},c_{n})_{n=1}^{N})+\lambda\sum_{n=1}^{N}\left(\sum_{c=1}^{C}\exp(f_{c}(x_{n}))-1\right)^{2}$ (103)
+or the more numerically stable version
+$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})=\tilde{\mathcal{L}}((x_{n},c_{n})_{n=1}^{N})+\lambda\sum_{n=1}^{N}\left(\log\sum_{c=1}^{C}\exp(f_{c}(x_{n}))\right)^{2}$ (104)
+for an appropriately tuned $\lambda$ hyperparameter.
+The latter formulation can be stably trained even at an ImageNet scale.
+However, there are fundamental limitations to the predictive given by Eq.
+10.
+Namely, for isotropic logit-space Gaussian distributions, $\nicefrac{{\sigma^{2}(x)}}{{2}}$ introduces a constant shift in the logits, under which the softmax activation function is invariant.
+Therefore, the predictive approximation collapses into the MAP prediction.
+The SNGP method predicts such isotropic logit-space Gaussians.
+Further, we empirically found that the Laplace method’s predictives for models trained with the regularised cross-entropy loss were also approximately isotropic; thus, this integral approximation yielded diminishing returns.
+For completeness, we share the ImageNet and CIFAR-10 results of $\varphi=\exp$ in Tables K.1 and K.2.
+Table K.1: Performance metrics of $\varphi=\exp$ on the CIFAR-10 validation dataset. We report the mean and two standard deviations.
+Metric Exp SNGP Exp HET Exp Laplace
+NLL $\phantom{-}0.376\pm 0.007$ $\phantom{-}0.353\pm 0.059$ $\phantom{-}0.370\pm 0.024$
+ECE $\phantom{-}0.041\pm 0.006$ $\phantom{-}0.029\pm 0.007$ $\phantom{-}0.045\pm 0.007$
+Log Prob. $-0.255\pm 0.005$ $-0.238\pm 0.038$ $-0.255\pm 0.016$
+Accuracy $\phantom{-}0.884\pm 0.013$ $\phantom{-}0.888\pm 0.023$ $\phantom{-}0.893\pm 0.007$
+AUROC $\phantom{-}0.592\pm 0.006$ $\phantom{-}0.603\pm 0.009$ $\phantom{-}0.596\pm 0.011$
+Table K.2: Performance metrics of $\varphi=\exp$ on the ImageNet validation dataset. We report the mean and two standard deviations.
+Metric Exp SNGP Exp HET Exp Laplace
+NLL $\phantom{-}0.866\pm 0.031$ $\phantom{-}0.929\pm 0.042$ $\phantom{-}0.986\pm 0.149$
+ECE $\phantom{-}0.058\pm 0.004$ $\phantom{-}0.022\pm 0.007$ $\phantom{-}0.058\pm 0.013$
+Log Prob. $-0.380\pm 0.009$ $-0.403\pm 0.011$ $-0.402\pm 0.014$
+Accuracy $\phantom{-}0.784\pm 0.001$ $\phantom{-}0.779\pm 0.001$ $\phantom{-}0.785\pm 0.002$
+AUROC $\phantom{-}0.606\pm 0.001$ $\phantom{-}0.611\pm 0.001$ $\phantom{-}0.615\pm 0.003$
+K.2 Vision Transformer Results on ImageNet
+Table K.3 shows results on ViT Little [11] backbones from the timm [37] library on the ImageNet validation set.
+As in the main paper, the hyperparameters are optimized using a ten-step Bayesian hyperparameter optimization scheme of Weights and Biases.
+Table K.3: NLL and accuracy metrics of ViT Little models on the ImageNet validation set. Error bars represent two standard deviations.
+Method NLL Accuracy (%)
+Laplace
+Softmax $0.81988\pm 0.0020$ $79.598\pm 0.145$
+NormCDF $0.79603\pm 0.0087$ $80.678\pm 0.042$
+Sigmoid $0.79479\pm 0.0002$ $80.104\pm 0.206$
+HET
+Softmax $0.87096\pm 0.0242$ $78.604\pm 0.378$
+NormCDF $0.79106\pm 0.0075$ $80.514\pm 0.425$
+Sigmoid $0.88175\pm 0.0356$ $78.738\pm 0.324$
+GP
+Softmax $0.86724\pm 0.0229$ $78.898\pm 0.187$
+NormCDF $0.76974\pm 0.0088$ $80.694\pm 0.575$
+Sigmoid $0.81178\pm 0.0115$ $79.980\pm 0.412$
+K.3 CIFAR-100 Results
+We repeat the NLL and accuracy evaluation on CIFAR-100 using WideResNet-28-5 models following a ten-step Bayesian hyperparameter sweep on Weights and Biases. Results are shown in Table K.4.
+Table K.4: NLL and accuracy metrics of WideResNet-28-5 models on the CIFAR-100 test set. Error bars represent two standard deviations.
+Method NLL Accuracy (%)
+Laplace
+Softmax $0.92176\pm 0.0128$ $78.225\pm 0.412$
+NormCDF $0.88823\pm 0.0094$ $78.900\pm 0.325$
+Sigmoid $0.94209\pm 0.0136$ $78.050\pm 0.436$
+HET
+Softmax HET $0.98513\pm 0.0254$ $77.050\pm 0.228$
+NormCDF HET $0.94643\pm 0.0022$ $78.117\pm 0.286$
+Sigmoid HET $0.95251\pm 0.0142$ $77.583\pm 0.175$
+SNGP
+Softmax SNGP $0.97156\pm 0.0348$ $77.750\pm 0.462$
+NormCDF SNGP $0.90011\pm 0.0106$ $79.010\pm 0.298$
+Sigmoid SNGP $0.99996\pm 0.0265$ $78.350\pm 0.215$
+K.4 Alignment with the True Predictives
+Table K.5 shows the divergence of predictives using various approximation techniques from the true predictive estimated using $10,000$ Monte Carlo samples.
+Our closed-form predictives, NormCDF and Sigmoid, are favorable to both the mean field and Laplace bridge approximations.
+Table K.5: Kullback-Leibler (KL) divergence to the true predictive distributions (estimated using $10,000$ Monte Carlo samples) for different approximation methods on the ImageNet validation set using ViT Little backbones.
+The mean KL divergence is calculated over the validation set.
+Error bars represent two standard deviations over independently trained models using different seeds.
+Method KL Divergence to True Predictive
+NormCDF $0.0057\pm 0.0008$
+Sigmoid $0.0064\pm 0.0009$
+Softmax mean field $0.0330\pm 0.0042$
+Softmax Laplace bridge $1.7900\pm 0.1254$
+K.5 BCE Loss Performance Gains
+Table K.6 shows that the Softmax and Sigmoid activation functions (and corresponding losses) already show improved performance on the vanilla models compared to softmax, highlighting that the performance gains we observe cannot only be attributed to our closed-form predictive approximations but also the favorable training dynamics of the class-wise BCE loss.
+Table K.6: NLL results on ImageNet using ResNet-50 backbones and different activation functions. Error bars represent two standard deviations.
+Activation Function NLL
+NormCDF $0.9345\pm 0.0072$
+Softmax $0.9369\pm 0.0025$
+Sigmoid $0.9146\pm 0.0046$
+K.6 Further Out-of-Distribution Detection Results
+Fig. K.1 shows OOD detection results across all ImageNet-C severity levels. Our closed-form predictives consistently outperform Softmax.
+Figure K.1: The OOD detection performance of all methods increases steadily as we increase the severity of the perturbed half of the mixed dataset on the ImageNet validation dataset.
+Our closed-form predictives consistently outperform Softmax.

### v3

@@
 We evaluate it combined with several approximate Gaussian inference methods (Laplace, HET, SNGP) on large- and small-scale datasets (ImageNet, CIFAR-100, CIFAR-10), demonstrating improved uncertainty quantification capabilities compared to softmax MC sampling.
+Our code is available at github.com/bmucsanyi/probit.
 1 Introduction
@@
 $\Delta^{C-1}=\{(p_{1},\dots,p_{C})\in[0,1]^{C}:p_{1}+\dots+p_{C}=1\}$ (1)
-is the $(C-1)$-dimensional probability simplex.
+is the (C1)(C-1)-dimensional probability simplex.
 The model $\bm{h}$ outputs the probability of inputs $x\in\mathcal{X}$ belonging to each class, $h_{c}(x)=p(c\mid x)$.
@@
 Notably, the output probability does not take into account the uncertainty of the model due to the finite nature of the data.
-In the uncertainty quantification formalism, such models can only estimate aleatoric uncertainty and disregard epistemic uncertainty [15].
+In the uncertainty quantification formalism, such models can only estimate aleatoric uncertainty and disregard epistemic uncertainty [hullermeier2021aleatoric].
 Figure 1: Mean KL divergence between the ‘true’ predictive (approximated with a $10000$ sample MC approximation) and different predictive approximations on a synthetic data set.
@@
 The MC approximations are capped at a fixed computational budget of $10000$ class samples, i.e.
-$\lceil 10000/C\rceil$ samples (see also Section C.1).
+$\lceil 10000/C\rceil$ samples (see also Section˜C.1).
 The probabilistic way to capture epistemic uncertainty is to require the model to output a second-order distribution (that is ‘a distribution over probability distributions’).
-Namely, for each input $x\in\mathcal{X}$, the model should output a probability measure over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$, as opposed to a point estimate in the simplex $\bm{h}(x)\in\Delta^{C-1}$ [32, 4].
+Namely, for each input $x\in\mathcal{X}$, the model should output a probability measure over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$, as opposed to a point estimate in the simplex $\bm{h}(x)\in\Delta^{C-1}$ [sensoy2018evidential, charpentier2020posterior].
 A number of methods for such distributional uncertainty quantification in classification rely on approximate Gaussian inference in logit space. That is, $\mathrm{h}$ is written as a composition
@@
 where $\mathrm{f}$ is learned from the training data, $\mathcal{G}(\mathbb{R}^{C})$ is the set of Gaussian measures on the logit space $\mathbb{R}^{C}$, and $\operatorname{\mathbf{softmax}}_{*}$ pushes forward the Gaussian probability measures through the softmax.
-For example, Heteroscedastic Classifiers (HET) [5] learn the output means and covariances of $\mathrm{f}$ explicitly with a neural network.
-In linearised Laplace approximations [8], only the means are learned explicitly with a neural network, while the covariances are obtained post-hoc from that neural network by approximating its parameter posterior distribution with a Gaussian, and then pushing it forward to logit space by linearising the network in the parameters.
-Last-layer Laplace approximations [16] are a variant of linearised Laplace approximations where only the posterior over the last-layer parameters is Gaussian-approximated, requiring no linearisation of the network.
-Spectral-Normalized Gaussian processes (SNGP) [20] are neural networks with last layers that are approximately Gaussian processes, through random feature expansions and Laplace approximations.
-To obtain predictive probabilities from a distributional classifier as in Eq.
-5, we marginalise it out by the measure-theoretic change of variables, using $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$:
+For example, Heteroscedastic Classifiers (HET) [collier2021correlated] learn the output means and covariances of $\mathrm{f}$ explicitly with a neural network.
+In linearised Laplace approximations [daxberger2021laplace], only the means are learned explicitly with a neural network, while the covariances are obtained post-hoc from that neural network by approximating its parameter posterior distribution with a Gaussian, and then pushing it forward to logit space by linearising the network in the parameters.
+Last-layer Laplace approximations [kristiadi_being_2020] are a variant of linearised Laplace approximations where only the posterior over the last-layer parameters is Gaussian-approximated, requiring no linearisation of the network.
+Spectral-Normalised Gaussian processes (SNGP) [liu2020simple] are neural networks with last layers that are approximately Gaussian processes, through random feature expansions and Laplace approximations.
+To obtain predictive probabilities from a distributional classifier as in Eq.˜5, we marginalise it out by the measure-theoretic change of variables, using $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$:
 $\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]$ $=\int_{\Delta^{C-1}}p\;\mathrm{h}(x)(dp)=\int_{\mathbb{R}^{C}}\operatorname{\mathbf{softmax}}(y)\;\mathrm{f}(x)(dy).$ (6)
-In addition to predictive probabilities, the distributional framework allows for the acquisition of other quantities of interest, such as variances, entropy, and information-theoretic decompositions of the predictive uncertainty into aleatoric and epistemic parts [10, 28, 39].
+In addition to predictive probabilities, the distributional framework allows for the acquisition of other quantities of interest, such as variances, entropy, and information-theoretic decompositions of the predictive uncertainty into aleatoric and epistemic parts [depeweg2018decomposition, mukhoti2023deep, wimmer2023quantifying].
 These are usually obtained by Monte Carlo (MC) sampling from the Gaussian measure $\mathrm{f}(x)$.
-Indeed, while the probability density of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ can be obtained in closed form using a change of variables formula [1], this does not seem of practical use—even predictives cannot be computed analytically.
+Indeed, while the probability density of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ can be obtained in closed form using a change of variables formula [atchison_logistic-normal_1980], this does not seem of practical use—even predictives cannot be computed analytically.
 On the other hand, given a variance, the runtime and memory cost required for an MC approximation to achieve an estimator of said variance grows linearly with the number of classes.
 Crucially, this computational cost is not limited to training time, but also to inference time, when the computational budget may be much more limited.
-This limits the application of such Gaussian inference methods in classification with a large number of classes, such as those common in computer vision and language modeling.
-1.1 Proposed Recipe
-We summarise our framework as a 3-step recipe:
+This limits the application of such Gaussian inference methods in classification with a large number of classes, such as those common in computer vision and language modelling.
+Summary of Contributions: Our work extends the classical probit approximation for the Gaussian sigmoid integral in binary classification [spiegelhalter_sequential_1990, mackay_evidence_1992] in the following manner:
   1. (i)
-    Choose an approximate Gaussian inference method that outputs logit means and covariances.
+    Generalising it to the multi-class setting (Sections˜3.1 and 3.2).
   2. (ii)
-    Choose an output activation that yields a closed-form Gaussian integral (Section 2).
+    Analysing it theoretically (Section˜3.3).
   3. (iii)
-    At inference time, use our closed-form formulas (see Section 2 for the predictives, Section 3 for other quantities of interest).
+    Analysing it empirically, and how it outperforms previous approximations (Figs.˜1 and 7).
+  4. (iv)
+    Generalising it to second-order distributional Dirichlet approximations (Section˜4).
+  5. (v)
+    Combining it with advantageous training (Sections˜5 and 7).
+Our proposed framework can be summarised as a 3-step recipe:
+(i) Choose an approximate Gaussian inference method that outputs logit means and covariances.
+(ii) Choose an output activation that yields a closed-form Gaussian integral (Section˜3).
+(iii) At inference time, use our closed-form formulas (see Section˜3 for the predictives, Section˜4 for other quantities of interest).
+2 Related Work
+In uncertainty quantification, methods that output logit Gaussians, which we refer to as approximate Gaussian inference methods, have become one of the dominant paradigms in uncertainty quantification [collier2021correlated, daxberger2021laplace, kristiadi_being_2020, liu2020simple].
+In classification tasks, such methods have the downside of requiring sampling at inference to transform Gaussian distributions into predictive probability vectors.
+In parallel, prior work has developed ways to propagate probabilistic uncertainty through neural networks [gast_lightweight_2018].
+Several works develop analytic propagation of covariances up to the logits [wright_analytic_2024, li_streamlining_2024]; our method complements these by addressing the propagation of the logit-space Gaussian distributions onto the probability simplex.
+Few works address this last step.
+The Laplace bridge [pmlr-v180-hobbhahn22a] proposes a map from Gaussians to Dirichlets on the probability simplex.
+However, this approach has no guarantees on the quality of its predictives, and it severely underperforms Monte Carlo sampling.
+Other approaches include the mean-field approximation to the Gaussian softmax integral [lu_mean-field_2021] and the approximations of shekhovtsov_feed-forward_2018.
+In this work, we propose approximate predictives that outperform existing closed-form predictive techniques and provide second-order Dirichlet distributions for other closed-form quantities of interest.
+Furthermore, this is the first work to provide a theoretical analysis of such an approximation.
+Other works have also proposed various activations for classification to replace the softmax.
+Particularly relevant is [galy-fajou_multi-class_2020], which composes element-wise sigmoids with normalisation and utilises auxiliary-variable augmentation to obtain closed-form ELBO terms during training; however, its predictives require Monte Carlo estimation.
+Related conjugate/augmentation-based objectives exist for stick-breaking and tree-structured multi-class GPs [linderman_dependent_2015, achituve_gp-tree_2021], and for BNNs, a tempered-likelihood/Dirichlet perspective with a factorised-Gaussian surrogate has been explored [kapoor_uncertainty_2022].
+Yet, all these approaches still require sampling or numerical quadrature at inference time.
+[spiegelhalter_sequential_1990, mackay_evidence_1992, kristiadi_being_2020] use the probit approximation in the binary case, which we generalise to the multi-class setting.
+[titsias_rc_aueb_one-vs-each_2016] approximates the softmax by providing lower bounds for point estimate training, but replacing the softmax with these bounds during training does not yield a normalised multi-class likelihood.
+The binary cross-entropy loss with the element-wise sigmoid activation has been reported to outperform cross-entropy in large-scale settings [wightman_resnet_2021, li_bce_2025].
+Our work is the first to use this advantageous loss while normalising the output to obtain proper multi-class probabilistic models with closed-form predictives from logit Gaussians.
+Evidential deep learning (EDL) approaches [sensoy2018evidential] offer sample-free outputs through Dirichlet distributions but have been shown to be unreliable for uncertainty estimation and disentanglement [NEURIPS2022_bc1d640f, jürgens2024epistemicuncertaintyfaithfullyrepresented, pandey2023learn, mucsányi2024benchmarking].
+We focus on making approximate Gaussian inference methods sample-free in a principled and performant manner; however, for completeness, we compare our approach with EDL methods in Section˜K.8.
+Finally, multiple authors have explored continuous approximations to categorical distributions, which enable gradient-based optimization for discrete stochastic models through sampling, such as the Gumbel-max trick [maddison_ast_2014, jang_categorical_2017, maddison_concrete_2017, huijben_review_2023].
+These methods principally optimise a single categorical distribution, while the present work is concerned with constructing probability densities over the (continuous) simplex of such categorical distributions.
 3 Gaussian Inference with Closed-Form Predictives
@@
 The softmax function is the composition $\bm{\operatorname{\mathbf{softmax}}}=\bm{n}\circ\bm{\exp}$, where $\bm{n}$ is the normalisation function given by $\bm{n}(\bm{q}):=\bm{q}/(q_{1}+\dots+q_{C})$ and $\bm{\exp}$ is applied element-wise.
-Thus, $\mathrm{h}$ can be written as the composition (see Eq.
-5)
+Thus, $\mathrm{h}$ can be written as the composition (see Eq.˜5)
 $\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\bm{\exp}_{*}}\mathcal{P}((0,\infty)^{C})\xrightarrow{\bm{n}_{*}}\mathcal{P}(\Delta^{C-1}).$ (7)
-Now it is straightforward to see that (Section A.1)
+Now it is straightforward to see that (Section˜A.1)
 $\int_{\mathbb{R}}\exp(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)=\exp\left(\mu+\frac{\sigma^{2}}{2}\right)$ (8)
@@
 The previous argumentation can be applied to a general family of output activation functions: suppose the output activation is of the form $\bm{n}\circ\bm{\varphi}$, where $\bm{\varphi}$ is the element-wise application of some activation function $\varphi\colon\mathbb{R}\to(0,\infty)$.
-Then we have (see Eq.
-7)
+Then we have (see Eq.˜7)
 $\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\bm{\varphi}_{*}}\mathcal{P}((0,\infty)^{C})\xrightarrow{\bm{n}_{*}}\mathcal{P}(\Delta^{C-1}).$ (11)
-Solving the one-dimensional integrals $\int_{\mathbb{R}}\varphi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)$ (see Eq. 8) allows us to obtain the approximate predictive (see Eq. 10)
+Solving the one-dimensional integrals $\int_{\mathbb{R}}\varphi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)$ (see Eq.˜8) allows us to obtain the approximate predictive (see Eq.˜10)
 $\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{Q}]}{\sum_{c=1}^{C}\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}\left[Q_{c}\right]}.$ (12)
-For instance, taking $\varphi=\Phi$, the standard Normal cumulative distribution function (normCDF) [^1] The resulting model is distinct from a multinomial probit model, see Appendix B., we have by a classical result (Section A.2)
+For instance, taking $\varphi=\Phi$, the standard Normal cumulative distribution function (normCDF) [^1] The resulting model is distinct from a multinomial probit model, see Appendix B., we have by a classical result (Section˜A.2)
 $\int_{\mathbb{R}}\Phi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{1+\sigma^{2}}}\right).$ (13)
@@
 $\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]\approx\frac{\bm{\Phi}\left(\frac{\bm{\mu}(x)}{\sqrt{1+\bm{\sigma}^{2}(x)}}\right)}{\sum_{c=1}^{C}\Phi\left(\frac{\mu_{c}(x)}{\sqrt{1+\sigma^{2}_{c}(x)}}\right)}.$ (14)
-Taking instead $\varphi=\rho$, the logistic sigmoid, we approximate Eq. 13 using the probit approximation (Section A.3),
+Taking instead $\varphi=\rho$, the logistic sigmoid, we approximate Eq.˜13 using the probit approximation (Section˜A.3),
 $\int_{\mathbb{R}}\rho(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)$ (15)
@@
 3.3 Theoretical Analysis of the Approximation’s Quality
-In Fig.
-1, we see empirically that the quality of our predictive approximations outperforms existing predictive approximations on synthetic data.
+In Fig.˜1, we see empirically that the quality of our predictive approximations outperforms existing predictive approximations on synthetic data.
 We now explore the quality of our predictive approximations theoretically.
@@
 $q=q(\mu,\sigma^{2}):=\mathbb{E}_{Q\sim\rho_{*}\mathcal{N}(\mu,\sigma^{2})}[Q]\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)=:\hat{q}(\mu,\sigma^{2})=\hat{q}$, (17)
-where $\hat{q}$ is the probit approximation (Section A.3).
+where $\hat{q}$ is the probit approximation (Section˜A.3).
 Moreover we write $q_{c}:=q(\mu_{c},\sigma_{c}^{2})$, $\hat{q}_{c}:=\hat{q}(\mu_{c},\sigma_{c}^{2})$ where $(\mu_{c},\sigma_{c}^{2})$ is the logit mean-variance pair for class $c$, $\bm{p}$ for the true predictive and $\bm{\hat{p}}$ for our approximate predictive, i.e.
-$\bm{p}:=\mathbb{E}_{\bm{P}\sim(\bm{n}\circ\bm{\rho})_{*}\mathcal{N}(\bm{\mu},\bm{\operatorname{\mathbf{diag}}}(\bm{\sigma}^{2}))}[\bm{P}],\quad\bm{\hat{p}}:=\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}}.$ (18)
-To analyse the quality of the approximation, we thus want to bound $\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$.
-Clearly, when all the $Q_{c}$ are deterministic (i.e.
-constant random variables), where $\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)$, we have $\bm{p}=\bm{\hat{p}}$.
-Alternatively, from the derivation Eq.
-10, we see that in the limiting case Eq.
-12 is exact when $\sum_{c=1}^{C}Q_{c}$ is deterministic.
-This is a weaker assumption than requiring all $Q_{c}$ to be deterministic individually.
-In fact, we can derive a rate of decay of $\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$ as a function of $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)$:
+$\bm{p}:=\mathbb{E}_{\bm{P}\sim(\bm{n}\circ\bm{\rho})_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}],\quad\bm{\hat{p}}:=\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}},$ (18)
+[…]
 Theorem 3.1.
@@
 as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, for some constant $M(\mathcal{K})\in[0,\infty]$ which only depends on $\mathcal{K}$, not on $C$, such that $M(\mathcal{K})<\infty$ when $\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$, and $M(\mathcal{K})=0$ when $q=\hat{q}$ for all $(\mu,\sigma^{2})\in\mathcal{K}$.
-Section C.2 provides a formula for $M(\mathcal{K})$ and a proof.
+Section˜C.2 provides a formula for $M(\mathcal{K})$ and a proof.
 We also conduct an empirical investigation of whether the assumptions made in the theorem hold in practice, namely whether the logit means and variances lie in a compact set, and whether $M(\mathcal{K})<\infty$.
-The $O$ term in Eq.
-19 is due to the splitting of the expectation into the numerator and denominator in the derivation Eq.
-10, while $M(\mathcal{K})$ results from the error due to the probit approximation.
-Importantly, Theorem 2.1 states that the latter does not increase with the number of classes.
+The $O$ term in Eq.˜19 is due to the splitting of the expectation into the numerator and denominator in the derivation Eq.˜10, while $M(\mathcal{K})$ results from the error due to the probit approximation.
+Importantly, Theorem˜3.1 states that the latter does not increase with the number of classes.
 4 Distributional Approximations of the Gaussian Pushforwards
@@
 As previously discussed, such a distribution allows, for instance, closed-form decompositions of aleatoric and epistemic uncertainties.
-The formulation of Eq. 11 does not yet allow this. However, we will now see that one can obtain good Dirichlet approximations by constructing a tractable approximation of Eq. 11 of the form
+The formulation of Eq.˜11 does not yet allow this. However, we will now see that one can obtain good Dirichlet approximations by constructing a tractable approximation of Eq.˜11 of the form
 $\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\quad\mathrm{a}\quad}\mathcal{D}(\Delta^{C-1}),$ (20)
 where $\mathcal{D}(\Delta^{C-1})$ is the space of Dirichlet distributions on $\Delta^{C-1}$.
-The family of Dirichlets presents an ideal choice of distributions over the simplex as, for such distributions, all quantities of interest are closed-form (see Appendix F).
+The family of Dirichlets presents an ideal choice of distributions over the simplex as, for such distributions, all quantities of interest are closed-form (see Appendix˜F).
 The map $\mathrm{a}$ is constructed by moment matching: we match the first two moments of the Dirichlet distribution to the moments of the Gaussian pushforward $\bm{\varphi}_{*}\mathrm{f}(x)$.
-Conveniently, like the first moments described in Section 2, we can obtain second moments for $\exp$
+Conveniently, like the first moments described in Section˜3, we can obtain second moments for $\exp$
 $\int_{\mathbb{R}}\exp(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)=\exp(2\mu+2\sigma^{2})$ (21)
-and for $\Phi$ (Owen [30, Eq. 20,010.4])
+and for $\Phi$ (owen_table_1980)
 $\int_{\mathbb{R}}\Phi(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{1+\sigma^{2}}}\right)-2T\left(\frac{\mu}{\sqrt{1+\sigma^{2}}},\frac{1}{\sqrt{1+2\sigma^{2}}}\right),$ (22)
-where $T$ is Owen’s T function. For $\rho$, we have the following approximation (Daunizeau [7, Equation 23])
+where $T$ is Owen’s T function. For $\rho$, we have the following approximation (daunizeau_semi-analytical_2017)
 $\int_{\mathbb{R}}\rho(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)-\frac{\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\left(1-\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\right)}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}.$ (23)
@@
 $\mathbb{E}_{\bm{P}\sim\bm{n}_{*}\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{P}^{2}]$ $=\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}\left[\frac{\bm{Q}^{2}}{\left(\sum_{c=1}^{C}Q_{c}\right)^{2}}\right]\approx\frac{\mathbb{E}_{\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)}[\bm{Q}^{2}]}{\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}\left[\sum_{c=1}^{C}Q_{c}\right]^{2}}.$ (24)
-As in Eq. 12, the approximation in Eq. 24 is exact when $\sum_{c=1}^{C}Q_{c}$ is a constant random variable.
-If we were to directly attempt to infer the parameters of the Dirichlet $\mathrm{h}(x)$ from Eqs.
-12 and 24, we would obtain an overparametrised system of equations with $2C$ equations and $C$ unknowns.
-Thus, we instead use a classical Dirichlet method of moments [25, Eqs.
+As in Eq.˜12, the approximation in Eq.˜24 is exact when $\sum_{c=1}^{C}Q_{c}$ is a constant random variable.
+If we were to directly attempt to infer the parameters of the Dirichlet $\mathrm{h}(x)$ from Eqs.˜12 and 24, we would obtain an overparametrised system of equations with $2C$ equations and $C$ unknowns.
+Thus, we instead use a classical Dirichlet method of moments [minka_2000, Eqs.
 19 & 23], which reduces these equations, as we will now describe.
@@
 for any $1\leq c\leq C$.
-So, a natural estimate of $\sum_{c=1}^{C}\gamma_{c}$ that uses all $P_{c}$ is the geometric mean of Eq.
-25.
-Since the mean of the Dirichlet is $\bm{\gamma}/\sum_{c=1}^{C}\gamma_{c}$, we obtain using Eq.
-25 an expression for the Dirichlet parameters
+So, a natural estimate of $\sum_{c=1}^{C}\gamma_{c}$ that uses all $P_{c}$ is the geometric mean of Eq.˜25.
+Since the mean of the Dirichlet is $\bm{\gamma}/\sum_{c=1}^{C}\gamma_{c}$, we obtain using Eq.˜25 an expression for the Dirichlet parameters
 $\left(\prod_{c=1}^{C}\frac{\mathbb{E}[P_{c}]-\mathbb{E}[P^{2}_{c}]}{\mathbb{E}[P^{2}_{c}]-\mathbb{E}[P_{c}]^{2}}\right)^{1/C}\mathbb{E}[\bm{P}].$ (26)
-Hence, using Eqs. 12 and 24, we obtain the computationally feasible Dirichlet parameters
+Hence, using Eqs.˜12 and 24, we obtain the computationally feasible Dirichlet parameters
 $\bm{\gamma}:=\left(\prod_{c=1}^{C}\frac{\mathbb{E}[Q_{c}]\cdot S-\mathbb{E}[Q^{2}_{c}]}{\mathbb{E}[Q^{2}_{c}]-\mathbb{E}[Q_{c}]^{2}}\right)^{1/C}\left(\frac{\mathbb{E}[\bm{Q}]}{\sum_{c=1}^{C}\mathbb{E}[Q_{c}]}\right)$ (27)
@@
 Taking the maximum with $1$ in the expression for $S$ is an additional approximation to ensure $\bm{\gamma}>\bm{0}$.
-Note that by performing such moment matching to obtain the Dirichlet parameters in Eq.
-27, the mean of this Dirichlet matches precisely the approximate predictive Eq.
-12.
-That is, the approximate predictive of the exact distributional model (Eq.
-11) matches the exact predictive of the approximate distributional model (Eq.
-20).
+Note that by performing such moment matching to obtain the Dirichlet parameters in Eq.˜27, the mean of this Dirichlet matches precisely the approximate predictive Eq.˜12.
+That is, the approximate predictive of the exact distributional model (Eq.˜11) matches the exact predictive of the approximate distributional model (Eq.˜20).
 4.2 Intermediate Beta Matching
@@
 With normCDF, we can match exact moments, whereas with sigmoid, we match approximate moments.
-Moreover, when $\mathrm{f}(x)$ is the standard Normal $\mathcal{N}(0,1)$ (on the left in blue), the approximation for the pushforward through normCDF is exact (Section D.1).
-When $\varphi=\Phi\text{ or }\rho$, noting that their range is $(0,1)$, we can decompose the approximate distributional model (Eq. 20) further:
+Moreover, when $\mathrm{f}(x)$ is the standard Normal $\mathcal{N}(0,1)$ (on the left in blue), the approximation for the pushforward through normCDF is exact (Section˜D.1).
+When $\varphi=\Phi\text{ or }\rho$, noting that their range is $(0,1)$, we can decompose the approximate distributional model (Eq.˜20) further:
 $\mathrm{h}\colon\mathcal{X}\xrightarrow{\quad\mathrm{f}\quad}\mathcal{G}(\mathbb{R}^{C})\xrightarrow{\mathrm{p}}\mathcal{B}((0,1))^{C}\xrightarrow{\mathrm{n}}\mathcal{D}(\Delta^{C-1}),$ (28)
 where $\mathcal{B}((0,1))$ is the space of Beta distributions on $(0,1)$.
-Like the output Dirichlets, we obtain the intermediate Beta distributions by moment matching (Appendix D).
-Figure 2 indicates that such Beta approximations tend to be of high quality.
-In Section D.1, we give an information-geometric motivation for this, which shows that the choice of activation $\Phi$ is ideal for approximating the Gaussian pushforwards with Beta distributions.
+Like the output Dirichlets, we obtain the intermediate Beta distributions by moment matching (Appendix˜D).
+Figure˜2 indicates that such Beta approximations tend to be of high quality.
+In Section˜D.1, we give an information-geometric motivation for this, which shows that the choice of activation $\Phi$ is ideal for approximating the Gaussian pushforwards with Beta distributions.
 The quality of such Beta approximations is encouraging for that of the output Dirichlet approximations.
@@
 However, we can instead leverage the fact that $\Phi$ and $\rho$ have values in $(0,1)$ to train using the binary cross-entropy (BCE) loss.
-The BCE loss is a valid choice of loss as it is a (negative) strictly proper scoring rule (see Appendix H).
-Moreover, it has been observed to provide advantageous training dynamics compared to the CE loss [37, 19].
-The choice of $\varphi=\Phi$ or $\rho$ not only fits our framework for closed-form predictives, but allows one to leverage such advantageous training dynamics (see experimental results in Section 6).
+The BCE loss is a valid choice of loss as it is a (negative) strictly proper scoring rule (see Appendix˜I).
+Moreover, it has been observed to provide advantageous training dynamics compared to the CE loss [rw2019timm, li_bce_2025].
+The choice of $\varphi=\Phi$ or $\rho$ not only fits our framework for closed-form predictives, but allows one to leverage such advantageous training dynamics (see experimental results in Section˜7).
 For example, in linearised Laplace and SNGP, a neural network $\bm{f}\colon\mathcal{X}\to\mathbb{R}^{C}$ is trained for the mean of $\mathrm{f}$, while the covariance is obtained post-hoc.
@@
 $\mathcal{L}((x_{n},c_{n})_{n=1}^{N})=-\sum_{n=1}^{N}\sum_{c=1}^{C}\big(\delta_{c_{n},c}\log(\varphi(f_{c}(x)))+(1-\delta_{c_{n},c})\log(1-\varphi(f_{c}(x))\big).$ (29)
+In contrast to the previously mentioned methods, HET directly trains not only for the mean $\bm{f}$ but also for the covariance matrix $\bm{\Sigma}$ of $\mathrm{f}$ through the predictives.
+Thus, in this case, we use the BCE with the pushforward means $\bar{Q}_{c}:=\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}[Q_{c}]$:
+$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})=-\sum_{n=1}^{N}\sum_{c=1}^{C}\big(\delta_{c_{n},c}\log(\bar{Q}_{c})+(1-\delta_{c_{n},c})\log(1-\bar{Q}_{c})\big).$ (30)
 6 Computational Gains
@@
 The $C^{2}$ term in memory comes from the logit covariance matrix; the $C^{3}$ term in runtime comes from the need to Cholesky-decompose such a covariance matrix before sampling.
-For example, on a ResNet-50, sampling one thousand logit vectors from the logit-space Gaussian means an approximately 7% overhead on the forward pass.
+For example, on a ResNet-50, sampling one thousand logit vectors from the logit-space Gaussian incurs approximately 7% overhead on the forward pass: see Appendix˜H for MC timing experiments.
 In embedded systems with much smaller models, this overhead is even larger.
 In contrast, the runtime and memory cost of calculating our predictives is $\mathcal{O}(C)$, which is negligible in comparison.
-Indeed, the approximation Eq.
-12 uses only the mean and the diagonal of the logit covariance to calculate the predictive.
+Indeed, the approximation Eq.˜12 uses only the mean and the diagonal of the logit covariance to calculate the predictive.
 That is, the logit quantities required to compute a predictive scale as $\mathcal{O}(C)$ instead of $\mathcal{O}(C^{2})$.
-This can provide further opportunities for computational savings in the methods used to obtain the Gaussian distributions.
-In HET, for example, the logit covariance is trained with both a diagonal term and a low-rank term. Our framework thus allows one to drop the low-rank term (Section I.2).
+This provides further opportunities for computational savings in the methods used to obtain the Gaussian distributions.
+E.g., in HET, the logit covariance is trained with both a diagonal term and a low-rank term.
+Our framework thus allows dropping the low-rank term (Section˜J.2).
 7 Experiments
@@
   1. (i)
-    Do we have to sacrifice performance for sample-free predictives? (Section 6.1)
+    Do we have to sacrifice performance for sample-free predictives? (Section˜7.1)
   2. (ii)
-    What are the effects of changing the learning objective? (Section 6.2)
+    What are the effects of changing the learning objective? (Section˜7.2)
 To answer the first question, we consider fixed (method, activation) pairs and verify that our methods perform on par with the Monte Carlo sampled predictives.
 In the main paper, we report results on Softmax ($\varphi=\exp$, trained with vanilla CE), NormCDF ($\varphi=\Phi$) and Sigmoid ($\varphi=\rho$) models.
-In Section K.1, we investigate the performance of Softmax models with closed-form predictives.
+In Section˜K.1, we investigate the performance of Softmax models with closed-form predictives.
 Table 1: Comparison of ECE results for different predictives using a fixed Laplace backbone.
@@
 MC 1000 0.0118 $\pm$ 0.0007
-We consider Heteroscedastic Classifiers (HET) [5], Spectral-Normalized Gaussian Processes (SNGP) [20], and last-layer Laplace-Approximated Networks [8] as backbones supplying logit-space Gaussian distributions.
-The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [9], CIFAR-10, and CIFAR-100 [17] on five metrics aligning with practical needs from uncertainty estimates [26].
+We consider Heteroscedastic Classifiers (HET) [collier2021correlated], Spectral-Normalised Gaussian Processes (SNGP) [liu2020simple], and last-layer Laplace-Approximated Networks [daxberger2021laplace] as backbones supplying logit-space Gaussian distributions.
+The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [deng2009imagenet], CIFAR-10, and CIFAR-100 [krizhevsky2009learning] on five metrics aligning with practical needs from uncertainty estimates [mucsanyi2023trustworthy].
 Further, we test the Out-of-Distribution (OOD) detection capabilities of the models on balanced mixtures of In-Distribution (ID) and OOD samples.
-For ImageNet, we treat ImageNet-C [13] samples with 15 corruption types and 5 severity levels as OOD samples.
+For ImageNet, we treat ImageNet-C [hendrycks2019benchmarking] samples with 15 corruption types and 5 severity levels as OOD samples.
 For CIFAR-10 and CIFAR-100, we use the CIFAR-10C corruptions.
-For the second question, we evaluate our closed-form predictives and moment-matched Dirichlets against softmax models with the Laplace Bridge [14] and Mean Field [21] approximations, as well as MC sampling, in Section 6.2.
+For the second question, we evaluate our closed-form predictives and moment-matched Dirichlets against softmax models with the Laplace Bridge [pmlr-v180-hobbhahn22a] and Mean Field [lu_mean-field_2021] approximations, the linear-time approximation by shekhovtsov_feed-forward_2018, and MC sampling; see Section˜7.2.
 To provide a fair comparison, we reimplement each method as simple-to-use wrappers around deterministic backbones.
-For ImageNet evaluation, we use a ResNet-50 backbone from the timm library [37] pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following Mucsányi et al.
-[27].
-For Vision Transformer [11] experiments, refer to Section K.2.
+For ImageNet evaluation, we use a ResNet-50 backbone from the timm library [rw2019timm] pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following mucsányi2024benchmarking.
+For Vision Transformer [dosovitskiy2020image] experiments, refer to Section˜K.2.
 On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs.
 The CIFAR-100 experiments train WideResNet-28-5 models for 200 epochs.
-We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [33] in Weights & Biases [3].
+We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [shahriari2015taking] in Weights & Biases [wandb].
 The main paper focuses on ImageNet results to highlight the scalability of our framework and only shows proper scoring results for CIFAR-10.
-For a complete overview of CIFAR-10 and CIFAR-100 results, refer to Appendix J and Section K.3, respectively.
+For a complete overview of CIFAR-10 and CIFAR-100 results, refer to Section˜K.3 and Section˜K.4, respectively.
 In all plots of the main paper and the Appendix, the error bars are calculated over five independent runs with different seeds and show the minimum and maximum metric value.
@@
 In this section, we investigate whether there is a price to pay for sample-free predictives and second-order distributions.
-To this end, we take the two best-performing methods on the Expected Calibration Error (ECE) metric [29]—NormCDF and Sigmoid Laplace (see Fig.
-4(a))—and evaluate their per-predictive performance.
-As Table 1 shows, the closed-form predictive is always on par with the (unbiased) MC estimates on ImageNet while being cheaper.
-Section K.4 reports the KL divergence of different approximations to the ‘true’ predictive (estimated using 10,000 MC samples) on ImageNet.
-Table 2 shows that the moment-matched second-order Dirichlet distributions consistently outperform both MC and sample-free predictives on the OOD detection task, showcasing the practical utility of second-order distributional knowledge.
+To this end, we take the two best-performing methods on the Expected Calibration Error (ECE) metric [naeini2015obtaining]—NormCDF and Sigmoid Laplace (see Fig.˜4(a))—and evaluate their per-predictive performance.
+As Table˜1 shows, the closed-form predictive is always on par with the (unbiased) MC estimates on ImageNet while being cheaper.
+Section˜K.5 reports the KL divergence of different approximations to the ‘true’ predictive (estimated using 10,000 MC samples) on ImageNet.
+Table˜2 shows that the moment-matched second-order Dirichlet distributions consistently outperform both MC and sample-free predictives on the OOD detection task, showcasing the practical utility of second-order distributional knowledge.
 (a) Closed-form predictives ( ) are either on par with or outperform Softmax ( ) across all methods.
@@
 The previous section shows that for models trained with Sigmoid or NormCDF, our closed-form predictives always perform on par with MC sampling while being more efficient.
-However, our proposed objective (see Eq.
-29) changes the training dynamics and, subsequently, the performance of models.
+However, our proposed objective (see Eq.˜29) changes the training dynamics and, subsequently, the performance of models.
 In this section, we showcase the performance of methods equipped with our closed-form predictives and learning objectives against Softmax models.
-For the latter, we use the best-performing predictive and estimator (see Appendices E and F for an overview of predictive approximations and estimators).
+For the latter, we use the best-performing predictive and estimator (see Appendices˜E and F for an overview of predictive approximations and estimators).
 We employ our methods with the closed-form predictives and second-order Dirichlet distributions to demonstrate their competitive nature while being more efficient than MC sampling.
-We first evaluate how calibrated the models are using the test NLL [12] and Expected Calibration Error (ECE) [29] metrics.
-Fig.
-3(b) shows that on CIFAR-10, the score of our closed-form predictives (Sigmoid, NormCDF) is consistently better than the corresponding Softmax results for all methods.
-On the large-scale ImageNet dataset, Sigmoid and NormCDF predictives either outperform or are on par with Softmax (see Fig.
-3(a)).
+We first evaluate how calibrated the models are using the test NLL [gneiting2007strictly] and Expected Calibration Error (ECE) [naeini2015obtaining] metrics.
+Fig.˜3(b) shows that on CIFAR-10, the NLL score of our closed-form predictives (Sigmoid, NormCDF) is consistently better than the corresponding Softmax results for all methods.
+On the large-scale ImageNet dataset, Sigmoid and NormCDF predictives either outperform or are on par with Softmax (see Fig.˜3(a)).
+The ECE metric requires the models’ confidence (i.e., maximum class probability) to match their accuracy.
+Fig.˜4(a) shows that on ImageNet, our closed-form predictives have a clear advantage over Softmax across all methods.
 (a) Closed-form predictives ( ) are more calibrated than Softmax ( ) across all methods. Note the restricted $y$-limits for readability.
@@
 Figure 4: ImageNet ECE (left) and log probability proper scoring results for the binary correctness prediction task (right).
-The ECE metric requires the models’ confidence (i.e., maximum class probability) to match their accuracy.
-Fig.
-4(a) shows that, on ImageNet, our closed-form predictives have a clear advantage over Softmax across all methods.
 Next, we turn to the correctness prediction task: whether the models can predict the correctness of their own predictions.
@@
 We then measure the binary log probability score of $\tilde{t}(x)$, given by $v(x)\log t(x)+(1-v(x))\log(1-t(x))$.
-Fig.
-4(b) shows that our closed-form predictives outperform all Softmax predictives across all methods.
+Fig.˜4(b) shows that our closed-form predictives outperform all Softmax predictives across all methods.
 (a) Closed-form predictives do not sacrifice accuracy. Closed-form predictives ( ) either outperform or are on par with Softmax ( ) across all methods.
@@
 Closed-form predictives do not sacrifice accuracy.
-Fig.
-5(a) evidences this claim on ImageNet: our closed-form predictives either outperform or are on par with Softmax predictives.
-These results are in line with the findings of Wightman et al.
-[38], who also recommend training with a per-class binary cross-entropy loss using the sigmoid activation function.
-Section K.5 shows that the BCE loss already yields performance gains in the model backbones (without the logit-space Gaussians).
+Fig.˜5(a) evidences this claim on ImageNet: our closed-form predictives either outperform or are on par with Softmax predictives.
+These results are in line with the findings of wightman_resnet_2021, who also recommend training with a per-class binary cross-entropy loss using the sigmoid activation function.
+Section˜K.6 shows that the BCE loss already yields performance gains in the model backbones (without the logit-space Gaussians).
 Finally, we consider another binary prediction task, where a general uncertainty estimator $u(x)\in\mathbb{R}$ (derived from predictives or second-order Dirichlet distributions) is tasked to separate ID and OOD samples from a balanced mixture thereof.
@@
 the uncertainty estimator.
-As OOD inputs, we consider corrupted ImageNet-C [13] samples.
-Fig.
-5(b) shows that our closed-form predictives outperform Softmax across all methods.
-Importantly, these closed-form predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5).
-See Section K.6 for the other severity levels.
+As OOD inputs, we consider corrupted ImageNet-C [hendrycks2019benchmarking] samples.
+Fig.˜5(b) shows that our closed-form predictives outperform Softmax across all methods.
+Importantly, these closed-form predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section˜6).
+See Section˜K.7 for the other severity levels.
 8 Conclusion and Limitations
@@
 Appendix A Gaussian Integral Derivations
-In this appendix, we derive the closed-form formula for the mean of Gaussian pushforwards through exp (Eq.
-8) and normCDF (Eq.
-13), as well as the approximations for pushforwards through sigmoid (Eq.
-15) and softmax.
+In this appendix, we derive the closed-form formula for the mean of Gaussian pushforwards through exp (Eq.˜8) and normCDF (Eq.˜13), as well as the approximations for pushforwards through sigmoid (Eq.˜15) and softmax.
 A.1 Gaussian Exp Integral
 By absorbing the exponential into the Gaussian probability density function, we get
-[…]
-[…]
+$\int_{\mathbb{R}}\exp(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)$ $=\int_{\mathbb{R}}\exp(y)\exp\left(-\frac{(y-\mu)^{2}}{2\sigma^{2}}\right)\;dy$ (31)
+$=\int_{\mathbb{R}}\exp\left(\mu+\frac{\sigma^{2}}{2}\right)\exp\left(-\frac{1}{2\sigma^{2}}(y-\mu-\sigma^{2})^{2}\right)\;dy$
 $=\int_{\mathbb{R}}\exp\left(\mu+\frac{\sigma^{2}}{2}\right)\;\mathcal{N}(\mu+\sigma^{2},\sigma^{2})(dy)$
@@
 A.2 Gaussian NormCDF Integral
-Here, we derive the classical normCDF Gaussian integration formula [30, Eq. 10,010.8].
+Here, we derive the classical normCDF Gaussian integration formula [owen_table_1980, Eq. 10,010.8].
 For $\lambda>0$, $Z\sim\mathcal{N}(0,1)$ and $Y\sim\mathcal{N}(\mu,\sigma^{2})$ with $Y$ and $Z$ independent,
@@
 A.3 Probit Approximation for Gaussian Sigmoid Integral
-Taylor expanding $\rho$ to first order about 00,
+Taylor expanding $\rho$ to first order about 0,
 $\rho(y)$ $=\frac{1}{2}+\frac{1}{4}y+o(y),$ (33)
@@
 Hence matching $\rho$ and $\Phi$ to first order we get the approximation $\rho(y)\approx\Phi\left(\sqrt{\frac{\pi}{8}}y\right)$.
-So using Eq.
-31 we derive the probit approximation [34, 23]
+So using Eq.˜32 we derive the probit approximation [spiegelhalter_sequential_1990, mackay_evidence_1992]
 $\int_{\mathbb{R}}\rho(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)\stackrel{{\scriptstyle(1)}}{{\approx}}\int_{\mathbb{R}}\Phi\left(\sqrt{\frac{\pi}{8}}y\right)\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{\frac{8}{\pi}+\sigma^{2}}}\right)\stackrel{{\scriptstyle(2)}}{{\approx}}\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right).$ (34)
@@
 A.4 Mean Field Approximation for Gaussian Softmax Integral
-For $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, the mean field approximation to the Gaussian softmax integral [21] is obtained as follows
+For $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, the mean field approximation to the Gaussian softmax integral [lu_mean-field_2021] is obtained as follows
 $\mathbb{E}_{\bm{Y}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\operatorname{softmax}_{c}\bm{Y}]$ $=\mathbb{E}\left[\left(2-C+\sum_{c^{\prime}\neq c}\rho(Y_{c}-Y_{c^{\prime}})^{-1}\right)^{-1}\right]$ (35)
@@
 $\mathbb{E}_{\bm{Y}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\operatorname{\mathbf{softmax}}\bm{Y}]\approx\operatorname{\mathbf{softmax}}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right).$ (36)
-(1) is the mean field approximation, and (3) uses the probit approximation Eq. 33. [21] provides two other variants of this approximation with other choices of approximation (2).
+(1) is the mean field approximation, and (3) uses the probit approximation Eq.˜34. [lu_mean-field_2021] provides two other variants of this approximation with other choices of approximation (2).
 Appendix B Comparison with the Multinomial Probit Model
-In this appendix, we show that a model whose output activation is a composition of an element-wise normCDF activation $\bm{\Phi}$ and a normalisation $\bm{n}$ is distinct from the classical multinomial probit model [6].
+In this appendix, we show that a model whose output activation is a composition of an element-wise normCDF activation $\bm{\Phi}$ and a normalisation $\bm{n}$ is distinct from the classical multinomial probit model [daganzo_chapter_1979].
 Given a logit $\bm{y}\in\mathbb{R}^{C}$, the multinomial probit model sets
@@
 $p(c\mid y)=\frac{p(y_{c}+\epsilon_{c}>0)}{\sum_{c^{\prime}=1}^{C}p(y_{c^{\prime}}+\epsilon_{c^{\prime}}>0)}=\frac{\Phi(y_{c})}{\sum_{c^{\prime}=1}^{C}\Phi(y_{c}^{\prime})}.$ (39)
-In the case $C=2$, the multinomial probit model Eq. 37 outputs closed-form probabilities. This allows us to construct an explicit counterexample to the equivalence of the two models Eq. 37 and Eq. 38:
+In the case $C=2$, the multinomial probit model Eq.˜38 outputs closed-form probabilities. This allows us to construct an explicit counterexample to the equivalence of the two models Eq.˜38 and Eq.˜39:
 $p(y_{1}+\epsilon_{1}>y_{2}+\epsilon_{2})$ $=p\left(\frac{y_{1}-y_{2}}{2}+\frac{\epsilon_{1}-\epsilon_{2}}{2}>0\right)=\Phi\left(\frac{y_{1}-y_{2}}{2}\right)\neq\frac{\Phi(y_{1})}{\Phi(y_{1})+\Phi(y_{2})}.$ (40)
@@
 In this appendix, we provide (formal and informal) theoretical analyses of the quality of various predictive approximations.
-This complements the empirical analyses, for instance in the synthetic experiment (Fig.
-1) or in [7].
+This complements the empirical analyses, for instance in the synthetic experiment (Fig.˜1) or in [daunizeau_semi-analytical_2017].
 Due to its information-theoretic interpretation, a natural divergence to equip the probability simplex $\Delta^{C-1}$ with is the Kullback-Leibler (KL) divergence
 $\operatorname{D_{KL}}(\bm{p},\bm{q})=\sum_{c=1}^{C}p_{c}(\log p_{c}-\log q_{c})$ (41)
-which is well defined if $\bm{p}$, $\bm{q}$ lie in the interior of the simplex ($p_{i},q_{i}\neq 0,1$ for all $i$).
-So, for a predictive approximation, $\bm{\hat{p}}$, we would like to analyse $\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$, where $\bm{p}:=\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}]$ is the true predictive, $\bm{\mu}$ and $\bm{\Sigma}$ are some logit space mean and covariance and $\bm{a}$ is an output activation (e.g.
-$\bm{a}=\bm{n}\circ\bm{\varphi}$).
+[…]
 C.1 Informal Analysis of Monte Carlo Approximations
@@
 Thus, for a fair assessment of the quality of such an estimate in terms of the number of classes, one should consider MC estimates $\bm{\hat{P}}^{\lceil S/C\rceil}$.
-We now give an informal theoretical argument for the linear growth of the KL divergence between $\bm{p}$ and $\bm{\hat{P}}^{\lceil S/C\rceil}$ in terms of $C$, under the distributional conditions of the synthetic experiment (Fig.
-1).
-Taylor expanding Eq. 40 about $\bm{p}$ to second order we obtain
+We now give an informal theoretical argument for the linear growth of the KL divergence between $\bm{p}$ and $\bm{\hat{P}}^{\lceil S/C\rceil}$ in terms of $C$, under the distributional conditions of the synthetic experiment (Fig.˜1).
+Taylor expanding Eq.˜41 about $\bm{p}$ to second order we obtain
 $\operatorname{D_{KL}}(\bm{p},\bm{q})$ $=\sum_{c=1}^{C}p_{c}(\log p_{c}-\log q_{c})$ (43)
@@
 where approximation (1) assumes $\|\bm{p}-\bm{q}\|_{2}$ is small, and (2) assumes $p_{c}\approx 1/C$.
-In the synthetic experiment (Fig.
-1), the logit class-wise means $\mu_{c}$ and variances $\sigma_{c}^{2}$ are sampled in an i.i.d.
+In the synthetic experiment (Fig.˜1), the logit class-wise means $\mu_{c}$ and variances $\sigma_{c}^{2}$ are sampled in an i.i.d.
 way.
@@
 We have
-$\mathrm{Var}[\bm{P}]=\mathbb{E}\left[\frac{\bm{Q}^{2}}{\left(\sum_{c=1}^{C}Q_{c}\right)^{2}}\right]-\mathbb{E}\left[\frac{\bm{Q}}{\sum_{c=1}^{C}Q_{c}}\right]^{2}\approx\frac{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}{\left(\sum_{c=1}^{C}\mathbb{E}[Q_{c}]\right)^{2}}=\frac{\mathrm{Var}[\bm{Q}]}{C^{2}\mathbb{E}[Q_{1}]^{2}}$ (43)
-where all operations are taken element-wise. Thus
-$\mathbb{E}\left[\|\bm{p}-\bm{P}\|^{2}_{2}\right]=\sum_{c=1}^{C}\mathrm{Var}[P_{c}]\approx\sum_{c=1}^{C}\frac{\mathrm{Var}[Q_{c}]}{C^{2}\mathbb{E}[Q_{1}]^{2}}=\frac{\mathrm{Var}[Q_{1}]}{C\mathbb{E}[Q_{1}]^{2}}.$ (44)
+$\mathrm{Var}[\bm{P}]=\mathbb{E}\left[\frac{\bm{Q}^{2}}{\left(\sum_{c=1}^{C}Q_{c}\right)^{2}}\right]-\mathbb{E}\left[\frac{\bm{Q}}{\sum_{c=1}^{C}Q_{c}}\right]^{2}\approx\frac{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}{\left(\sum_{c=1}^{C}\mathbb{E}[Q_{c}]\right)^{2}}\approx\frac{\mathrm{Var}[\bm{Q}]}{C^{2}\mathbb{E}[Q_{1}]^{2}}$ (44)
+where all operations are taken element-wise, and $\mathbb{E}[Q_{c}]\approx\mathbb{E}[Q_{1}]$ follows from the fact that the $\mu_{c}$ and $\sigma_{c}^{2}$ are i.i.d. Thus
+$\mathbb{E}\left[\|\bm{p}-\bm{P}\|^{2}_{2}\right]=\sum_{c=1}^{C}\mathrm{Var}[P_{c}]\approx\sum_{c=1}^{C}\frac{\mathrm{Var}[Q_{c}]}{C^{2}\mathbb{E}[Q_{1}]^{2}}\approx\frac{\mathrm{Var}[Q_{1}]}{C\mathbb{E}[Q_{1}]^{2}}.$ (45)
 Now the MC samples $\bm{\hat{P}}^{(s)}$ are i.i.d. copies of $\bm{P}$. So we have
-$\mathbb{E}\big{[}\|\bm{p}-\bm{\hat{P}}^{\lceil S/C\rceil}\|^{2}_{2}\big{]}=\frac{1}{\lceil S/C\rceil}\mathbb{E}\left[\|\bm{p}-\bm{P}\|^{2}_{2}\right]\approx\frac{\mathrm{Var}[Q_{1}]}{S\mathbb{E}[Q_{1}]^{2}}.$ (45)
-Plugging this into Eq. 42 we get
-$\mathbb{E}[\operatorname{D_{KL}}(\bm{p},\bm{\hat{P}}^{\lceil S/C\rceil})]\approx\frac{\mathrm{Var}[Q_{1}]}{2\mathbb{E}[Q_{1}]^{2}}\cdot\frac{C}{S}$ (46)
-which grows linearly with the number of classes, as observed in Fig. 1.
-C.2 Analysis of the Closed-Form Approximations (Theorem 2.1)
-In this section we prove Theorem 2.1 and discuss whether its underlying assumptions are fulfilled in practice. We start with a lemma.
+$\mathbb{E}\big[\|\bm{p}-\bm{\hat{P}}^{\lceil S/C\rceil}\|^{2}_{2}\big]=\frac{1}{\lceil S/C\rceil}\mathbb{E}\left[\|\bm{p}-\bm{P}\|^{2}_{2}\right]\approx\mathbb{E}\left[\frac{\mathrm{Var}[Q_{1}]}{S\mathbb{E}[Q_{1}]^{2}}\right].$ (46)
+where the outer expectation is over the (third-order) distributions over $\bm{P}$. Plugging this into Eq.˜43 we get
+$\mathbb{E}[\operatorname{D_{KL}}(\bm{p},\bm{\hat{P}}^{\lceil S/C\rceil})]\approx\mathbb{E}\left[\frac{\mathrm{Var}[Q_{1}]}{2\mathbb{E}[Q_{1}]^{2}}\right]\cdot\frac{C}{S}$ (47)
+which grows linearly with the number of classes, as observed in Fig.˜1.
+C.2 Analysis of the Closed-Form Approximations (Theorem˜3.1)
+In this section we prove Theorem˜3.1 and discuss whether its underlying assumptions are fulfilled in practice. We start with a lemma.
 Lemma C.1.
@@
 $\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{4}]$ $\leq M^{2}\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{2}].$
-So from Eq. 49 we obtain
+So from Eq.˜50 we obtain
 $\left|\mathbb{E}\left[\frac{X_{n}}{Y_{n}}\right]-\frac{\mathbb{E}[X_{n}]}{\mathbb{E}[Y_{n}]}\right|=O(\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{2}]^{1/2})=O(\mathrm{Var}(Y_{n})^{1/2})$ (52)
@@
 $q=q(\mu,\sigma^{2}):=\mathbb{E}_{Q\sim\rho_{*}\mathcal{N}(\mu,\sigma^{2})}[Q]\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)=:\hat{q}(\mu,\sigma^{2})=\hat{q}$ (53)
-where $\hat{q}$ is the probit approximation (Section A.3).
+where $\hat{q}$ is the probit approximation (Section˜A.3).
 Moreover we write $q_{c}:=q(\mu_{c},\sigma_{c}^{2})$, $\hat{q}_{c}:=\hat{q}(\mu_{c},\sigma_{c}^{2})$, $\bm{p}$ for the true predictive, $\bm{\hat{p}}$ for our approximate predictive, and in addition $\bm{\tilde{p}}$ for the approximate predictive using the exact one dimensional integrals, i.e.
-$\bm{p}:=\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}],\quad\bm{\hat{p}}:=\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}},\quad\bm{\tilde{p}}:=\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}.$ (53)
-We now restate Theorem 2.1 with an explicit expression for $M(\mathcal{K})$:
+$\bm{p}:=\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}],\quad\bm{\hat{p}}:=\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}},\quad\bm{\tilde{p}}:=\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}},$ (54)
+where $\bm{\Sigma}$ has $\bm{\sigma}$ as diagonal. We now restate Theorem˜3.1 with an explicit expression for $M(\mathcal{K})$:
 Theorem C.2.
@@
 $\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})\leq\log\left(\frac{1+\delta/u}{1-\Delta}\right)+O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$ (55)
-as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$
+as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$.
 Proof.
@@
 $=\log\left(\frac{1+\delta/u}{1-\Delta}\right).$
-Now for (1), first note that $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$ where [^2] In the case $\varphi=\Phi$ and $Q_{c}\sim\Phi_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$, due to the fast decay of the tail of $\Phi$ we may have $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]=\infty$, making the proof strategy fail in that case.
+Now for (1), first note that $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$ where [^2] In the case $\varphi=\Phi$ and $Q_{c}\sim\Phi_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$, due to the fast decay of the tail of $\Phi$ we may have $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]=\infty$, and the proof strategy fails in that case.
 $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$.
 By compactness of $\mathcal{K}$, we have in fact $\sup_{(\bm{\mu},\bm{\sigma})\in\mathcal{K}^{C}}\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$.
-Thus we can apply Lemma C.1 with $(X_{n})_{n\geq 1}$ a sequence of $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$ and $(Y_{n})_{n\geq 1}$ a sequence of $\sum_{c=1}^{C}Q_{c}$, to get
-$|p_{c}-\tilde{p}_{c}|=O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)^{1/2}\right)$ (57)
+Thus we can apply Lemma˜C.1 with $(X_{n})_{n\geq 1}$ a sequence of $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$ and $(Y_{n})_{n\geq 1}$ a sequence of $\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}$, to get
+$|p_{c}-\tilde{p}_{c}|=O\left(\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{1/2}\right)$ (58)
 as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$. Thus Taylor expanding each term of (1) around $p_{c}$ we get
-$(1)$ $=\sum_{c=1}^{C}p_{c}\left(\frac{p_{c}-\tilde{p}_{c}}{p_{c}}+\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega(\tilde{p}_{c})^{2}}\right)$ (58)
-$=\underbrace{\sum_{c=1}^{C}p_{c}}_{=1}-\underbrace{\sum_{c=1}^{C}\tilde{p}_{c}}_{=1}+\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega(\tilde{p}_{c})^{2}}$
-$=\frac{O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)^{1/2}\right)^{2}}{2\omega(\tilde{p}_{c})^{2}}$
+$(1)$ $=\sum_{c=1}^{C}p_{c}\left(\frac{p_{c}-\tilde{p}_{c}}{p_{c}}+\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega_{c}(\tilde{p}_{c})^{2}}\right)$ (59)
+$=\underbrace{\sum_{c=1}^{C}p_{c}}_{=1}-\underbrace{\sum_{c=1}^{C}\tilde{p}_{c}}_{=1}+\sum_{c=1}^{C}p_{c}\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega_{c}(\tilde{p}_{c})^{2}}$
+$=\sum_{c=1}^{C}\frac{p_{c}}{2\omega_{c}(\tilde{p}_{c})^{2}}\cdot O\left(\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{1/2}\right)^{2}$
 $=O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$
-as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, where $\omega_{c}(\tilde{p}_{c})\in[p_{c},\tilde{p}_{c}]\text{ or }[\tilde{p}_{c},p_{c}]$.
+as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, where $\omega_{c}(\tilde{p}_{c})\in[p_{c},\tilde{p}_{c}]\text{ or }[\tilde{p}_{c},p_{c}]$, where we used that $p_{c}$, $\tilde{p}_{c}$, and hence $\omega_{c}(\tilde{p}_{c})$ is bounded away from 0 by compactness of $\mathcal{K}$.
 Figure C.1: Scatter plots of logit mean and variance pairs $(\mu_{c},\sigma_{c}^{2})$ on real models.
 We see that these seem to remain constrained to a compact set, with no much difference in the support as we increase the number of classes by an order of magnitude (from CIFAR-10 to CIFAR-100).
-A key assumption in Theorem C.2 is the logit means and variances being restricted to a compact set.
-In Fig.
-C.1 we see empirically that this is approximately the case for HET and SNGP models, independently of the number of classes.
-Theorem C.2 is of practical value as $M(\mathcal{K}):=\log\left(\frac{1+\delta/u}{1-\Delta}\right)$ is well-behaved:
+A key assumption in Theorem˜C.2 is the logit means and variances being restricted to a compact set.
+In Fig.˜C.1 we see empirically that this is approximately the case for HET and SNGP models, independently of the number of classes.
+Theorem˜C.2 is of practical value as $M(\mathcal{K}):=\log\left(\frac{1+\delta/u}{1-\Delta}\right)$ is well-behaved:
   1. 1.
@@
     $M(\mathcal{K})\to 0$ as $\delta\to 0$.
-In other words, given knowledge of the worst case error in the approximation Eq.
-15 on the compact set $\mathcal{K}$, we can bound the KL divergence in terms of that error independently of the number of classes.
+In other words, given knowledge of the worst case error in the approximation Eq.˜15 on the compact set $\mathcal{K}$, we can bound the KL divergence in terms of that error independently of the number of classes.
 Due to the simplicity of our assumptions, the bound remains quite raw and could be strengthened with further distributional assumptions on the means and variances.
-Finally, to obtain a meaningful bound in Theorem C.2, we needed to assume $\Delta(\mathcal{K}):=\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$.
-In Fig.
-C.2 we see that this can be assumed to hold on the compact sets on which logit means and variances tend to live (see also Fig.
-C.1).
+Finally, to obtain a meaningful bound in Theorem˜C.2, we needed to assume $\Delta(\mathcal{K}):=\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$.
+In Fig.˜C.2 we see that this can be assumed to hold on the compact sets on which logit means and variances tend to live (see also Fig.˜C.1).
 Figure C.2: Plot of the ‘true’ predictive $q$ (approximated with a $50000$ sample MC approximation) versus the probit approximation $\hat{q}$, as well as $\frac{q-\hat{q}}{q}$.
-We see that $\frac{q-\hat{q}}{q}$ tends to be negative, making the assumption $\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$ from Theorem C.2 easily fulfilled.
+We see that $\frac{q-\hat{q}}{q}$ tends to be negative, making the assumption $\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$ from Theorem˜C.2 easily fulfilled.
+Note that, in the proof of Theorem˜C.2, we did not use the form of the probit approximation explicitly, and the theorem holds for any approximation scheme $\hat{q}$ to the one dimensional integrals $q$.
 Appendix D Moment Matching Beta Distributions
-As noted in Section 3.2, when $\varphi=\Phi\text{ or }\rho$, we can construct a mapping
+As noted in Section˜4.2, when $\varphi=\Phi\text{ or }\rho$, we can construct a mapping
 $\mathrm{p}\colon\mathcal{G}(\mathbb{R}^{C})\to\mathcal{B}((0,1))^{C}$ (60)
@@
 where all vector operations are element-wise.
-Multiplying out the denominators on the right-hand side of the equations of Eq.
-60, we obtain a system of two linear equations with two unknowns (for each $c$), which can be solved uniquely, yielding
-$\bm{\alpha}$ $:=\frac{\mathbb{E}[\bm{Q}]-\mathbb{E}[\bm{Q}^{2}]}{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}\mathbb{E}[\bm{Q}],$ (61)
-$\bm{\beta}$ $:=\frac{\mathbb{E}[\bm{Q}]-\mathbb{E}[\bm{Q}^{2}]}{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}\left(1-\mathbb{E}[\bm{Q}]\right).$
+Multiplying out the denominators on the right-hand side of the equations of Eq.˜61, we obtain a system of two linear equations with two unknowns (for each $c$), which can be solved uniquely, yielding
+$\bm{\alpha}$ $=\frac{\mathbb{E}[\bm{Q}]-\mathbb{E}[\bm{Q}^{2}]}{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}\mathbb{E}[\bm{Q}],$ (62)
+$\bm{\beta}$ $=\frac{\mathbb{E}[\bm{Q}]-\mathbb{E}[\bm{Q}^{2}]}{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}\left(1-\mathbb{E}[\bm{Q}]\right).$
 which give us the parameters of the Beta distributions $\mathrm{p}(\mathrm{f})$.
@@
 NormCDF, unlike the sigmoid, makes the manifolds intersect at the point $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$.
-Here, we argue that the normCDF activation is an ideal choice for approximating Gaussian pushforwards with Beta distributions by interpreting Fig.
-D.1.
-This extends the work from [22], as it shows how one can make sense of the ‘right’ basis for performing Laplace approximations in the classification setting.
+Here, we argue that the normCDF activation is an ideal choice for approximating Gaussian pushforwards with Beta distributions by interpreting Fig.˜D.1.
+This extends the work from [mackay_choice_1998], as it shows how one can make sense of the ‘right’ basis for performing Laplace approximations in the classification setting.
 $\Phi_{*}\mathcal{G}(\mathbb{R})$, $\rho_{*}\mathcal{G}(\mathbb{R})$ the space of pushforwards of Gaussian distributions by normCDF and sigmoid respectively, and $\mathcal{B}((0,1))$, the space of Beta distributions, are statistical manifolds naturally equipped with Riemannian metrics, that is their respective Fisher information metrics.
@@
 The key observation is that $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$ are isometric to $\mathcal{G}(\mathbb{R})$.
-This is because $\Phi$ and $\rho$ are diffeomorphisms, so in particular sufficient statistics, and the Fisher information metric is invariant under sufficient statistics [2, Section 5.1.4].
+This is because $\Phi$ and $\rho$ are diffeomorphisms, so in particular sufficient statistics, and the Fisher information metric is invariant under sufficient statistics [ay_information_2017, Section 5.1.4].
 Visually, this means that $\varphi_{*}\mathcal{G}(\mathbb{R})$ has the same shape irrespectively of the diffeomorphism activation function $\varphi$.
@@
 Taking $\mu=0$, $\sigma^{2}=1$, $\alpha=\beta=1$ we get $\Phi_{0,1}=\Phi$ and $F_{1,1}=\operatorname{id}_{(0,1)}$, yielding $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}=\Phi$.
-Now $\mathcal{G}(\mathbb{R})$, and hence $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$, is isometric to the hyperbolic plane Ay et al.
-[2, Example 3.1].
-We can embed this isometrically into Minkowski space with the classical hyperboloid model of the hyperbolic plane [31],
+Now $\mathcal{G}(\mathbb{R})$, and hence $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$, is isometric to the hyperbolic plane ay_information_2017.
+We can embed this isometrically into Minkowski space with the classical hyperboloid model of the hyperbolic plane [reynolds_hyperbolic_1993],
 $\mathcal{G}(\mathbb{R})\hookrightarrow\mathbb{R}^{2,1}.$ (63)
-For $\mathcal{B}((0,1))$, we use the isometric embedding from Le Brigant et al. [18, Proposition 2]:
+For $\mathcal{B}((0,1))$, we use the isometric embedding from le_brigant_fisher-rao_2021:
 $\mathcal{B}((0,1))$ $\hookrightarrow\mathbb{R}^{2,1}$ (64)
 $(\alpha,\beta)$ $\mapsto(\eta(\alpha),\eta(\beta),\eta(\alpha+\beta))$
-[…]
-Finally, we choose our embedding Eq.
-62 for $\Phi_{*}\mathcal{G}(\mathbb{R})$ such that it intersects the embedding Eq.
-63 at a point, to highlight that the statistical manifolds $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ intersect at a point in the infinite-dimensional ambient space $\mathcal{M}((0,1))$, while $\rho_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ do not.
-Moreover, since $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$, our moment matching approximation Eq. 60 is exact when $\mathrm{f}(x)$ is the standard Normal distribution.
+where $\eta(a):=\int_{1}^{a}\sqrt{\psi^{\prime}(r)}\;dr$ and $\psi$ is the digamma function.
+Finally, we choose our embedding Eq.˜63 for $\Phi_{*}\mathcal{G}(\mathbb{R})$ such that it intersects the embedding Eq.˜64 at a point, to highlight that the statistical manifolds $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ intersect at a point in the infinite-dimensional ambient space $\mathcal{M}((0,1))$, while $\rho_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ do not.
+Moreover, since $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$, our moment matching approximation Eq.˜61 is exact when $\mathrm{f}(x)$ is the standard Normal distribution.
 Appendix E List of Predictive and Dirichlet Parameters Formulas
 E.1 Predictive Formulas
-We gather formulas for all predictive estimators $\bm{\hat{p}}$ of the true predictive $\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}]$, $\bm{\Sigma}=\bm{\operatorname{\mathbf{diag}}}(\bm{\sigma}^{2})$, used in our experiments (Section 6).
+We gather formulas for all predictive estimators $\bm{\hat{p}}$ of the true predictive $\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}]$, $\bm{\Sigma}=\bm{\operatorname{\mathbf{diag}}}(\bm{\sigma}^{2})$, used in our experiments (Section˜7).
 E.1.1 Softmax
@@
 One can Monte Carlo estimate the true predictive as follows:
-$\bm{\hat{p}}:=\frac{1}{S}\sum_{s=1}^{S}\bm{\hat{p}}^{(s)}$ (64)
+$\bm{\hat{p}}=\frac{1}{S}\sum_{s=1}^{S}\bm{\hat{p}}^{(s)}$ (65)
 where the $\bm{\hat{p}}^{(s)}$ are sampled i.i.d. from $\mathcal{N}(\bm{\mu},\bm{\Sigma})$.
-Mean Field (Section A.4)
-The Mean Field predictive [21] uses the following approximation for the true predictive:
+Mean Field (Section˜A.4)
+The Mean Field predictive [lu_mean-field_2021] uses the following approximation for the true predictive:
 $\bm{\hat{p}}:=\bm{\operatorname{\mathbf{softmax}}}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right).$ (66)
 Laplace Bridge
-The Laplace Bridge predictive [14] approximates the true predictive as follows:
+The Laplace Bridge predictive [pmlr-v180-hobbhahn22a] approximates the true predictive as follows:
 $\bm{\hat{p}}:=\frac{\frac{1}{\bm{\tilde{\sigma}}^{2}}\left(1-\frac{2}{C}+\frac{e^{\bm{\tilde{\mu}}}}{C^{2}}\sum_{c=1}^{C}e^{-\tilde{\mu}_{c}}\right)}{\sum_{c=1}^{C}\frac{1}{\tilde{\sigma}^{2}_{c}}\left(1-\frac{2}{C}+\frac{e^{\bm{\tilde{\mu}}}}{C^{2}}\sum_{c^{\prime}=1}^{C}e^{-\tilde{\mu}_{c^{\prime}}}\right)}$ (67)
@@
 $\bm{\tilde{\mu}}^{2}:=\sqrt{\frac{\sqrt{C/2}}{\sum_{c=1}^{C}\sigma^{2}_{c}}}\bm{\mu},\;\bm{\tilde{\sigma}^{2}}:=\frac{\sqrt{C/2}}{\sum_{c=1}^{C}\sigma^{2}_{c}}\bm{\sigma}^{2}.$ (68)
+Shekhovtsov & Flach.
+The shekhovtsov_feed-forward_2018 approximation of the Gaussian-softmax integral yields:
+$\bm{\hat{p}}=\frac{\bm{n}}{\bm{n}+\big(A-\bm{b}\big)^{s/\bm{s}}}$ (69)
+where
+$A:=\sum_{c=1}^{C}\exp(\tilde{\mu}_{c}/s),\;\bm{n}:=\exp(\bm{\tilde{\mu}}/\bm{s}),\;\bm{b}:=\exp(\bm{\tilde{\mu}}/s),$ (70)
+$\bm{\tilde{\mu}}:=\bm{\mu}-\max_{c}\mu_{c}$ for numerical stability,
+$\bm{\tau}^{2}:=\bm{\sigma}^{2}+G,\;\tau_{\text{min}}^{2}:=\min_{c}\tau_{c}^{2},\;s:=\sqrt{\tfrac{2\,\tau_{\text{min}}^{2}}{L}},\;\bm{s}:=\sqrt{\tfrac{\bm{\tau}^{2}+\tau_{\text{min}}^{2}}{L}}$ (71)
+and $L:=\pi^{2}/3$, $G:=L/2$.
 E.1.2 Exp
-Our closed-form predictive for the exp activation function (Section 2.1) is given by
+Our closed-form predictive for the exp activation function (Section˜3.1) is given by
 $\bm{\hat{p}}:=\frac{\bm{\exp}\left(\bm{\mu}+\frac{\bm{\sigma}^{2}}{2}\right)}{\sum_{c=1}^{C}\exp\left(\mu_{c}+\frac{\sigma^{2}_{c}}{2}\right)}.$ (72)
 E.1.3 NormCDF
-The closed-form predictive for the normCDF activation function (Section 2.2) is computed as
+The closed-form predictive for the normCDF activation function (Section˜3.2) is computed as
 $\bm{\hat{p}}:=\frac{\bm{\Phi}\left(\frac{\bm{\mu}}{\sqrt{1+\bm{\sigma}^{2}}}\right)}{\sum_{c=1}^{C}\Phi\left(\frac{\mu_{c}}{\sqrt{1+\sigma^{2}_{c}}}\right)}.$ (73)
 E.1.4 Sigmoid
-For the sigmoid activation function (Section 2.2), the closed-form predictive can be computed as
+For the sigmoid activation function (Section˜3.2), the closed-form predictive can be computed as
 $\bm{\hat{p}}:=\frac{\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)}{\sum_{c=1}^{C}\rho\left(\frac{\mu_{c}}{\sqrt{1+\frac{\pi}{8}\sigma^{2}_{c}}}\right)}.$ (74)
@@
 E.2.1 Softmax
-The Laplace Bridge method [14] uses the following Dirichlet parameters:
+The Laplace Bridge method [pmlr-v180-hobbhahn22a] uses the following Dirichlet parameters:
 $\bm{\gamma}:=\frac{1}{\bm{\tilde{\sigma}}^{2}}\left(1-\frac{2}{C}+\frac{e^{\bm{\tilde{\mu}}}}{C^{2}}\sum_{c=1}^{C}e^{-\tilde{\mu}_{c}}\right)$ (75)
@@
 E.2.2 Exp
-Exp uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
+Exp uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section˜4.1):
 $\bm{\gamma}:=\left(\prod_{c=1}^{C}\frac{a_{c}\cdot\max(\sum_{c^{\prime}=1}^{C}a_{c^{\prime}},1)-b_{c}}{b_{c}-a_{c}^{2}}\right)^{1/C}\frac{\bm{a}}{\sum_{c=1}^{C}a_{c}}$ (77)
@@
 E.2.3 NormCDF
-NormCDF uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
+NormCDF uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section˜4.1):
 $\bm{\gamma}:=\left(\prod_{c=1}^{C}\frac{a_{c}\cdot\max(\sum_{c^{\prime}=1}^{C}a_{c^{\prime}},1)-b_{c}}{b_{c}-a_{c}^{2}}\right)^{1/C}\frac{\bm{a}}{\sum_{c=1}^{C}a_{c}}$ (79)
@@
 E.2.4 Sigmoid
-Sigmoid uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
+Sigmoid uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section˜4.1):
 $\bm{\gamma}:=\left(\prod_{c=1}^{C}\frac{a_{c}\cdot\max(\sum_{c^{\prime}=1}^{C}a_{c^{\prime}},1)-b_{c}}{b_{c}-a_{c}^{2}}\right)^{1/C}\frac{\bm{a}}{\sum_{c=1}^{C}a_{c}}$ (81)
@@
 Appendix F List of Uncertainty Estimators
-In this section, we list the uncertainty estimators used in our experiments (Section 6).
+In this section, we list the uncertainty estimators used in our experiments (Section˜7).
 F.1 Predictive
@@
 Entropy
-$-\sum_{c=1}^{C}p_{c}\log p_{c}.$ (80)
+$\sum_{c=1}^{C}p_{c}\log p_{c}.$ (84)
 F.2 Monte Carlo
-Given $S$ Monte Carlo samples $\bm{\hat{p}}^{(1)},\dots,\bm{\hat{p}}^{(S)}$ with mean $\bm{\hat{p}}$, one can calculate a predictive as their average and derive the estimators in Section F.1.
+Given $S$ Monte Carlo samples $\bm{\hat{p}}^{(1)},\dots,\bm{\hat{p}}^{(S)}$ with mean $\bm{\hat{p}}$, one can calculate a predictive as their average and derive the estimators in Section˜F.1.
 However, Monte Carlo samples allow one to calculate two additional estimators detailed below.
 Expected Entropy
-$-\frac{1}{S}\sum_{s=1}^{S}\sum_{c=1}^{C}\hat{p}^{(s)}_{c}\log\hat{p}^{(s)}_{c}.$ (81)
+$\frac{1}{S}\sum_{s=1}^{S}\sum_{c=1}^{C}\hat{p}^{(s)}_{c}\log\hat{p}^{(s)}_{c}$. (85)
 Mutual Information/Jensen-Shannon Divergence
-$-\sum_{c=1}^{C}{\hat{p}}_{c}\log{\hat{p}}_{c}+\frac{1}{S}\sum_{s=1}^{S}\sum_{c=1}^{C}\hat{p}^{(s)}_{c}\log\hat{p}^{(s)}_{c}.$ (82)
+$\sum_{c=1}^{C}{\hat{p}}_{c}\log{\hat{p}}_{c}+\frac{1}{S}\sum_{s=1}^{S}\sum_{c=1}^{C}\hat{p}^{(s)}_{c}\log\hat{p}^{(s)}_{c}$. (86)
 F.3 Dirichlet
@@
 Expected Entropy
-$-\sum_{c=1}^{C}\frac{\gamma_{c}}{\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}}\left(\psi(\gamma_{c}+1)-\psi\left(\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}+1\right)\right)$ (83)
+$\sum_{c=1}^{C}\frac{\gamma_{c}}{\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}}\left(\psi(\gamma_{c}+1)-\psi\left(\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}+1\right)\right)$ (87)
 where $\psi$ is the digamma function.
@@
   • Do we have to sacrifice performance for sample-free predictives?
-To answer the first question, we evaluate our closed-form predictives (Sigmoid, NormCDF, Exp) and moment-matched Dirichlet distributions against softmax models equipped with approximate inference tools (Laplace Bridge [14], Mean Field [21], Monte Carlo sampling).
-We consider Heteroscedastic Classifiers (HET) [5], Spectral-Normalized Gaussian Processes (SNGP) [20], and last-layer Laplace approximation methods [8] as backbones (see Appendix I for details).
-The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [9] and CIFAR-10 [17] on five metrics aligning with practical needs from uncertainty estimates [26]:
+To answer the first question, we evaluate our closed-form predictives (Sigmoid, NormCDF, Exp) and moment-matched Dirichlet distributions against softmax models equipped with approximate inference tools (Laplace Bridge [pmlr-v180-hobbhahn22a], Mean Field [lu_mean-field_2021], Monte Carlo sampling).
+We consider Heteroscedastic Classifiers (HET) [collier2021correlated], Spectral-Normalised Gaussian Processes (SNGP) [liu2020simple], and last-layer Laplace approximation methods [daxberger2021laplace] as backbones (see Appendix˜J for details).
+The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [deng2009imagenet] and CIFAR-10 [krizhevsky2009learning] on five metrics aligning with practical needs from uncertainty estimates [mucsanyi2023trustworthy]:
   1. 1.
@@
     AUROC for the out-of-distribution (OOD) detection task.
-See Appendix H for details.
-For ImageNet, we treat ImageNet-C [13] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
+See Appendix˜I for details.
+For ImageNet, we treat ImageNet-C [hendrycks2019benchmarking] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
 For the second question, we consider fixed (method, activation) pairs and test whether our methods perform on par with the Monte Carlo sampled predictives.
 To provide a fair comparison, we reimplement each method as simple-to-use wrappers around deterministic backbones.
-For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following Mucsányi et al.
-[27].
-We train with the LAMB optimiser [40] using a batch size of $128$ and gradient accumulation across $16$ batches, resulting in an effective batch size of $2048$, following Tran et al.
-[36].
+For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following mucsányi2024benchmarking.
+We train with the LAMB optimiser [you2019large] using a batch size of $128$ and gradient accumulation across $16$ batches, resulting in an effective batch size of $2048$, following tran2022plex.
 We further use a cosine learning rate schedule with a single warmup epoch using a warmup learning rate of $0.0001$.
@@
 During training, we keep track of the best-performing checkpoint on the validation set and load it before testing.
-We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [33] in Weights & Biases [3] based on the negative log-likelihood.
+We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [shahriari2015taking] in Weights & Biases [wandb] based on the negative log-likelihood.
 On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs.
-The only exceptions are the SNGP models that are trained for 125 epochs [20].
+The only exceptions are the SNGP models that are trained for 125 epochs [liu2020simple].
 We train with Momentum SGD using a batch size of $128$ and no gradient accumulation.
@@
 The optimal learning rates for Sigmoid and NormCDF are generally larger, as the class-wise binary cross-entropies are averaged instead of summed.
-The weight decay is selected from the interval $[1\mathrm{e}{-6},1\mathrm{e}{-4}]$.
+The weight decay is selected from the interval $1\mathrm{e}{-6},1\mathrm{e}{-4}]$.
 Similarly to ImageNet, we use the best-performing checkpoint in the tests and use a ten-step Bayesian Optimization scheme to select performant hyperparameters.
@@
 The individual runs required no more than $50$ GB of RAM and $3$ days of runtime.
-Appendix H Benchmark Metrics
+Appendix H Runtime Overhead of Monte Carlo Sampling
+We measure the time to obtain the predictives from the logit-space Gaussians compared to the time of a forward pass on an NVIDIA RTX 2080 Ti GPU (with similar results on Tesla A100 GPUs).
+As shown in Table˜H.1, ResNet-50 with 1000 samples shows 6-8% overhead depending on the batch size, confirming our claim in the main text.
+In contrast, the runtime cost of computing our closed-form predictives is negligible compared to a forward pass, especially for larger batch sizes.
+These results demonstrate that even with a diagonal covariance structure, MC sampling yields a nontrivial overhead, especially for large numbers of classes and samples.
+In comparison, our closed-form approach has negligible computational cost.
+Table H.1: Runtime overhead of Monte Carlo sampling on ImageNet as a percentage of total inference time across models and batch sizes.
+Batch Size Mean Field 1 sample 10 samples 100 samples 1000 samples
+ResNet-26 (C=10)
+1 2.75% 9.82% 10.63% 10.70% 10.70%
+16 2.04% 10.61% 11.51% 11.51% 11.92%
+32 2.11% 10.98% 11.94% 11.96% 12.42%
+64 2.11% 10.88% 11.90% 11.95% 12.34%
+128 2.04% 8.12% 8.84% 8.96% 9.61%
+256 1.17% 4.31% 4.73% 4.93% 7.14%
+Wide-ResNet-26-5 (C=100)
+1 1.29% 8.54% 9.36% 9.40% 9.28%
+16 1.16% 5.94% 6.58% 6.84% 8.56%
+32 0.86% 3.36% 3.76% 3.98% 6.95%
+64 0.47% 1.83% 2.03% 2.12% 5.82%
+128 0.23% 0.90% 1.04% 1.04% 4.88%
+256 0.12% 0.46% 0.53% 0.84% 4.47%
+ResNet-50 (C=1000)
+1 1.15% 5.43% 5.88% 6.09% 6.30%
+16 0.32% 1.25% 1.43% 1.70% 8.04%
+32 0.16% 0.64% 0.72% 1.34% 7.51%
+64 0.09% 0.36% 0.41% 1.13% 7.75%
+128 0.05% 0.20% 0.22% 1.00% 7.82%
+256 0.03% 0.10% 0.18% 0.95% 8.05%
+Appendix I Benchmark Metrics
 Our experiments use five tasks/metrics:
@@
 Below, we describe these metrics and their respective tasks.
-H.1 Log Probability Proper Scoring Rule for the Predictive
-First, we briefly discuss proper and strictly proper scoring rules over general probability measures based on [26].
+I.1 Log Probability Proper Scoring Rule for the Predictive
+First, we briefly discuss proper and strictly proper scoring rules over general probability measures based on [mucsanyi2023trustworthy].
 Consider a function $S\colon\mathcal{Q}\times\mathcal{Y}\to\mathbb{R}$ where $\mathcal{Q}$ is a family of probability distributions over the space $\mathcal{Y}$, called the label space.
@@
 The score above is equivalent to the negative cross-entropy loss.
-H.2 Expected Calibration Error
-To set up the required quantities for the Expected Calibration Error (ECE) metric [29], we follow the steps below, based on [26].
+I.2 Expected Calibration Error
+To set up the required quantities for the Expected Calibration Error (ECE) metric [naeini2015obtaining], we follow the steps below, based on [mucsanyi2023trustworthy].
   1. 1.
@@
     $\tilde{c}_{n}\in\left(\frac{m-1}{M},\frac{m}{M}\right].$ (92)
-The Expected Calibration Error (ECE) metric [29] is then defined as
+The Expected Calibration Error (ECE) metric [naeini2015obtaining] is then defined as
 $\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\right|$ (93)
@@
 Intuitively, the ECE is high when the model’s per-bin confidences match its accuracy on the bin. We use $M=15$ bins in this paper.
-H.3 Binary Log Probability Proper Scoring Rule for Correctness Prediction
+I.3 Binary Log Probability Proper Scoring Rule for Correctness Prediction
 The correctness prediction task measures the models’ ability to predict the correctness of their own predictions.
@@
 Dropping the dependency on $x\in\mathcal{X}$ for brevity, the log probability score for binary targets $\ell\in\{0,1\}$ and estimators $\tilde{c}\in[0,1]$ is defined as
-$S(\tilde{c},\ell)=\begin{cases}\log c&\mathrm{if}\ \ell=1\\ \log(1-\tilde{c})&\mathrm{if}\ \ell=0\end{cases}=\ell\log\tilde{c}+(1-\ell)\log(1-\tilde{c}).$ (93)
-One can show that this is indeed a strictly proper scoring rule [26].
-H.4 Accuracy
+[…]
+One can show that this is indeed a strictly proper scoring rule [mucsanyi2023trustworthy].
+I.4 Accuracy
 For completeness, the accuracy of a predictive $h$ on a dataset $(x_{n},c_{n})_{n=1}^{N}$ is
 $\operatorname{acc}(h;(x_{n},c_{n})_{n=1}^{N})=\frac{1}{N}\sum_{n=1}^{N}1\left(\operatorname*{arg\,max}_{c\in\{1,\dotsc,C\}}h_{c}(x_{n})=\tilde{c}_{n}\right).$ (98)
-H.5 Area Under the Receiver Operating Characteristic Curve for Out-of-Distribution Detection
+I.5 Area Under the Receiver Operating Characteristic Curve for Out-of-Distribution Detection
 Out-of-distribution detection is another binary prediction task where a general uncertainty estimator $u(x)\in\mathbb{R}$ (derived from predictives or second-order Dirichlet distributions) is tasked to separate ID and OOD samples from a balanced mixture.
@@
 the uncertainty estimator.
-Given uncertainty estimates $u_{n}\equiv u(x_{n})$ and target binary labels $o_{i}$ on a balanced dataset $(x_{n},o_{n})_{n=1}^{N}$, as well as a threshold $t\in\mathbb{R}$, we predict $1$ (out-of-distribution) when $u_{n}\geq t$ and 00 (in-distribution) when $u_{n}<t$.
+Given uncertainty estimates $u_{n}\equiv u(x_{n})$ and target binary labels $o_{i}$ on a balanced dataset $(x_{n},o_{n})_{n=1}^{N}$, as well as a threshold $t\in\mathbb{R}$, we predict $1$ (out-of-distribution) when $u_{n}\geq t$ and 0 (in-distribution) when $u_{n}<t$.
 This lets us define the following index sets:
-[…]
-[…]
-[…]
-[…]
+$\text{True positives: }\mathrm{TP}(t)$ $=\left\{n:o_{n}=1\land u_{n}\geq t\right\}$ (99)
+$\text{False positives: }\mathrm{FP}(t)$ $=\left\{n:o_{n}=0\land u_{n}\geq t\right\}$
+$\text{False negatives: }\mathrm{FN}(t)$ $=\left\{n:o_{n}=1\land u_{n}<t\right\}$
+$\text{True negatives: }\mathrm{TN}(t)$ $=\left\{n:o_{n}=0\land u_{n}<t\right\}.$
 The Receiver Operating Characteristic (ROC) curve compares the following quantities:
@@
 Here, FPR tells us how many of the actual negative samples in the dataset are recalled (predicted positive) at threshold $t$.
-One can draw a curve of $(\mathrm{FPR}(t),\mathrm{TPR}(t))$ for all $t$ from $\infty$- ∞ to +$\infty$+ ∞.
+One can draw a curve of $(\mathrm{FPR}(t),\mathrm{TPR}(t))$ for all $t$ from -$\infty$ to $\infty$.
 This is the ROC curve.
 The area under this curve quantifies how well the uncertainty estimator $u(x)$ can separate in-distribution and out-of-distribution inputs.
-Appendix I Benchmarked Methods
+Appendix J Benchmarked Methods
 This section describes our benchmarked methods and provides further implementation details.
-I.1 Spectral Normalized Gaussian Process
-Spectral normalized Gaussian processes (SNGP) [20] use spectral normalization of the parameter tensors for distance-awareness and a last-layer Gaussian process approximated by Fourier features to capture uncertainty.
+J.1 Spectral Normalised Gaussian Process
+Spectral normalised Gaussian processes (SNGP) [liu2020simple] use spectral normalization of the parameter tensors for distance-awareness and a last-layer Gaussian process approximated by Fourier features to capture uncertainty.
 For an input $x\in\mathcal{X}$ and number of classes $C$, they predict a $C$-variate Gaussian distribution
@@
 The method applies spectral normalization to the hidden weights in each layer using a power iteration scheme with a single iteration per batch to obtain the largest singular value.
-Liu et al.
-[20] claim this helps with input distance awareness.
-I.2 Heteroscedastic Classifier
-Heteroscedastic classifiers (HET) [5] construct a Gaussian distribution in the logit space to model per-input uncertainties:
+liu2020simple claim this helps with input distance awareness.
+J.2 Heteroscedastic Classifier
+Heteroscedastic classifiers (HET) [collier2021correlated] construct a Gaussian distribution in the logit space to model per-input uncertainties:
 $\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x)),$ (102)
@@
 is a (positive definite) covariance matrix. Both the low-rank term $\bm{V}(x)$ and the diagonal term $\bm{d}(x)$ are calculated as a linear function of the pre-logit layer’s output.
-To learn the per-input covariance matrices from the training set, one has to construct a predictive estimate from $\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x))$ using any of the methods in Appendix E.
+To learn the per-input covariance matrices from the training set, one has to construct a predictive estimate from $\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x))$ using any of the methods in Appendix˜E.
 This predictive estimate is then trained using a standard cross-entropy (NLL) loss.
 HET uses a temperature parameter to scale the logits before calculating the BMA. This is chosen using a validation set.
-The off-diagonal terms of the covariance matrix do not affect the approximate predictive (Eq.
-12).
+The off-diagonal terms of the covariance matrix do not affect the approximate predictive (Eq.˜12).
 This means that, in our framework, one can discard the low-rank term $\bm{V}(x)$ and only model the diagonal term $\bm{d}(x)$ without a decrease in expressivity.
 To keep comparisons fair and use the same backbone with the same number of parameters, we also only model $\bm{d}(x)$ for softmax-based predictives.
-I.3 Laplace Approximation
-The Laplace approximation [8] approximates the posterior $p(\bm{\theta}\mid\mathcal{D})$ over the network parameters $\bm{\theta}$ for a Gaussian prior $p(\bm{\theta})$ and likelihood defined by the network architecture by a Gaussian.
-In its simplest form, it uses the maximum a posteriori (MAP) weights $\bm{\theta}_{\text{MAP}}\in\mathbb{R}^{P}$ as the mean and the inverse Hessian of the regularised loss over the training set ~$\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})=\mathcal{L}(\bm{\theta};\mathcal{D})+\lambda\|\bm{\theta}\|_{2}^{2}$ evaluated at the MAP as the covariance matrix:
+J.3 Laplace Approximation
+The Laplace approximation [daxberger2021laplace] approximates the posterior $p(\bm{\theta}\mid\mathcal{D})$ over the network parameters $\bm{\theta}$ for a Gaussian prior $p(\bm{\theta})$ and likelihood defined by the network architecture by a Gaussian.
+In its simplest form, it uses the maximum a posteriori (MAP) weights $\bm{\theta}_{\text{MAP}}\in\mathbb{R}^{P}$ as the mean and the inverse Hessian of the regularised loss over the training set $\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})=\mathcal{L}(\bm{\theta};\mathcal{D})+\lambda\|\bm{\theta}\|_{2}^{2}$ evaluated at the MAP as the covariance matrix:
 $\mathcal{N}\left(\bm{\theta}_{\text{MAP}},\left(\left.\frac{\partial^{2}\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})}{\partial\theta_{i}\partial\theta_{j}}\right|_{\bm{\theta}_{\text{MAP}}}\right)^{-1}\right)=\mathcal{N}\left(\bm{\theta}_{\text{MAP}},\left(\left.\frac{\partial^{2}\mathcal{L}(\bm{\theta};\mathcal{D})}{\partial\theta_{i}\partial\theta_{j}}\right|_{\bm{\theta}_{\text{MAP}}}+\lambda\bm{I}_{P}\right)^{-1}\right).$ (104)
-This is a locally optimal post-hoc Gaussian approximation of the true posterior $p(\bm{\theta}\mid\mathcal{D})$ based on a second-order Taylor approximation. For details, see [35].
-For deep neural networks, the Hessian matrix is often replaced with the Generalized Gauss-Newton (GGN) matrix.
-The GGN is guaranteed to be positive semidefinite even for suboptimal weights and has efficient approximation schemes, such as Kronecker-Factored Approximate Curvature [24] or low-rank approximations.
+This is a locally optimal post-hoc Gaussian approximation of the true posterior $p(\bm{\theta}\mid\mathcal{D})$ based on a second-order Taylor approximation. For details, see [tatzel2025debiasing].
+For deep neural networks, the Hessian matrix is often replaced with the Generalised Gauss-Newton (GGN) matrix.
+The GGN is guaranteed to be positive semidefinite even for suboptimal weights and has efficient approximation schemes, such as Kronecker-Factored Approximate Curvature [martens2015optimizing] or low-rank approximations.
 Denoting our curvature estimate of choice by $\bm{G}$, the logit-space Gaussian is obtained by pushing forward the weight-space Gaussian measure through the linearised model around $\bm{\theta}_{\text{MAP}}$.
@@
 where $\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\in\mathbb{R}^{C\times P}$ is the model Jacobian matrix.
-We use a last-layer KFAC Laplace variant in our experiments and use the full training set for calculating the GGN instead of a mini-batch based on recent works on the bias in mini-batch estimates [35].
-Appendix J CIFAR-10 Experiments
-This appendix section repeats the experiments presented in the main paper on the CIFAR-10 dataset.
-For a detailed description of the experimental setup, refer to Appendix G.
-Appendix H describes the used tasks and metrics.
-As stated in the main paper, our two research questions are:
-  • What are the effects of changing the learning objective? (Section J.2)
-  • Do we have to sacrifice performance for sample-free predictives? (Section J.1)
-J.1 Quality of Sample-Free Predictives
-Table J.1: Comparison of ECE results for different predictives using a fixed Laplace backbone.
-Method Mean $\pm$ 2 std
-Softmax Laplace
-MC 100 0.0096 $\pm$ 0.0013
-MC 1000 0.0102 $\pm$ 0.0016
-MC 10 0.0120 $\pm$ 0.0013
-Mean Field 0.0121 $\pm$ 0.0029
-Laplace Bridge Predictive 0.5933 $\pm$ 0.0105
-NormCDF Laplace
-Closed-form 0.0074 $\pm$ 0.0012
-MC 1000 0.0092 $\pm$ 0.0022
-MC 100 0.0095 $\pm$ 0.0015
-MC 10 $0$$\pm$ 0.0020
-Similarly to the main paper, in this section, we investigate our first research question: whether there is a price to pay for sample-free predictives.
-Table J.1 showcases the two best-performing (activation, method) pairs on the ECE metric and the CIFAR-10 dataset: Softmax and NormCDF Laplace.
-Mean Field (MF) is a strong alternative for sample-free predictives, but it has no guarantees and can fall behind MC sampling (see also Fig.
-1).
-Empirically, our closed-form predictives always perform on par with MC sampling.
-J.2 Effects of Changing the Learning Objective
-As in the main paper, in this section, we use the best-performing predictive and estimator (see Appendix E) for softmax models and employ our methods with the closed-form predictives.
-J.2.1 Calibration and Proper Scoring
-We first evaluate calibration using the log probability scoring rule [12] and the Expected Calibration Error (ECE) metric [29].
-Fig.
-3(b) shows that on CIFAR-10, the score of our closed-form predictives (Sigmoid, NormCDF) are consistently better than the corresponding softmax results for all methods.
-Figure J.1: CIFAR-10 ECE results.
-Our closed-form predictives ( ) outperform Softmax ( ) on HET and SNGP.
-Laplace tunes its hyperparameters based on the ECE metric – NormCDF Laplace is the overall best method.
-Note the restricted $y$-limits for readability.
-Fig.
-J.1 shows that on CIFAR-10, our closed-form predictives have a clear advantage on HET and SNGP.
-Laplace is a post-hoc method that tunes its hyperparameters on the ECE metric, hence its enhanced performance.
-Our NormCDF predictive is on par with Softmax.
-J.2.2 Correctness Prediction
-Figure J.2: CIFAR-10 log probability proper scoring results for the binary correctness prediction task. Our closed-form predictives ( ) consistently outperform Softmax ( ) on all methods.
-Fig. J.2 shows that on the correctness prediction task, our closed-form predictives outperform all Softmax predictives across all methods, as measured by the log probability proper scoring rule.
-J.2.3 Accuracy
-Figure J.3: Closed-form predictives do not sacrifice accuracy. CIFAR-10 accuracies. Our closed-form predictives ( ) either outperform or are on par with Softmax ( ) across all methods.
-Closed-form predictives do not sacrifice accuracy.
-Fig.
-J.3 evidences this claim on CIFAR-10: our closed-form predictives either outperform or are on par with Softmax predictives.
-The most accurate method is Sigmoid HET.
-These results support the findings of Wightman et al.
-[38] that showcase desirable training dynamics of the class-wise cross-entropy loss.
-J.2.4 Out-of-Distribution Detection
-Figure J.4: CIFAR-10C OOD detection AUROC results for severity level one. Across all methods, the best-performing predictive is closed-form ( ).
-Finally, we consider the OOD detection task on a balanced mixture of ID (CIFAR-10) and OOD inputs.
-As OOD inputs, we consider corrupted CIFAR-10C samples.
-We use the AUROC metric to evaluate the methods’ performance.
-As shown in Fig.
-J.4, the best-performing method is Sigmoid SNGP, a closed-form method.
-Generally, Softmax performs on par with our closed-form predictives.
-Intuitively, separating ID and OOD samples does not require a fine-grained representation of uncertainty, unlike the ECE or proper scoring rules.
-Nevertheless, the closed-form predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5).
+We use a last-layer KFAC Laplace variant in our experiments and use the full training set for calculating the GGN instead of a mini-batch based on recent works on the bias in mini-batch estimates [tatzel2025debiasing].
 Appendix K Additional Results
 K.1 Closed-Form Softmax Predictives
-One can directly apply Eq.
-10 to a neural network trained with the softmax cross-entropy loss in Eq.
-4 to obtain closed-form predictives.
-However, we empirically found this approach to decrease performance, as the denominator of Eq.
-10 increased to the order of billions during training.
+One can directly apply Eq.˜10 to a neural network trained with the softmax cross-entropy loss in Eq.˜4 to obtain closed-form predictives.
+However, we empirically found this approach to decrease performance, as the denominator of Eq.˜10 increased to the order of billions during training.
 Considering a multivariate normal random variable $\bm{x}\sim\mathcal{N}(\bm{\mu},\Sigma)$ representing a logit distribution,
@@
 i.e., when $e^{\mu_{j}}$ is large for some $j\in\{1,\dots,C\}$, the variance of the denominator necessarily explodes, violating our assumption.
-To mitigate this issue, one may optimize the regularised cross-entropy loss
+To mitigate this issue, one may optimise the regularised cross-entropy loss
 $\mathcal{L}((x_{n},c_{n})_{n=1}^{N})=\tilde{\mathcal{L}}((x_{n},c_{n})_{n=1}^{N})+\lambda\sum_{n=1}^{N}\left(\sum_{c=1}^{C}\exp(f_{c}(x_{n}))-1\right)^{2}$ (107)
@@
 The latter formulation can be stably trained even at an ImageNet scale.
-However, there are fundamental limitations to the predictive given by Eq.
-10.
+However, there are fundamental limitations to the predictive given by Eq.˜10.
 Namely, for isotropic logit-space Gaussian distributions, $\nicefrac{{\sigma^{2}(x)}}{{2}}$ introduces a constant shift in the logits, under which the softmax activation function is invariant.
@@
 Further, we empirically found that the Laplace method’s predictives for models trained with the regularised cross-entropy loss were also approximately isotropic; thus, this integral approximation yielded diminishing returns.
-For completeness, we share the ImageNet and CIFAR-10 results of $\varphi=\exp$ in Tables K.1 and K.2.
+For completeness, we share the ImageNet and CIFAR-10 results of $\varphi=\exp$ in Tables˜K.1 and K.2.
 Table K.1: Performance metrics of $\varphi=\exp$ on the CIFAR-10 validation dataset. We report the mean and two standard deviations.
@@
 ECE $\phantom{-}0.041\pm 0.006$ $\phantom{-}0.029\pm 0.007$ $\phantom{-}0.045\pm 0.007$
-Log Prob. $-0.255\pm 0.005$ $-0.238\pm 0.038$ $-0.255\pm 0.016$
+Log Prob. $0.255\pm 0.005$ $0.238\pm 0.038$ $0.255\pm 0.016$
 Accuracy $\phantom{-}0.884\pm 0.013$ $\phantom{-}0.888\pm 0.023$ $\phantom{-}0.893\pm 0.007$
@@
 ECE $\phantom{-}0.058\pm 0.004$ $\phantom{-}0.022\pm 0.007$ $\phantom{-}0.058\pm 0.013$
-Log Prob. $-0.380\pm 0.009$ $-0.403\pm 0.011$ $-0.402\pm 0.014$
+Log Prob. $0.380\pm 0.009$ $0.403\pm 0.011$ $0.402\pm 0.014$
 Accuracy $\phantom{-}0.784\pm 0.001$ $\phantom{-}0.779\pm 0.001$ $\phantom{-}0.785\pm 0.002$
@@
 K.2 Vision Transformer Results on ImageNet
-Table K.3 shows results on ViT Little [11] backbones from the timm [37] library on the ImageNet validation set.
-As in the main paper, the hyperparameters are optimized using a ten-step Bayesian hyperparameter optimization scheme of Weights and Biases.
+Table˜K.3 shows results on ViT Little [dosovitskiy2020image] backbones from the timm [rw2019timm] library on the ImageNet validation set.
+As in the main paper, the hyperparameters are optimised using a ten-step Bayesian hyperparameter optimization scheme of Weights and Biases.
 Table K.3: NLL and accuracy metrics of ViT Little models on the ImageNet validation set. Error bars represent two standard deviations.
@@
 Sigmoid $0.81178\pm 0.0115$ $79.980\pm 0.412$
-K.3 CIFAR-100 Results
-We repeat the NLL and accuracy evaluation on CIFAR-100 using WideResNet-28-5 models following a ten-step Bayesian hyperparameter sweep on Weights and Biases. Results are shown in Table K.4.
+K.3 CIFAR-10 Results
+This appendix section repeats the experiments presented in the main paper on the CIFAR-10 dataset.
+For a detailed description of the experimental setup, refer to Appendix˜G.
+Appendix˜I describes the used tasks and metrics.
+As stated in the main paper, our two research questions are:
+  • What are the effects of changing the learning objective? (Section˜K.3.2)
+  • Do we have to sacrifice performance for sample-free predictives? (Section˜K.3.1)
+K.3.1 Quality of Sample-Free Predictives
+Table K.4: Comparison of ECE results for different predictives using a fixed Laplace backbone.
+Method Mean $\pm$ 2 std
+Softmax Laplace
+MC 100 0.0096 $\pm$ 0.0013
+MC 1000 0.0102 $\pm$ 0.0016
+MC 10 0.0120 $\pm$ 0.0013
+Mean Field 0.0121 $\pm$ 0.0029
+Laplace Bridge Predictive 0.5933 $\pm$ 0.0105
+NormCDF Laplace
+Closed-form 0.0074 $\pm$ 0.0012
+MC 1000 0.0092 $\pm$ 0.0022
+MC 100 0.0095 $\pm$ 0.0015
+MC 10 0.0100 $\pm$ 0.0020
+Similarly to the main paper, in this section, we investigate our first research question: whether there is a price to pay for sample-free predictives.
+Table˜K.4 showcases the two best-performing (activation, method) pairs on the ECE metric and the CIFAR-10 dataset: Softmax and NormCDF Laplace.
+Mean Field (MF) is a strong alternative for sample-free predictives, but it has no guarantees and can fall behind MC sampling (see also Fig.˜1).
+Empirically, our closed-form predictives always perform on par with MC sampling.
+K.3.2 Effects of Changing the Learning Objective
+As in the main paper, in this section, we use the best-performing predictive and estimator (see Appendix˜E) for softmax models and employ our methods with the closed-form predictives.
+Calibration and Proper Scoring
+We first evaluate calibration using the log probability scoring rule [gneiting2007strictly] and the Expected Calibration Error (ECE) metric [naeini2015obtaining].
+Fig.˜3(b) shows that on CIFAR-10, the score of our closed-form predictives (Sigmoid, NormCDF) are consistently better than the corresponding softmax results for all methods.
+Figure K.1: CIFAR-10 ECE results.
+Our closed-form predictives ( ) outperform Softmax ( ) on HET and SNGP.
+Laplace tunes its hyperparameters based on the ECE metric – NormCDF Laplace is the overall best method.
+Note the restricted $y$-limits for readability.
+Fig.˜K.1 shows that on CIFAR-10, our closed-form predictives have a clear advantage on HET and SNGP.
+Laplace is a post-hoc method that tunes its hyperparameters on the ECE metric, hence its enhanced performance.
+Our NormCDF predictive is on par with Softmax.
+Correctness Prediction
+Figure K.2: CIFAR-10 log probability proper scoring results for the binary correctness prediction task. Our closed-form predictives ( ) consistently outperform Softmax ( ) on all methods.
+Fig.˜K.2 shows that on the correctness prediction task, our closed-form predictives outperform all Softmax predictives across all methods, as measured by the log probability proper scoring rule.
+Accuracy
+Figure K.3: Closed-form predictives do not sacrifice accuracy. CIFAR-10 accuracies. Our closed-form predictives ( ) either outperform or are on par with Softmax ( ) across all methods.
+Closed-form predictives do not sacrifice accuracy.
+Fig.˜K.3 evidences this claim on CIFAR-10: our closed-form predictives either outperform or are on par with Softmax predictives.
+The most accurate method is Sigmoid HET.
+These results support the findings of wightman_resnet_2021 that showcase desirable training dynamics of the class-wise cross-entropy loss.
+Out-of-Distribution Detection
+Figure K.4: CIFAR-10C OOD detection AUROC results for severity level one. Across all methods, the best-performing predictive is closed-form ( ).
+Finally, we consider the OOD detection task on a balanced mixture of ID (CIFAR-10) and OOD inputs.
+As OOD inputs, we consider corrupted CIFAR-10C samples.
+We use the AUROC metric to evaluate the methods’ performance.
+As shown in Fig.˜K.4, the best-performing method is Sigmoid SNGP, a closed-form method.
+Generally, Softmax performs on par with our closed-form predictives.
+Intuitively, separating ID and OOD samples does not require a fine-grained representation of uncertainty, unlike the ECE or proper scoring rules.
+Nevertheless, the closed-form predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section˜6).
+K.4 CIFAR-100 Results
+We repeat the NLL and accuracy evaluation on CIFAR-100 using WideResNet-28-5 models following a ten-step Bayesian hyperparameter sweep on Weights and Biases. Results are shown in Table˜K.5.
 Table K.5: NLL and accuracy metrics of WideResNet-28-5 models on the CIFAR-100 test set. Error bars represent two standard deviations.
@@
 Sigmoid SNGP $0.99996\pm 0.0265$ $78.350\pm 0.215$
-K.4 Alignment with the True Predictives
-Table K.5 shows the divergence of predictives using various approximation techniques from the true predictive estimated using $10,000$ Monte Carlo samples.
+K.5 Alignment with the True Predictives
+Table˜K.6 shows the divergence of predictives using various approximation techniques from the true predictive estimated using $10,000$ Monte Carlo samples.
 Our closed-form predictives, NormCDF and Sigmoid, are favorable to both the mean field and Laplace bridge approximations.
@@
 Softmax Laplace bridge $1.7900\pm 0.1254$
-K.5 BCE Loss Performance Gains
-Table K.6 shows that the Softmax and Sigmoid activation functions (and corresponding losses) already show improved performance on the vanilla models compared to softmax, highlighting that the performance gains we observe cannot only be attributed to our closed-form predictive approximations but also the favorable training dynamics of the class-wise BCE loss.
+K.6 BCE Loss Performance Gains
+Table˜K.7 shows that the Softmax and Sigmoid activation functions (and corresponding losses) already show improved performance on the vanilla models compared to softmax, highlighting that the performance gains we observe cannot only be attributed to our closed-form predictive approximations but also the favorable training dynamics of the class-wise BCE loss.
 Table K.7: NLL results on ImageNet using ResNet-50 backbones and different activation functions. Error bars represent two standard deviations.
@@
 Sigmoid $0.9146\pm 0.0046$
-K.6 Further Out-of-Distribution Detection Results
-Fig. K.1 shows OOD detection results across all ImageNet-C severity levels. Our closed-form predictives consistently outperform Softmax.
+K.7 Further Out-of-Distribution Detection Results
+Fig.˜K.5 shows OOD detection results across all ImageNet-C severity levels. Our closed-form predictives consistently outperform Softmax.
 Figure K.5: The OOD detection performance of all methods increases steadily as we increase the severity of the perturbed half of the mixed dataset on the ImageNet validation dataset.
 Our closed-form predictives consistently outperform Softmax.
+K.8 Comparison to Evidential Deep Learning Methods
+Table˜K.8 provides results for two evidential deep learning methods, EDL [sensoy2018evidential] and PostNet [charpentier2020posterior].
+(We adopt the naming convention of [mucsányi2024benchmarking].) We make the comparison because these methods also yield closed-form predictives and second-order quantities of interest, as they build Dirichlet distributions whose advantages we discuss in Section˜4.1.
+Following our experimental setup (Section 6), we optimise the hyperparameters of both methods using a ten-step Bayesian Optimization scheme in Weights & Biases.
+Compared with the results presented in the main text, our approach outperforms the two EDL methods across all metrics, significantly so on NLL and OOD AUROC.
+The results on the latter metric indicate that our moment-matched Dirichlet distributions capture second-order epistemic information more effectively than the loss-based EDL methods, which aligns with the findings of NEURIPS2022_bc1d640f.
+Table K.8: Evaluation results of PostNet and EDL across CIFAR-10, CIFAR-100, and ImageNet. Values are mean $\pm$ two standard deviations.
+Dataset Metric PostNet EDL
+ImageNet
+NLL $1.10202\pm 0.052$ $1.25544\pm 0.076$
+Accuracy $75.602\pm 0.627$ $75.632\pm 0.326$
+ECE $058168\pm 0.002 0.0$$033219\pm 0.001$
+OOD AUROC $0.60675\pm 0.013$ $0.60167\pm 0.018$
+CIFAR-100
+NLL $1.17973\pm 0.081$ $1.3645\pm 0.107$
+Accuracy $75.183\pm 0.192$ $75.383\pm 0.273$
+ECE $12295\pm 0.012 0.0$$088213\pm 0.005$
+OOD AUROC $0.59152\pm 0.017$ $0.59471\pm 0.02$
+CIFAR-10
+NLL $0.37406\pm 0.013$ $0.40403\pm 0.023$
+Accuracy $91.757\pm 0.021$ $90.274\pm 0.043$
+ECE $046192\pm 0.002 0.0$$041158\pm 0.002$
+OOD AUROC $0.5931\pm 0.014$ $0.60808\pm 0.006$
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The claims of improved uncertainty quantification capabilities are thoroughly verified in Sections˜7, K.3, K.4 and K.7.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: The paper’s limitations are discussed in Section˜8.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: All theoretical claims include the full set of assumptions and are proved in the appendix.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: The paper accurately describes the experimental setup in Sections˜7 and G and the code is provided as supplementary material.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: See above.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The main training, implementation, and hyperparameter optimization details are discussed in Sections˜7, G and J.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We run each experiment on five different seeds and report the min, mean, and max metric values.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed-form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: The compute resources and time of execution are discussed at the end of Appendix˜G.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We fulfill all requirements.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Our paper conducts no societally harmful research.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: There is no risk in abusing ImageNet, CIFAR-100, or CIFAR-10 classifiers.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We cite the ImageNet, ImageNet-ReaL, CIFAR-10, CIFAR-100, and CIFAR-10H datasets and the timm library for the code and the pretrained ResNet-50s.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: We do not release new assets; we only record metrics.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: We do not use crowdsourcing.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: We do not conduct experiments with humans.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The paper involves no use of LLMs.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v4

@@
 Notably, the output probability does not take into account the uncertainty of the model due to the finite nature of the data.
-In the uncertainty quantification formalism, such models can only estimate aleatoric uncertainty and disregard epistemic uncertainty [hullermeier2021aleatoric].
+In the uncertainty quantification formalism, such models can only estimate aleatoric uncertainty and disregard epistemic uncertainty [21].
 Figure 1: Mean KL divergence between the ‘true’ predictive (approximated with a $10000$ sample MC approximation) and different predictive approximations on a synthetic data set.
@@
 The probabilistic way to capture epistemic uncertainty is to require the model to output a second-order distribution (that is ‘a distribution over probability distributions’).
-Namely, for each input $x\in\mathcal{X}$, the model should output a probability measure over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$, as opposed to a point estimate in the simplex $\bm{h}(x)\in\Delta^{C-1}$ [sensoy2018evidential, charpentier2020posterior].
+Namely, for each input $x\in\mathcal{X}$, the model should output a probability measure over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$, as opposed to a point estimate in the simplex $\bm{h}(x)\in\Delta^{C-1}$ [44, 6].
 A number of methods for such distributional uncertainty quantification in classification rely on approximate Gaussian inference in logit space. That is, $\mathrm{h}$ is written as a composition
@@
 where $\mathrm{f}$ is learned from the training data, $\mathcal{G}(\mathbb{R}^{C})$ is the set of Gaussian measures on the logit space $\mathbb{R}^{C}$, and $\operatorname{\mathbf{softmax}}_{*}$ pushes forward the Gaussian probability measures through the softmax.
-For example, Heteroscedastic Classifiers (HET) [collier2021correlated] learn the output means and covariances of $\mathrm{f}$ explicitly with a neural network.
-In linearised Laplace approximations [daxberger2021laplace], only the means are learned explicitly with a neural network, while the covariances are obtained post-hoc from that neural network by approximating its parameter posterior distribution with a Gaussian, and then pushing it forward to logit space by linearising the network in the parameters.
-Last-layer Laplace approximations [kristiadi_being_2020] are a variant of linearised Laplace approximations where only the posterior over the last-layer parameters is Gaussian-approximated, requiring no linearisation of the network.
-Spectral-Normalised Gaussian processes (SNGP) [liu2020simple] are neural networks with last layers that are approximately Gaussian processes, through random feature expansions and Laplace approximations.
+For example, Heteroscedastic Classifiers (HET) [7] learn the output means and covariances of $\mathrm{f}$ explicitly with a neural network.
+In linearised Laplace approximations [11], only the means are learned explicitly with a neural network, while the covariances are obtained post-hoc from that neural network by approximating its parameter posterior distribution with a Gaussian, and then pushing it forward to logit space by linearising the network in the parameters.
+Last-layer Laplace approximations [24] are a variant of linearised Laplace approximations where only the posterior over the last-layer parameters is Gaussian-approximated, requiring no linearisation of the network.
+Spectral-Normalised Gaussian processes (SNGP) [30] are neural networks with last layers that are approximately Gaussian processes, through random feature expansions and Laplace approximations.
 To obtain predictive probabilities from a distributional classifier as in Eq.˜5, we marginalise it out by the measure-theoretic change of variables, using $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$:
 $\mathbb{E}_{\bm{P}\sim\mathrm{h}(x)}[\bm{P}]$ $=\int_{\Delta^{C-1}}p\;\mathrm{h}(x)(dp)=\int_{\mathbb{R}^{C}}\operatorname{\mathbf{softmax}}(y)\;\mathrm{f}(x)(dy).$ (6)
-In addition to predictive probabilities, the distributional framework allows for the acquisition of other quantities of interest, such as variances, entropy, and information-theoretic decompositions of the predictive uncertainty into aleatoric and epistemic parts [depeweg2018decomposition, mukhoti2023deep, wimmer2023quantifying].
+In addition to predictive probabilities, the distributional framework allows for the acquisition of other quantities of interest, such as variances, entropy, and information-theoretic decompositions of the predictive uncertainty into aleatoric and epistemic parts [13, 39, 53].
 These are usually obtained by Monte Carlo (MC) sampling from the Gaussian measure $\mathrm{f}(x)$.
-Indeed, while the probability density of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ can be obtained in closed form using a change of variables formula [atchison_logistic-normal_1980], this does not seem of practical use—even predictives cannot be computed analytically.
+Indeed, while the probability density of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ can be obtained in closed form using a change of variables formula [2], this does not seem of practical use—even predictives cannot be computed analytically.
 On the other hand, given a variance, the runtime and memory cost required for an MC approximation to achieve an estimator of said variance grows linearly with the number of classes.
@@
 This limits the application of such Gaussian inference methods in classification with a large number of classes, such as those common in computer vision and language modelling.
-Summary of Contributions: Our work extends the classical probit approximation for the Gaussian sigmoid integral in binary classification [spiegelhalter_sequential_1990, mackay_evidence_1992] in the following manner:
+Summary of Contributions: Our work extends the classical probit approximation for the Gaussian sigmoid integral in binary classification [47, 32] in the following manner:
   1. (i)
@@
 (ii) Choose an output activation that yields a closed-form Gaussian integral (Section˜3).
-(iii) At inference time, use our closed-form formulas (see Section˜3 for the predictives, Section˜4 for other quantities of interest).
+(iii) At inference time, use our closed-form formulas (Section˜3 for the predictives, Section˜4 for other quantities of interest).
 2 Related Work
-In uncertainty quantification, methods that output logit Gaussians, which we refer to as approximate Gaussian inference methods, have become one of the dominant paradigms in uncertainty quantification [collier2021correlated, daxberger2021laplace, kristiadi_being_2020, liu2020simple].
+In uncertainty quantification, methods that output logit Gaussians, which we refer to as approximate Gaussian inference methods, have become one of the dominant paradigms in uncertainty quantification [7, 11, 24, 30].
 In classification tasks, such methods have the downside of requiring sampling at inference to transform Gaussian distributions into predictive probability vectors.
-In parallel, prior work has developed ways to propagate probabilistic uncertainty through neural networks [gast_lightweight_2018].
-Several works develop analytic propagation of covariances up to the logits [wright_analytic_2024, li_streamlining_2024]; our method complements these by addressing the propagation of the logit-space Gaussian distributions onto the probability simplex.
+In parallel, prior work has developed ways to propagate probabilistic uncertainty through neural networks [16].
+Several works develop analytic propagation of covariances up to the logits [54, 28]; our method complements these by addressing the propagation of the logit-space Gaussian distributions onto the probability simplex.
 Few works address this last step.
-The Laplace bridge [pmlr-v180-hobbhahn22a] proposes a map from Gaussians to Dirichlets on the probability simplex.
+The Laplace bridge [19] proposes a map from Gaussians to Dirichlets on the probability simplex.
 However, this approach has no guarantees on the quality of its predictives, and it severely underperforms Monte Carlo sampling.
-Other approaches include the mean-field approximation to the Gaussian softmax integral [lu_mean-field_2021] and the approximations of shekhovtsov_feed-forward_2018.
+Other approaches include the mean-field approximation to the Gaussian softmax integral [31] and the approximations of Shekhovtsov and Flach [46].
 In this work, we propose approximate predictives that outperform existing closed-form predictive techniques and provide second-order Dirichlet distributions for other closed-form quantities of interest.
@@
 Other works have also proposed various activations for classification to replace the softmax.
-Particularly relevant is [galy-fajou_multi-class_2020], which composes element-wise sigmoids with normalisation and utilises auxiliary-variable augmentation to obtain closed-form ELBO terms during training; however, its predictives require Monte Carlo estimation.
-Related conjugate/augmentation-based objectives exist for stick-breaking and tree-structured multi-class GPs [linderman_dependent_2015, achituve_gp-tree_2021], and for BNNs, a tempered-likelihood/Dirichlet perspective with a factorised-Gaussian surrogate has been explored [kapoor_uncertainty_2022].
+Particularly relevant is [15], which composes element-wise sigmoids with normalisation and utilises auxiliary-variable augmentation to obtain closed-form ELBO terms during training; however, its predictives require Monte Carlo estimation.
+Related conjugate/augmentation-based objectives exist for stick-breaking and tree-structured multi-class GPs [29, 1], and for BNNs, a tempered-likelihood/Dirichlet perspective with a factorised-Gaussian surrogate has been explored [23].
 Yet, all these approaches still require sampling or numerical quadrature at inference time.
-[spiegelhalter_sequential_1990, mackay_evidence_1992, kristiadi_being_2020] use the probit approximation in the binary case, which we generalise to the multi-class setting.
-[titsias_rc_aueb_one-vs-each_2016] approximates the softmax by providing lower bounds for point estimate training, but replacing the softmax with these bounds during training does not yield a normalised multi-class likelihood.
-The binary cross-entropy loss with the element-wise sigmoid activation has been reported to outperform cross-entropy in large-scale settings [wightman_resnet_2021, li_bce_2025].
+[47, 32, 24] use the probit approximation in the binary case, which we generalise to the multi-class setting.
+[49] approximates the softmax by providing lower bounds for point estimate training, but replacing the softmax with these bounds during training does not yield a normalised multi-class likelihood.
+The binary cross-entropy loss with the element-wise sigmoid activation has been reported to outperform cross-entropy in large-scale settings [51, 27].
 Our work is the first to use this advantageous loss while normalising the output to obtain proper multi-class probabilistic models with closed-form predictives from logit Gaussians.
-Evidential deep learning (EDL) approaches [sensoy2018evidential] offer sample-free outputs through Dirichlet distributions but have been shown to be unreliable for uncertainty estimation and disentanglement [NEURIPS2022_bc1d640f, jürgens2024epistemicuncertaintyfaithfullyrepresented, pandey2023learn, mucsányi2024benchmarking].
+Evidential deep learning (EDL) approaches [44] offer sample-free outputs through Dirichlet distributions but have been shown to be unreliable for uncertainty estimation and disentanglement [4, jürgens2024epistemicuncertaintyfaithfullyrepresented, 42, mucsányi2024benchmarking].
 We focus on making approximate Gaussian inference methods sample-free in a principled and performant manner; however, for completeness, we compare our approach with EDL methods in Section˜K.8.
-Finally, multiple authors have explored continuous approximations to categorical distributions, which enable gradient-based optimization for discrete stochastic models through sampling, such as the Gumbel-max trick [maddison_ast_2014, jang_categorical_2017, maddison_concrete_2017, huijben_review_2023].
+Finally, multiple authors have explored continuous approximations to categorical distributions, which enable gradient-based optimization for discrete stochastic models through sampling, such as the Gumbel-max trick [35, 22, 34, 20].
 These methods principally optimise a single categorical distribution, while the present work is concerned with constructing probability densities over the (continuous) simplex of such categorical distributions.
@@
 Suppose the means and variances $(\mu_{c},\sigma_{c}^{2})$ lie in some compact set $\mathcal{K}\subset\mathbb{R}\times[0,\infty)$ for each class $c$. Then
-$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})=M(\mathcal{K})+O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$ (19)
-as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, for some constant $M(\mathcal{K})\in[0,\infty]$ which only depends on $\mathcal{K}$, not on $C$, such that $M(\mathcal{K})<\infty$ when $\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$, and $M(\mathcal{K})=0$ when $q=\hat{q}$ for all $(\mu,\sigma^{2})\in\mathcal{K}$.
-Section˜C.2 provides a formula for $M(\mathcal{K})$ and a proof.
+$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})\leq M(\mathcal{K})+O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$ (19)
+as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$ and $M(\mathcal{K})\in[0,\infty]$ some constant which only depends on $\mathcal{K}$, not on $C$, such that $M(\mathcal{K})<\infty$ when $\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}<1$, and $M(\mathcal{K})=0$ when $q=\hat{q}$ for all $(\mu,\sigma^{2})\in\mathcal{K}$.
+Section˜C.2 and Theorem˜C.2 provide a formula for $M(\mathcal{K})$ and a proof.
 We also conduct an empirical investigation of whether the assumptions made in the theorem hold in practice, namely whether the logit means and variances lie in a compact set, and whether $M(\mathcal{K})<\infty$.
-The $O$ term in Eq.˜19 is due to the splitting of the expectation into the numerator and denominator in the derivation Eq.˜10, while $M(\mathcal{K})$ results from the error due to the probit approximation.
-Importantly, Theorem˜3.1 states that the latter does not increase with the number of classes.
+The term $M(\mathcal{K})$ in Eq.˜19 results from the error due to the probit approximation, while the $O$ term is due to the splitting of the expectation into the numerator and denominator in the derivation Eq.˜10.
+Importantly, Theorem˜3.1 states that the former does not increase with the number of classes.
+For the latter, requiring $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$ is weaker than requiring $\mathrm{Var}(Q_{c})\to 0$ for all $c$ (see Eq.˜63 in Section˜C.2); but we show in Section˜C.2 that assuming $\mathrm{Var}(Q_{c})\to 0$ for all $c$ gives almost twice as strong a decay rate in $\operatorname{D_{KL}}$.
+We include this additional result in Theorem˜C.2.
 4 Distributional Approximations of the Gaussian Pushforwards
@@
 $\int_{\mathbb{R}}\exp(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)=\exp(2\mu+2\sigma^{2})$ (21)
-and for $\Phi$ (owen_table_1980)
+and for $\Phi$ (Owen [41, Eq. 20,010.4])
 $\int_{\mathbb{R}}\Phi(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{1+\sigma^{2}}}\right)-2T\left(\frac{\mu}{\sqrt{1+\sigma^{2}}},\frac{1}{\sqrt{1+2\sigma^{2}}}\right),$ (22)
-where $T$ is Owen’s T function. For $\rho$, we have the following approximation (daunizeau_semi-analytical_2017)
+where $T$ is Owen’s T function. For $\rho$, we have the following approximation (Daunizeau [10, Equation 23])
 $\int_{\mathbb{R}}\rho(y)^{2}\;\mathcal{N}(\mu,\sigma^{2})(dy)\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)-\frac{\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\left(1-\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)\right)}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}.$ (23)
@@
 If we were to directly attempt to infer the parameters of the Dirichlet $\mathrm{h}(x)$ from Eqs.˜12 and 24, we would obtain an overparametrised system of equations with $2C$ equations and $C$ unknowns.
-Thus, we instead use a classical Dirichlet method of moments [minka_2000, Eqs.
+Thus, we instead use a classical Dirichlet method of moments [37, Eqs.
 19 & 23], which reduces these equations, as we will now describe.
@@
 The BCE loss is a valid choice of loss as it is a (negative) strictly proper scoring rule (see Appendix˜I).
-Moreover, it has been observed to provide advantageous training dynamics compared to the CE loss [rw2019timm, li_bce_2025].
+Moreover, it has been observed to provide advantageous training dynamics compared to the CE loss [52, 27].
 The choice of $\varphi=\Phi$ or $\rho$ not only fits our framework for closed-form predictives, but allows one to leverage such advantageous training dynamics (see experimental results in Section˜7).
@@
 MC 1000 0.0118 $\pm$ 0.0007
-We consider Heteroscedastic Classifiers (HET) [collier2021correlated], Spectral-Normalised Gaussian Processes (SNGP) [liu2020simple], and last-layer Laplace-Approximated Networks [daxberger2021laplace] as backbones supplying logit-space Gaussian distributions.
-The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [deng2009imagenet], CIFAR-10, and CIFAR-100 [krizhevsky2009learning] on five metrics aligning with practical needs from uncertainty estimates [mucsanyi2023trustworthy].
+We consider Heteroscedastic Classifiers (HET) [7], Spectral-Normalised Gaussian Processes (SNGP) [30], and last-layer Laplace-Approximated Networks [11] as backbones supplying logit-space Gaussian distributions.
+The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [12], CIFAR-10, and CIFAR-100 [25] on five metrics aligning with practical needs from uncertainty estimates [38].
 Further, we test the Out-of-Distribution (OOD) detection capabilities of the models on balanced mixtures of In-Distribution (ID) and OOD samples.
-For ImageNet, we treat ImageNet-C [hendrycks2019benchmarking] samples with 15 corruption types and 5 severity levels as OOD samples.
+For ImageNet, we treat ImageNet-C [18] samples with 15 corruption types and 5 severity levels as OOD samples.
 For CIFAR-10 and CIFAR-100, we use the CIFAR-10C corruptions.
-For the second question, we evaluate our closed-form predictives and moment-matched Dirichlets against softmax models with the Laplace Bridge [pmlr-v180-hobbhahn22a] and Mean Field [lu_mean-field_2021] approximations, the linear-time approximation by shekhovtsov_feed-forward_2018, and MC sampling; see Section˜7.2.
+For the second question, we evaluate our closed-form predictives and moment-matched Dirichlets against softmax models with the Laplace Bridge [19] and Mean Field [31] approximations, the linear-time approximation by Shekhovtsov and Flach [46], and MC sampling; see Section˜7.2.
 To provide a fair comparison, we reimplement each method as simple-to-use wrappers around deterministic backbones.
-For ImageNet evaluation, we use a ResNet-50 backbone from the timm library [rw2019timm] pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following mucsányi2024benchmarking.
-For Vision Transformer [dosovitskiy2020image] experiments, refer to Section˜K.2.
+For ImageNet evaluation, we use a ResNet-50 backbone from the timm library [52] pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following mucsányi2024benchmarking.
+For Vision Transformer [14] experiments, refer to Section˜K.2.
 On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs.
 The CIFAR-100 experiments train WideResNet-28-5 models for 200 epochs.
-We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [shahriari2015taking] in Weights & Biases [wandb].
+We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [45] in Weights & Biases [5].
 The main paper focuses on ImageNet results to highlight the scalability of our framework and only shows proper scoring results for CIFAR-10.
@@
 In this section, we investigate whether there is a price to pay for sample-free predictives and second-order distributions.
-To this end, we take the two best-performing methods on the Expected Calibration Error (ECE) metric [naeini2015obtaining]—NormCDF and Sigmoid Laplace (see Fig.˜4(a))—and evaluate their per-predictive performance.
+To this end, we take the two best-performing methods on the Expected Calibration Error (ECE) metric [40]—NormCDF and Sigmoid Laplace (see Fig.˜4(a))—and evaluate their per-predictive performance.
 As Table˜1 shows, the closed-form predictive is always on par with the (unbiased) MC estimates on ImageNet while being cheaper.
@@
 We employ our methods with the closed-form predictives and second-order Dirichlet distributions to demonstrate their competitive nature while being more efficient than MC sampling.
-We first evaluate how calibrated the models are using the test NLL [gneiting2007strictly] and Expected Calibration Error (ECE) [naeini2015obtaining] metrics.
+We first evaluate how calibrated the models are using the test NLL [17] and Expected Calibration Error (ECE) [40] metrics.
 Fig.˜3(b) shows that on CIFAR-10, the NLL score of our closed-form predictives (Sigmoid, NormCDF) is consistently better than the corresponding Softmax results for all methods.
@@
 Fig.˜5(a) evidences this claim on ImageNet: our closed-form predictives either outperform or are on par with Softmax predictives.
-These results are in line with the findings of wightman_resnet_2021, who also recommend training with a per-class binary cross-entropy loss using the sigmoid activation function.
+These results are in line with the findings of Wightman et al.
+[51], who also recommend training with a per-class binary cross-entropy loss using the sigmoid activation function.
 Section˜K.6 shows that the BCE loss already yields performance gains in the model backbones (without the logit-space Gaussians).
@@
 the uncertainty estimator.
-As OOD inputs, we consider corrupted ImageNet-C [hendrycks2019benchmarking] samples.
+As OOD inputs, we consider corrupted ImageNet-C [18] samples.
 Fig.˜5(b) shows that our closed-form predictives outperform Softmax across all methods.
@@
 A.2 Gaussian NormCDF Integral
-Here, we derive the classical normCDF Gaussian integration formula [owen_table_1980, Eq. 10,010.8].
+Here, we derive the classical normCDF Gaussian integration formula [41, Eq. 10,010.8].
 For $\lambda>0$, $Z\sim\mathcal{N}(0,1)$ and $Y\sim\mathcal{N}(\mu,\sigma^{2})$ with $Y$ and $Z$ independent,
@@
 Hence matching $\rho$ and $\Phi$ to first order we get the approximation $\rho(y)\approx\Phi\left(\sqrt{\frac{\pi}{8}}y\right)$.
-So using Eq.˜32 we derive the probit approximation [spiegelhalter_sequential_1990, mackay_evidence_1992]
+So using Eq.˜32 we derive the probit approximation [47, 32]
 $\int_{\mathbb{R}}\rho(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)\stackrel{{\scriptstyle(1)}}{{\approx}}\int_{\mathbb{R}}\Phi\left(\sqrt{\frac{\pi}{8}}y\right)\;\mathcal{N}(\mu,\sigma^{2})(dy)=\Phi\left(\frac{\mu}{\sqrt{\frac{8}{\pi}+\sigma^{2}}}\right)\stackrel{{\scriptstyle(2)}}{{\approx}}\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right).$ (34)
@@
 A.4 Mean Field Approximation for Gaussian Softmax Integral
-For $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, the mean field approximation to the Gaussian softmax integral [lu_mean-field_2021] is obtained as follows
+For $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, the mean field approximation to the Gaussian softmax integral [31] is obtained as follows
 $\mathbb{E}_{\bm{Y}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\operatorname{softmax}_{c}\bm{Y}]$ $=\mathbb{E}\left[\left(2-C+\sum_{c^{\prime}\neq c}\rho(Y_{c}-Y_{c^{\prime}})^{-1}\right)^{-1}\right]$ (35)
@@
 $\mathbb{E}_{\bm{Y}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\operatorname{\mathbf{softmax}}\bm{Y}]\approx\operatorname{\mathbf{softmax}}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right).$ (36)
-(1) is the mean field approximation, and (3) uses the probit approximation Eq.˜34. [lu_mean-field_2021] provides two other variants of this approximation with other choices of approximation (2).
+(1) is the mean field approximation, and (3) uses the probit approximation Eq.˜34. [31] provides two other variants of this approximation with other choices of approximation (2).
 Appendix B Comparison with the Multinomial Probit Model
-In this appendix, we show that a model whose output activation is a composition of an element-wise normCDF activation $\bm{\Phi}$ and a normalisation $\bm{n}$ is distinct from the classical multinomial probit model [daganzo_chapter_1979].
+In this appendix, we show that a model whose output activation is a composition of an element-wise normCDF activation $\bm{\Phi}$ and a normalisation $\bm{n}$ is distinct from the classical multinomial probit model [9].
 Given a logit $\bm{y}\in\mathbb{R}^{C}$, the multinomial probit model sets
@@
 In this appendix, we provide (formal and informal) theoretical analyses of the quality of various predictive approximations.
-This complements the empirical analyses, for instance in the synthetic experiment (Fig.˜1) or in [daunizeau_semi-analytical_2017].
+This complements the empirical analyses, for instance in the synthetic experiment (Fig.˜1) or in [10].
 Due to its information-theoretic interpretation, a natural divergence to equip the probability simplex $\Delta^{C-1}$ with is the Kullback-Leibler (KL) divergence
@@
 C.2 Analysis of the Closed-Form Approximations (Theorem˜3.1)
-In this section we prove Theorem˜3.1 and discuss whether its underlying assumptions are fulfilled in practice. We start with a lemma.
+In this section we extend and prove Theorem˜3.1, and discuss whether its underlying assumptions are fulfilled in practice. We start with a lemma.
 Lemma C.1.
-Let $(X_{n})_{n\geq 1},(Y_{n})_{n\geq 1}$ be sequences of $(0,M)$-valued random variables for some $M>0$.
-Suppose $\sup_{n}\mathbb{E}[Y_{n}^{-8}]<\infty$ and that $\mathrm{Var}(Y_{n})\to 0$ as $n\to\infty$.
-Then
-$\left|\mathbb{E}\left[\frac{X_{n}}{Y_{n}}\right]-\frac{\mathbb{E}[X_{n}]}{\mathbb{E}[Y_{n}]}\right|=O(\mathrm{Var}(Y_{n})^{1/2})$ (48)
-as $n\to\infty$.
+Let $\mathcal{Y}$ be a family of $(0,M)$-valued random variables such that $\sup_{Y\in\mathcal{Y}}\mathbb{E}[Y^{-k}]<\infty$ for all $k>0$.
+Then for $X,Y$ $(0,M)$-valued random variables with $Y\in\mathcal{Y}$ and any $\epsilon>0$,
+$\left|\mathbb{E}\left[\frac{X}{Y}\right]-\frac{\mathbb{E}[X]}{\mathbb{E}[Y]}\right|=O\left(\mathrm{Var}(X)^{\frac{1}{2}-\epsilon}\mathrm{Var}(Y)^{\frac{1}{2}}+\mathrm{Var}(Y)^{1-\epsilon}\right)$ (48)
+as $\mathrm{Var}(Y)\to 0$ or $\mathrm{Var}(X),\mathrm{Var}(Y)\to 0$.
 Proof.
@@
 where $\eta(X)\in[\mathbb{E}[X],X]\text{ or }[X,\mathbb{E}[X]]$ and $\xi(Y)\in[\mathbb{E}[Y],Y]\text{ or }[Y,\mathbb{E}[Y]]$, using the Lagrange form of the remainder.
-So by the Cauchy-Schwarz inequality,
+So by the multivariate Hölder inequality we have for any $0<\epsilon<1/2$,
 $\left|\mathbb{E}\left[\frac{X}{Y}\right]-\frac{\mathbb{E}[X]}{\mathbb{E}[Y]}\right|$ $=\left|-\mathbb{E}\left[\frac{1}{\xi(Y)^{2}}(X-\mathbb{E}[X])(Y-\mathbb{E}[Y])\right]+\mathbb{E}\left[\frac{\eta(X)}{\xi(Y)^{3}}(Y-\mathbb{E}[Y])^{2}\right]\right|$ (50)
-$\leq\mathbb{E}\left[\frac{1}{\xi(Y)^{8}}\right]^{1/4}\mathbb{E}[(X-\mathbb{E}[X])^{4}]^{1/4}\mathbb{E}[(Y-\mathbb{E}[Y])^{2}]^{1/2}$
-$\quad+\mathbb{E}\left[\frac{\eta(X)^{2}}{\xi(Y)^{6}}\right]^{1/2}\mathbb{E}[(Y-\mathbb{E}[Y])^{4}]^{1/2}.$
+$\leq\mathbb{E}\left[\frac{1}{\xi(Y)^{2/\epsilon}}\right]^{\epsilon}\mathbb{E}[(X-\mathbb{E}[X])^{\frac{2}{1-2\epsilon}}]^{\frac{1}{2}-\epsilon}\mathbb{E}[(Y-\mathbb{E}[Y])^{2}]^{\frac{1}{2}}$
+$\quad+\mathbb{E}\left[\frac{\eta(X)^{1/\epsilon}}{\xi(Y)^{3/\epsilon}}\right]^{\epsilon}\mathbb{E}[(Y-\mathbb{E}[Y])^{\frac{2}{1-\epsilon}}]^{1-\epsilon}.$
 Note that
-$\eta(X_{n})$ $\leq M,$ (51)
-$\sup_{n}\mathbb{E}\left[\frac{1}{\xi(Y_{n})^{k}}\right]$ $\leq\sup_{n}\mathbb{E}\left[\frac{1}{\min(Y_{n}^{k},\mathbb{E}[Y_{n}]^{k})}\right]\leq\sup_{n}\mathbb{E}\left[\frac{1}{Y_{n}^{k}}+\frac{1}{\mathbb{E}[Y_{n}]^{k}}\right]<\infty\text{ for }k\in\{6,8\},$
-$\mathbb{E}[(X_{n}-\mathbb{E}[X_{n}])^{4}]$ $\leq M^{4},$
-$\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{4}]$ $\leq M^{2}\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{2}].$
+$\mathbb{E}[(X-\mathbb{E}[X])^{\frac{2}{1-2\epsilon}}]$ $\leq M^{\frac{4\epsilon}{1-2\epsilon}}\mathbb{E}[(X-\mathbb{E}[X])^{2}]=M^{\frac{4\epsilon}{1-2\epsilon}}\mathrm{Var}(X),$ (51)
+$\eta(X)^{1/\epsilon}$ $\leq M^{1/\epsilon},$
+$\sup_{Y\in\mathcal{Y}}\mathbb{E}\left[\frac{1}{\xi(Y)^{k}}\right]\leq\sup_{Y\in\mathcal{Y}}\mathbb{E}$ $\left[\frac{1}{\min(Y,\mathbb{E}[Y]^{k})}\right]\leq\sup_{Y\in\mathcal{Y}}\mathbb{E}\left[\frac{1}{Y^{k}}+\frac{1}{\mathbb{E}[Y]^{k}}\right]<\infty\text{ for }k\in\{2/\epsilon,3/\epsilon\}.$
 So from Eq.˜50 we obtain
-$\left|\mathbb{E}\left[\frac{X_{n}}{Y_{n}}\right]-\frac{\mathbb{E}[X_{n}]}{\mathbb{E}[Y_{n}]}\right|=O(\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{2}]^{1/2})=O(\mathrm{Var}(Y_{n})^{1/2})$ (52)
-as $n\to\infty$. ∎
+$\left|\mathbb{E}\left[\frac{X}{Y}\right]-\frac{\mathbb{E}[X]}{\mathbb{E}[Y]}\right|=O\left(\mathrm{Var}(X)^{\frac{1}{2}-\epsilon}\mathrm{Var}(Y)^{\frac{1}{2}}+\mathrm{Var}(Y)^{1-\epsilon}\right)$ (52)
+as $\mathrm{Var}(Y)\to 0$ or $\mathrm{Var}(X),\mathrm{Var}(Y)\to 0$. ∎
 Now recall our notation
@@
 $\bm{p}:=\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}],\quad\bm{\hat{p}}:=\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}},\quad\bm{\tilde{p}}:=\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}},$ (54)
-where $\bm{\Sigma}$ has $\bm{\sigma}$ as diagonal. We now restate Theorem˜3.1 with an explicit expression for $M(\mathcal{K})$:
+where $\bm{\Sigma}$ has $\bm{\sigma}$ as diagonal. We now restate Theorem˜3.1 with an explicit expression for $M(\mathcal{K})$ and an additional decay result for the $O$ term:
 Theorem C.2.
@@
 If $\Delta(\mathcal{K})>1$ then
-$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})\leq\log\left(\frac{1+\delta/u}{1-\Delta}\right)+O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$ (55)
-as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$.
+$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})\leq\log\left(\frac{1+\delta/u}{1-\Delta}\right)+O$ (55)
+where
+$O=O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)\qquad\text{as }\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$ (56)
+and
+$O=O\left(\max_{1\leq c\leq C}\mathrm{Var}(Q_{c})^{2-\epsilon}\right)\;\forall\>\epsilon>0\qquad\text{as }\mathrm{Var}(Q_{c})\to 0\;\forall\>1\leq c\leq C.$ (57)
 Proof.
@@
 $=\log\left(\frac{1+\delta/u}{1-\Delta}\right).$
-Now for (1), first note that $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$ where [^2] In the case $\varphi=\Phi$ and $Q_{c}\sim\Phi_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$, due to the fast decay of the tail of $\Phi$ we may have $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]=\infty$, and the proof strategy fails in that case.
-$Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$.
-By compactness of $\mathcal{K}$, we have in fact $\sup_{(\bm{\mu},\bm{\sigma})\in\mathcal{K}^{C}}\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$.
-Thus we can apply Lemma˜C.1 with $(X_{n})_{n\geq 1}$ a sequence of $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$ and $(Y_{n})_{n\geq 1}$ a sequence of $\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}$, to get
-$|p_{c}-\tilde{p}_{c}|=O\left(\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{1/2}\right)$ (58)
-as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$. Thus Taylor expanding each term of (1) around $p_{c}$ we get
+Now for (1), first note that $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-k}\right]<\infty$ where [^2] In the case $\varphi=\Phi$ and $Q_{c}\sim\Phi_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$, due to the fast decay of the tail of $\Phi$ we may have $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-k}\right]=\infty$, and the proof strategy fails in that case.
+$Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$ and $k>0$.
+By compactness of $\mathcal{K}$, we have in fact $\sup_{(\bm{\mu},\bm{\sigma}^{2})\in\mathcal{K}^{C}}\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-k}\right]<\infty$ for any $k>0$.
+Thus we can apply Lemma˜C.1 with $\mathcal{Y}=\left\{\sum_{c=1}^{C}Q_{c}:(\bm{\mu},\bm{\sigma}^{2})\in\mathcal{K}^{C}\right\}$ to get that for any $\epsilon>0$,
+$|p_{c}-\tilde{p}_{c}|=O\left(\mathrm{Var}(Q_{c})^{\frac{1}{2}-\epsilon}\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{\frac{1}{2}}+\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{1-\epsilon}\right)$ (60)
+as $\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)\to 0$ or $\mathrm{Var}(Q_{c}),\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)\to 0$.
+Thus Taylor expanding each term of (1) around $p_{c}$ we get
 $(1)$ $=\sum_{c=1}^{C}p_{c}\left(\frac{p_{c}-\tilde{p}_{c}}{p_{c}}+\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega_{c}(\tilde{p}_{c})^{2}}\right)$ (61)
 $=\underbrace{\sum_{c=1}^{C}p_{c}}_{=1}-\underbrace{\sum_{c=1}^{C}\tilde{p}_{c}}_{=1}+\sum_{c=1}^{C}p_{c}\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega_{c}(\tilde{p}_{c})^{2}}$
-$=\sum_{c=1}^{C}\frac{p_{c}}{2\omega_{c}(\tilde{p}_{c})^{2}}\cdot O\left(\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{1/2}\right)^{2}$
-$=O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$
-as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, where $\omega_{c}(\tilde{p}_{c})\in[p_{c},\tilde{p}_{c}]\text{ or }[\tilde{p}_{c},p_{c}]$, where we used that $p_{c}$, $\tilde{p}_{c}$, and hence $\omega_{c}(\tilde{p}_{c})$ is bounded away from 0 by compactness of $\mathcal{K}$.
+$=\sum_{c=1}^{C}\frac{p_{c}}{2\omega_{c}(\tilde{p}_{c})^{2}}\cdot O\Bigg(\mathrm{Var}(Q_{c})^{\frac{1}{2}-\epsilon}\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{\frac{1}{2}}+\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{1-\epsilon}\Bigg)^{2}$
+$=\sum_{c=1}^{C}\frac{p_{c}}{2\omega_{c}(\tilde{p}_{c})^{2}}\cdot O\Bigg(\mathrm{Var}(Q_{c})^{1-2\epsilon}\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)$
+$\quad+\mathrm{Var}(Q_{c})^{\frac{1}{2}-\epsilon}\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{\frac{3}{2}-\epsilon}+\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{2-2\epsilon}\Bigg)$
+$=O\Bigg(\sum_{c=1}^{C}\mathrm{Var}(Q_{c})^{1-2\epsilon}\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)$
+$\quad+\sum_{c=1}^{C}\mathrm{Var}(Q_{c})^{\frac{1}{2}-\epsilon}\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{\frac{3}{2}-\epsilon}+\mathrm{Var}\left(\sum_{c^{\prime}=1}^{C}Q_{c^{\prime}}\right)^{2-2\epsilon}\Bigg)$
+where $\omega_{c}(\tilde{p}_{c})\in[p_{c},\tilde{p}_{c}]\text{ or }[\tilde{p}_{c},p_{c}]$, where we used that $p_{c}$, $\tilde{p}_{c}$, and hence $\omega_{c}(\tilde{p}_{c})$ is bounded away from 0 by compactness of $\mathcal{K}$.
+We now distinguish two cases.
+If we only assume $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, then we see from Eq.˜61 that
+$(1)=O\Bigg(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\Bigg).$ (62)
+If we make the stronger assumption $\mathrm{Var}(Q_{c})\to 0$ for all $c$, then using the inequality
+$\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)=\sum_{1\leq c_{1},c_{2}\leq C}\operatorname{Cov}(Q_{c_{1}},Q_{c_{2}})\leq\sum_{1\leq c_{1},c_{2}\leq C}\mathrm{Var}(Q_{c_{1}})^{1/2}\mathrm{Var}(Q_{c_{2}})^{1/2}$ (63)
+we see that $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)=O\left(\max_{1\leq c\leq C}\mathrm{Var}(Q_{c})\right)$ and hence, from Eq.˜61,
+$(1)=O\left(\max_{1\leq c\leq C}\mathrm{Var}(Q_{c})^{2-2\epsilon}\right)$ (64)
+and the result follows since $\epsilon$ is arbitrary. ∎
 Figure C.1: Scatter plots of logit mean and variance pairs $(\mu_{c},\sigma_{c}^{2})$ on real models.
 We see that these seem to remain constrained to a compact set, with no much difference in the support as we increase the number of classes by an order of magnitude (from CIFAR-10 to CIFAR-100).
+Equation˜63 shows that requiring $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$ is weaker than requiring $\mathrm{Var}(Q_{c})\to 0$ for all $c$; but we see in Theorem˜C.2 that assuming the latter gives almost twice as strong a decay rate in $\operatorname{D_{KL}}$.
+This means that neither Eq.˜56 nor Eq.˜57 can be proven from the other; both results are thus of interest.
 A key assumption in Theorem˜C.2 is the logit means and variances being restricted to a compact set.
@@
 where all vector operations are element-wise.
-Multiplying out the denominators on the right-hand side of the equations of Eq.˜61, we obtain a system of two linear equations with two unknowns (for each $c$), which can be solved uniquely, yielding
+Multiplying out the denominators on the right-hand side of the equations of Eq.˜66, we obtain a system of two linear equations with two unknowns (for each $c$), which can be solved uniquely, yielding
 $\bm{\alpha}$ $=\frac{\mathbb{E}[\bm{Q}]-\mathbb{E}[\bm{Q}^{2}]}{\mathbb{E}[\bm{Q}^{2}]-\mathbb{E}[\bm{Q}]^{2}}\mathbb{E}[\bm{Q}],$ (67)
@@
 Here, we argue that the normCDF activation is an ideal choice for approximating Gaussian pushforwards with Beta distributions by interpreting Fig.˜D.1.
-This extends the work from [mackay_choice_1998], as it shows how one can make sense of the ‘right’ basis for performing Laplace approximations in the classification setting.
+This extends the work from [33], as it shows how one can make sense of the ‘right’ basis for performing Laplace approximations in the classification setting.
+Also see the discussion in [8].
 $\Phi_{*}\mathcal{G}(\mathbb{R})$, $\rho_{*}\mathcal{G}(\mathbb{R})$ the space of pushforwards of Gaussian distributions by normCDF and sigmoid respectively, and $\mathcal{B}((0,1))$, the space of Beta distributions, are statistical manifolds naturally equipped with Riemannian metrics, that is their respective Fisher information metrics.
@@
 The key observation is that $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$ are isometric to $\mathcal{G}(\mathbb{R})$.
-This is because $\Phi$ and $\rho$ are diffeomorphisms, so in particular sufficient statistics, and the Fisher information metric is invariant under sufficient statistics [ay_information_2017, Section 5.1.4].
+This is because $\Phi$ and $\rho$ are diffeomorphisms, so in particular sufficient statistics, and the Fisher information metric is invariant under sufficient statistics [3, Section 5.1.4].
 Visually, this means that $\varphi_{*}\mathcal{G}(\mathbb{R})$ has the same shape irrespectively of the diffeomorphism activation function $\varphi$.
@@
 Taking $\mu=0$, $\sigma^{2}=1$, $\alpha=\beta=1$ we get $\Phi_{0,1}=\Phi$ and $F_{1,1}=\operatorname{id}_{(0,1)}$, yielding $F_{\alpha,\beta}^{-1}\circ\Phi_{\mu,\sigma^{2}}=\Phi$.
-Now $\mathcal{G}(\mathbb{R})$, and hence $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$, is isometric to the hyperbolic plane ay_information_2017.
-We can embed this isometrically into Minkowski space with the classical hyperboloid model of the hyperbolic plane [reynolds_hyperbolic_1993],
+Now $\mathcal{G}(\mathbb{R})$, and hence $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\rho_{*}\mathcal{G}(\mathbb{R})$, is isometric to the hyperbolic plane Ay et al.
+[3, Example 3.1].
+We can embed this isometrically into Minkowski space with the classical hyperboloid model of the hyperbolic plane [43],
 $\mathcal{G}(\mathbb{R})\hookrightarrow\mathbb{R}^{2,1}.$ (68)
-For $\mathcal{B}((0,1))$, we use the isometric embedding from le_brigant_fisher-rao_2021:
+For $\mathcal{B}((0,1))$, we use the isometric embedding from Le Brigant et al. [26, Proposition 2]:
 $\mathcal{B}((0,1))$ $\hookrightarrow\mathbb{R}^{2,1}$ (69)
@@
 where $\eta(a):=\int_{1}^{a}\sqrt{\psi^{\prime}(r)}\;dr$ and $\psi$ is the digamma function.
-Finally, we choose our embedding Eq.˜63 for $\Phi_{*}\mathcal{G}(\mathbb{R})$ such that it intersects the embedding Eq.˜64 at a point, to highlight that the statistical manifolds $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ intersect at a point in the infinite-dimensional ambient space $\mathcal{M}((0,1))$, while $\rho_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ do not.
-Moreover, since $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$, our moment matching approximation Eq.˜61 is exact when $\mathrm{f}(x)$ is the standard Normal distribution.
+Finally, we choose our embedding Eq.˜68 for $\Phi_{*}\mathcal{G}(\mathbb{R})$ such that it intersects the embedding Eq.˜69 at a point, to highlight that the statistical manifolds $\Phi_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ intersect at a point in the infinite-dimensional ambient space $\mathcal{M}((0,1))$, while $\rho_{*}\mathcal{G}(\mathbb{R})$ and $\mathcal{B}((0,1))$ do not.
+Moreover, since $\Phi_{*}\mathcal{N}(0,1)=B(1,1)$, our moment matching approximation Eq.˜66 is exact when $\mathrm{f}(x)$ is the standard Normal distribution.
 Appendix E List of Predictive and Dirichlet Parameters Formulas
@@
 Mean Field (Section˜A.4)
-The Mean Field predictive [lu_mean-field_2021] uses the following approximation for the true predictive:
+The Mean Field predictive [31] uses the following approximation for the true predictive:
 $\bm{\hat{p}}:=\bm{\operatorname{\mathbf{softmax}}}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right).$ (71)
 Laplace Bridge
-The Laplace Bridge predictive [pmlr-v180-hobbhahn22a] approximates the true predictive as follows:
+The Laplace Bridge predictive [19] approximates the true predictive as follows:
 $\bm{\hat{p}}:=\frac{\frac{1}{\bm{\tilde{\sigma}}^{2}}\left(1-\frac{2}{C}+\frac{e^{\bm{\tilde{\mu}}}}{C^{2}}\sum_{c=1}^{C}e^{-\tilde{\mu}_{c}}\right)}{\sum_{c=1}^{C}\frac{1}{\tilde{\sigma}^{2}_{c}}\left(1-\frac{2}{C}+\frac{e^{\bm{\tilde{\mu}}}}{C^{2}}\sum_{c^{\prime}=1}^{C}e^{-\tilde{\mu}_{c^{\prime}}}\right)}$ (72)
@@
 Shekhovtsov & Flach.
-The shekhovtsov_feed-forward_2018 approximation of the Gaussian-softmax integral yields:
+The Shekhovtsov and Flach [46] approximation of the Gaussian-softmax integral yields:
 $\bm{\hat{p}}=\frac{\bm{n}}{\bm{n}+\big(A-\bm{b}\big)^{s/\bm{s}}}$ (74)
@@
 E.2.1 Softmax
-The Laplace Bridge method [pmlr-v180-hobbhahn22a] uses the following Dirichlet parameters:
+The Laplace Bridge method [19] uses the following Dirichlet parameters:
 $\bm{\gamma}:=\frac{1}{\bm{\tilde{\sigma}}^{2}}\left(1-\frac{2}{C}+\frac{e^{\bm{\tilde{\mu}}}}{C^{2}}\sum_{c=1}^{C}e^{-\tilde{\mu}_{c}}\right)$ (80)
@@
   • Do we have to sacrifice performance for sample-free predictives?
-To answer the first question, we evaluate our closed-form predictives (Sigmoid, NormCDF, Exp) and moment-matched Dirichlet distributions against softmax models equipped with approximate inference tools (Laplace Bridge [pmlr-v180-hobbhahn22a], Mean Field [lu_mean-field_2021], Monte Carlo sampling).
-We consider Heteroscedastic Classifiers (HET) [collier2021correlated], Spectral-Normalised Gaussian Processes (SNGP) [liu2020simple], and last-layer Laplace approximation methods [daxberger2021laplace] as backbones (see Appendix˜J for details).
-The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [deng2009imagenet] and CIFAR-10 [krizhevsky2009learning] on five metrics aligning with practical needs from uncertainty estimates [mucsanyi2023trustworthy]:
+To answer the first question, we evaluate our closed-form predictives (Sigmoid, NormCDF, Exp) and moment-matched Dirichlet distributions against softmax models equipped with approximate inference tools (Laplace Bridge [19], Mean Field [31], Monte Carlo sampling).
+We consider Heteroscedastic Classifiers (HET) [7], Spectral-Normalised Gaussian Processes (SNGP) [30], and last-layer Laplace approximation methods [11] as backbones (see Appendix˜J for details).
+The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [12] and CIFAR-10 [25] on five metrics aligning with practical needs from uncertainty estimates [38]:
   1. 1.
@@
 See Appendix˜I for details.
-For ImageNet, we treat ImageNet-C [hendrycks2019benchmarking] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
+For ImageNet, we treat ImageNet-C [18] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
 For the second question, we consider fixed (method, activation) pairs and test whether our methods perform on par with the Monte Carlo sampled predictives.
@@
 For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following mucsányi2024benchmarking.
-We train with the LAMB optimiser [you2019large] using a batch size of $128$ and gradient accumulation across $16$ batches, resulting in an effective batch size of $2048$, following tran2022plex.
+We train with the LAMB optimiser [55] using a batch size of $128$ and gradient accumulation across $16$ batches, resulting in an effective batch size of $2048$, following Tran et al.
+[50].
 We further use a cosine learning rate schedule with a single warmup epoch using a warmup learning rate of $0.0001$.
@@
 During training, we keep track of the best-performing checkpoint on the validation set and load it before testing.
-We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [shahriari2015taking] in Weights & Biases [wandb] based on the negative log-likelihood.
+We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [45] in Weights & Biases [5] based on the negative log-likelihood.
 On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs.
-The only exceptions are the SNGP models that are trained for 125 epochs [liu2020simple].
+The only exceptions are the SNGP models that are trained for 125 epochs [30].
 We train with Momentum SGD using a batch size of $128$ and no gradient accumulation.
@@
 I.1 Log Probability Proper Scoring Rule for the Predictive
-First, we briefly discuss proper and strictly proper scoring rules over general probability measures based on [mucsanyi2023trustworthy].
+First, we briefly discuss proper and strictly proper scoring rules over general probability measures based on [38].
 Consider a function $S\colon\mathcal{Q}\times\mathcal{Y}\to\mathbb{R}$ where $\mathcal{Q}$ is a family of probability distributions over the space $\mathcal{Y}$, called the label space.
@@
 I.2 Expected Calibration Error
-To set up the required quantities for the Expected Calibration Error (ECE) metric [naeini2015obtaining], we follow the steps below, based on [mucsanyi2023trustworthy].
+To set up the required quantities for the Expected Calibration Error (ECE) metric [40], we follow the steps below, based on [38].
   1. 1.
@@
     $\tilde{c}_{n}\in\left(\frac{m-1}{M},\frac{m}{M}\right].$ (97)
-The Expected Calibration Error (ECE) metric [naeini2015obtaining] is then defined as
+The Expected Calibration Error (ECE) metric [40] is then defined as
 $\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\right|$ (98)
@@
 […]
-One can show that this is indeed a strictly proper scoring rule [mucsanyi2023trustworthy].
+One can show that this is indeed a strictly proper scoring rule [38].
 I.4 Accuracy
@@
 J.1 Spectral Normalised Gaussian Process
-Spectral normalised Gaussian processes (SNGP) [liu2020simple] use spectral normalization of the parameter tensors for distance-awareness and a last-layer Gaussian process approximated by Fourier features to capture uncertainty.
+Spectral normalised Gaussian processes (SNGP) [30] use spectral normalization of the parameter tensors for distance-awareness and a last-layer Gaussian process approximated by Fourier features to capture uncertainty.
 For an input $x\in\mathcal{X}$ and number of classes $C$, they predict a $C$-variate Gaussian distribution
@@
 The method applies spectral normalization to the hidden weights in each layer using a power iteration scheme with a single iteration per batch to obtain the largest singular value.
-liu2020simple claim this helps with input distance awareness.
+Liu et al.
+[30] claim this helps with input distance awareness.
 J.2 Heteroscedastic Classifier
-Heteroscedastic classifiers (HET) [collier2021correlated] construct a Gaussian distribution in the logit space to model per-input uncertainties:
+Heteroscedastic classifiers (HET) [7] construct a Gaussian distribution in the logit space to model per-input uncertainties:
 $\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x)),$ (107)
@@
 J.3 Laplace Approximation
-The Laplace approximation [daxberger2021laplace] approximates the posterior $p(\bm{\theta}\mid\mathcal{D})$ over the network parameters $\bm{\theta}$ for a Gaussian prior $p(\bm{\theta})$ and likelihood defined by the network architecture by a Gaussian.
+The Laplace approximation [11] approximates the posterior $p(\bm{\theta}\mid\mathcal{D})$ over the network parameters $\bm{\theta}$ for a Gaussian prior $p(\bm{\theta})$ and likelihood defined by the network architecture by a Gaussian.
 In its simplest form, it uses the maximum a posteriori (MAP) weights $\bm{\theta}_{\text{MAP}}\in\mathbb{R}^{P}$ as the mean and the inverse Hessian of the regularised loss over the training set $\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})=\mathcal{L}(\bm{\theta};\mathcal{D})+\lambda\|\bm{\theta}\|_{2}^{2}$ evaluated at the MAP as the covariance matrix:
 $\mathcal{N}\left(\bm{\theta}_{\text{MAP}},\left(\left.\frac{\partial^{2}\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})}{\partial\theta_{i}\partial\theta_{j}}\right|_{\bm{\theta}_{\text{MAP}}}\right)^{-1}\right)=\mathcal{N}\left(\bm{\theta}_{\text{MAP}},\left(\left.\frac{\partial^{2}\mathcal{L}(\bm{\theta};\mathcal{D})}{\partial\theta_{i}\partial\theta_{j}}\right|_{\bm{\theta}_{\text{MAP}}}+\lambda\bm{I}_{P}\right)^{-1}\right).$ (109)
-This is a locally optimal post-hoc Gaussian approximation of the true posterior $p(\bm{\theta}\mid\mathcal{D})$ based on a second-order Taylor approximation. For details, see [tatzel2025debiasing].
+This is a locally optimal post-hoc Gaussian approximation of the true posterior $p(\bm{\theta}\mid\mathcal{D})$ based on a second-order Taylor approximation. For details, see [48].
 For deep neural networks, the Hessian matrix is often replaced with the Generalised Gauss-Newton (GGN) matrix.
-The GGN is guaranteed to be positive semidefinite even for suboptimal weights and has efficient approximation schemes, such as Kronecker-Factored Approximate Curvature [martens2015optimizing] or low-rank approximations.
+The GGN is guaranteed to be positive semidefinite even for suboptimal weights and has efficient approximation schemes, such as Kronecker-Factored Approximate Curvature [36] or low-rank approximations.
 Denoting our curvature estimate of choice by $\bm{G}$, the logit-space Gaussian is obtained by pushing forward the weight-space Gaussian measure through the linearised model around $\bm{\theta}_{\text{MAP}}$.
@@
 where $\bm{J}_{\bm{\theta}_{\text{MAP}}}\bm{f}(x)\in\mathbb{R}^{C\times P}$ is the model Jacobian matrix.
-We use a last-layer KFAC Laplace variant in our experiments and use the full training set for calculating the GGN instead of a mini-batch based on recent works on the bias in mini-batch estimates [tatzel2025debiasing].
+We use a last-layer KFAC Laplace variant in our experiments and use the full training set for calculating the GGN instead of a mini-batch based on recent works on the bias in mini-batch estimates [48].
 Appendix K Additional Results
@@
 K.2 Vision Transformer Results on ImageNet
-Table˜K.3 shows results on ViT Little [dosovitskiy2020image] backbones from the timm [rw2019timm] library on the ImageNet validation set.
+Table˜K.3 shows results on ViT Little [14] backbones from the timm [52] library on the ImageNet validation set.
 As in the main paper, the hyperparameters are optimised using a ten-step Bayesian hyperparameter optimization scheme of Weights and Biases.
@@
 Calibration and Proper Scoring
-We first evaluate calibration using the log probability scoring rule [gneiting2007strictly] and the Expected Calibration Error (ECE) metric [naeini2015obtaining].
+We first evaluate calibration using the log probability scoring rule [17] and the Expected Calibration Error (ECE) metric [40].
 Fig.˜3(b) shows that on CIFAR-10, the score of our closed-form predictives (Sigmoid, NormCDF) are consistently better than the corresponding softmax results for all methods.
@@
 The most accurate method is Sigmoid HET.
-These results support the findings of wightman_resnet_2021 that showcase desirable training dynamics of the class-wise cross-entropy loss.
+These results support the findings of Wightman et al.
+[51] that showcase desirable training dynamics of the class-wise cross-entropy loss.
 Out-of-Distribution Detection
@@
 K.8 Comparison to Evidential Deep Learning Methods
-Table˜K.8 provides results for two evidential deep learning methods, EDL [sensoy2018evidential] and PostNet [charpentier2020posterior].
+Table˜K.8 provides results for two evidential deep learning methods, EDL [44] and PostNet [6].
 (We adopt the naming convention of [mucsányi2024benchmarking].) We make the comparison because these methods also yield closed-form predictives and second-order quantities of interest, as they build Dirichlet distributions whose advantages we discuss in Section˜4.1.
@@
 Compared with the results presented in the main text, our approach outperforms the two EDL methods across all metrics, significantly so on NLL and OOD AUROC.
-The results on the latter metric indicate that our moment-matched Dirichlet distributions capture second-order epistemic information more effectively than the loss-based EDL methods, which aligns with the findings of NEURIPS2022_bc1d640f.
+The results on the latter metric indicate that our moment-matched Dirichlet distributions capture second-order epistemic information more effectively than the loss-based EDL methods, which aligns with the findings of Bengs et al.
+[4].
 Table K.8: Evaluation results of PostNet and EDL across CIFAR-10, CIFAR-100, and ImageNet. Values are mean $\pm$ two standard deviations.